### v1

# High Dimensional Ensemble Kalman Filter

Abstract
The ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of earth science, engineering and beyond. However, there are several unknown theoretical aspects of the EnKF, especially when the state variable is of high dimensional accompanied with high resolution observation and physical models. This paper first proposes several high dimensional EnKF methods which provide consistent estimators for the important forecast error covariance and the Kalman gain matrix. It then studies the theoretical properties of the EnKF under both the fixed and high dimensional state variables, which provides the mean square errors of the analysis states to the underlying oracle states offered by the Kalman filter and gives the much needed insight into the roles played by forecast error covariance on the accuracy of the EnKF. The accuracy of the data assimilation under the misspecified physical model is also considered. Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed high dimensional EnKF algorithms perform better than the standard EnKF methods as they provide more robust and accurate assimilated results. The high dimensional EnKF is applied to assimilate sea temperature in Northwest Pacific, which showed more accurate out-sample performance than the existing methods.
Data Assimilation,
Ensemble Kalman Filter,
High Dimension Imputation,
keywords:
[…]
, and
1 Introduction
Kalman Filter (Kalman, 1960) is a fundamental approach for forecasting and data assimilation for a partially observed dynamic system. Kalman Filter and its later developments (Ensembled Kalman Filter, three and four–dimensional variational assimilation (3D-Var, 4D-Var) (Pailleux, 1990; Courtier et al., 1998) are based on the optimal estimation theory under the multivariate normal distribution. Data assimilation is a statistical procedure that revises a forecast based on the physical model to the state variables of the dynamic system upon receiving new observations. It combines information from numerical physical model forecasts with observed data to obtain the best possible description of a dynamical system and its uncertainty (Evensen, Vossepoel and van Leeuwen, 2022). The assimilated state variables at a time are used to generate the forecast for the next time step. For chaotic systems sensitive to initial values, data assimilation can significantly improve the accuracy of the forecast by using optimized state variables as initial values (Lorenz, 1963).
Although much applied in numerical prediction for weather and oceanography, Kalman Filter and its later versions are the methodological core for scientific algorithms over a range of fields, including hydrology, remote sensing data utilization, air quality prediction, carbon data assimilation, robotic control and automatic driving. Data assimilation methods also provide the underlying algorithms and predictive techniques for building digital twins for social and economic systems.
The limitations of the Kalman Filter are (i) it assumes a linear model system which is known to be restrictive for many non-linear dynamic systems; and (ii) the forecast error covariance matrix and Kalman Gain matrix are expensive to store and update when the dimension of the state vector is large. The Ensemble Kalman Filter (EnKF; Evensen, 1994) was proposed to overcome these limitations, where the forecast error covariance matrix is estimated by the sample covariance based on generated forecast ensembles. The EnKF only requires storing and operating on reduced size state vectors and the ensembles can be generated parallelly. These advantages have made the EnKF an effective data assimilation scheme and has been widely used in atmospheric and oceanic sciences, and engineering.
Estimating the forecast error covariance matrices is a key in the Kalman Filter and the EnKF schemes, which is made in the EnKF by the sample covariance matrix of the forecast ensembles. Past works on the EnKF found that the sampling error can lead to underestimation of the forecast error covariance and can eventually cause the filter to diverge (Anderson and Anderson, 1999; Constantinescu et al., 2007). The filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation. One reason for the filter divergence is the variance of the ensemble forecast error becomes too small when there are unaccounted model errors (Lorenc, 2003); the other is due to the substantial mismatch between the estimated and true error covariance matrices (Houtekamer and Mitchell, 1998; Constantinescu et al., 2007).
Along with advances in remote sensing technology in measuring the earth system, high spatial resolution observations of the earth system are increasingly available, which leads to high dimensional geophysical models with high dimensional state vectors (Stroud et al., 2010; Guinehut et al., 2012; Lellouche et al., 2013). High resolution data assimilation and forecasting techniques are much needed in modern meteorology and atmospheric environment systems. For example, in global CO2 calculation, the dimension of the state variable can reach tens of millions. The high dimensional state vectors lead to high dimensional  forecast error covariance and Kalman Gain matrices. The high dimensionality is known to cause the non-convergence of the forecast error covariance matrix (Bai and Yin, 1993). In practice, the high dimensional effect is felt in the form of the filter divergence and unreliable data assimilation. Geophysicists have proposed two methods to counter the problems. One is the localization method (Furrer, Genton and Nychka, 2006; Furrer and Bengtsson, 2007) that sets covariances to zero if the geographical distances between two sites are bigger than a threshold; and the other is the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term to avoid filter divergence. However, these methods lack support of statistical theory.
The last two decades have witnessed rapid development in high dimensional statistical inference. Consistent high dimensional  covariance estimators have been proposed, which include the banding, tapering and thresholding estimators (Bickel and Levina, 2008a, b; Cai, Zhang and Zhou, 2010). Both the banding and tapering estimators require a bandwidth parameter, and the thresholding estimator needs the threshold level. Qiu and Chen (2015) proposed bandwidth and thresholding level selection methods by minimizing an empirical estimate of the expected squared Frobenius norms of the estimation error matrix.
This paper proposes several high dimensional EnKF methods that utilize the high dimensional covariance estimators, and conducts a theoretical analysis of the proposed high dimensional EnKF. We also provide a theoretical analysis of the performance of the EnKF under the fixed dimensional setting since such a study appears lacking and is important. It is shown that with the use of the high dimensional covariance matrix estimators, consistent estimation of the Kalman Gain matrix and the underlying analysis states can be attained by the proposed high dimensional EnKF procedure. Specifically, we establish the mean square error (MSE) bounds of the analysis states from the EnKF methods to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models.
Our analysis shows that consistent estimator of the forecast error covariance leads to the consistency of the Kalman Gain matrix by establishing an inequality that bounds the error of estimation of the Kalman Gain matrix by that of the forecast error covariance, which implies that, even at the high dimensional setting, the MSEs of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators. Thus, consistent high dimensional covariance estimators would make the high dimensional EnKF consistent under certain conditions. Indeed, our theoretical studies show that both the forecast and analysis states of the EnKF converge to their underlying counterparts of the Kalman Filter if the forecast error covariance can be estimated consistently.
The investigation of the high dimensional EnKF under the misspecified model gives a more realistic description of the performance of the EnKF. Under the misspecified model, the MSE not only depends on the estimation error of the forecast error covariance matrices but also the discrepancy between the misspecified and the true model. Simulation studies on the Lorenz-96 model and the Shallow Water Equation model suggest that the proposed high dimensional EnKF algorithm could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method. The application to the sea temperature data assimilation in Northwest Pacific provides a comprehensive and accurate description of the spatiotemporal distribution of the three-dimensional sea temperature field by combining the ocean model with satellite data, which also illustrates the superiority and usefulness of the proposed high dimensional EnKF. This means the proposed high dimensional EnKF has wide application prospects in earth science like global carbon calculation, weather forecasting and the production of high-resolution data sets.
The paper is structured as follows. Section 2 reviews the Kalman Filter and the EnKF. Theoretical properties of the EnKF for the fixed dimensional case are given in Section 3. Section 4 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman gain matrix, and establishes the theoretical properties of the EnKF for high dimensional case. The effects of imperfect or misspecified models on the accuracy of the EnKF are provided in Section 5. Results of empirical data assimilation based on two well known geophysical models are reported in Section 6 and the application to the sea temperature data assimilation is presented in Section 7 , followed by a conclusion in Section 8. Details of the proposed high dimensional EnKF algorithms, proofs of the theoretical results and additional results are given in the Supplementary Material (SM).
2 Review on Kalman and Ensemble Kalman Filters
The Kalman Filter (Kalman, 1960) assumes the following linear Gaussian state-space model (which was generalized to nonlinear models in the EnKF)
$\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathbf{M}_{t}\mathbf{x}_{t-1}+\mathbf{w}_{t},\quad\ \quad\mathbf{w}_{t}\sim\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\,\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}(\mathbf{0},\mathbf{R}_{t}),\end{array}\right.$ (2.1)
where $\mathbf{x}_{t}$ is the $p-$dimensional not fully observed state vector of a dynamic system at time $t$, $\mathbf{M}_{t}$ is the linear model operator (also called evolution or forecast operator), $\mathbf{w}_{t}$ is the error term that follows a zero-mean multivariate normal distribution with $\mathbf{Q}_{t}$ as the covariance, $\mathbf{y}_{t}$ is the $q-$dimensional observation vector at time $t$, $\mathbf{H}_{t}$ is a linear observation operator which relates the state $\mathbf{x}_{t}$ to the observation $\mathbf{y}_{t}$, and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix.
Assuming $\mathbf{M}_{t}$, $\mathbf{Q}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are known, the Kalman Filter provides a two-step alternative updating framework based on the theory of multivariate normal distribution. In the forecast step, suppose that $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p}\left(\mathbf{x}_{t-1}^{a},\boldsymbol{\Sigma}_{t-1}^{a}\right)$, where $\mathbf{x}_{t-1}^{a}$ is the analysis state vector at the previous time $t-1$ and $\mathbf{y}_{1:t-1}$ denotes the observation $\{\mathbf{y}_{i}\}_{i=1}^{t-1}$ up to time $t-1$, the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{y}_{1:t-1}$ is
$\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p}\big{(}\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f}\big{)}\equiv{\mathrm{N}}_{p}\big{(}{\mathbf{M}}_{t}\mathbf{x}_{t-1}^{a},{\mathbf{M}}_{t}\boldsymbol{\Sigma}_{t-1}^{a}{\mathbf{M}}_{t}^{\mathrm{T}}+\mathbf{Q}_{t}\big{)},$ (2.2)
where $\mathbf{x}_{t}^{f}=\mathbf{M}_{t}\mathbf{x}_{t-1}^{a}$ is the forecast state and
$\boldsymbol{\Sigma}_{t}^{f}=\mathbf{M}_{t}\boldsymbol{\Sigma}_{t-1}^{a}\mathbf{M}_{t}^{\mathrm{T}}+\mathbf{Q}_{t},$ (2.3)
is the variance of the forecast error $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$, also called the forecast error covariance matrix.
In the update (assimilation) step, the goal is to update $\mathbf{x}_{t}^{f}$ according to the new observation $\mathbf{y}_{t}$ and to reduce the difference between the true state $\mathbf{x}_{t}$ and the forecast state $\mathbf{x}_{t}^{f}$. The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is
$\left(\begin{array}[]{c}\mathbf{x}_{t}\\ \mathbf{y}_{t}\end{array}\right)\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p+q}\left(\left(\begin{array}[]{c}\mathbf{x}_{t}^{f}\\ \mathbf{H}_{t}\mathbf{x}_{t}^{f}\end{array}\right),\left(\begin{array}[]{cc}\boldsymbol{\Sigma}_{t}^{f}&\boldsymbol{\Sigma}_{t}^{f}\mathbf{H}_{t}^{T}\\ \mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}&\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}\mathbf{H}_{t}^{T}+\mathbf{R}_{t}\end{array}\right)\right),$ (2.4)
which implies that
$\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right)$ (2.5)
where, by the conditional Gaussian distribution formulae,
$\mathbf{x}_{t}^{a}$ $={\mathbf{x}}_{t}^{f}+\mathbf{K}_{t}\big{(}\mathbf{y}_{t}-\mathbf{H}_{t}{\mathbf{x}}_{t}^{f}\big{)},\,\boldsymbol{\Sigma}_{t}^{a}=\left(\mathbf{I}_{p}-\mathbf{K}_{t}\mathbf{H}_{t}\right)\boldsymbol{\Sigma}_{t}^{f}\text{ and }$ (2.6)
$\mathbf{K}_{t}$ $={\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1},$ (2.7)
which are, respectively, the analysis (assimilation) state and the associated analysis covariance matrix, and the Kalman gain matrix. It is noted that the analysis state $\mathbf{x}_{t}^{a}$ is the minimum variance unbiased estimator for the mean of $\mathbf{x}_{t}$ under the linear state-space model (2.1).
However, models in earth science or engineering are often nonlinear, which implies that the evolution model in (2.1) may not be linear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$, where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2). Besides, for large dynamic systems, the dimension $p$ of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store. To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error variance matrix is no longer updated as (2.3) but by the sample covariance based on $n$ forecast ensembles. It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing, and thus improves the computing efficiency and reduces the storage cost.
The EnKF assumes possibly nonlinear model system with linear observation operator:
$\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1}\right)+\mathbf{w}_{t},\quad\quad\ \mathbf{w}_{t}\sim\mathrm{N}_{p}(\mathbf{0},\mathbf{Q}_{t}),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\quad\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}\left(\mathbf{0},\mathbf{R}_{t}\right),\end{array}\right.$ (2.8)
where $\mathcal{M}_{t}$ is the possibly nonlinear model operator, $\mathbf{H}_{t}$ is a linear observation operator and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix. It further assumes that the forecast error $\boldsymbol{\eta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$ follows a Gaussian distribution given $\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$, where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)$. Here $\mathcal{M}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are assumed to be known for the time being as in the Kalman Filter.
A key feature of the EnKF is to generate an ensemble of perturbed forecast states $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ of size $n$ from the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{y}_{1:t-1}$, which are used to estimate the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$. Given the analysis ensemble at the previous time step $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$, the EnKF consists of the following three steps.
  • Step (i):
    Run the model forward in time to get the perturbed forecast state and their average:
    $\mathbf{x}_{t,j}^{f}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1,j}^{a}\right)+\mathbf{w}_{t,j}\text{ and }\hat{\mathbf{x}}_{t}^{f}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f},\text{ where }\mathbf{w}_{t,j}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right).$ (2.9)
  • Step (ii):
    Estimate the forecast error covariance matrix $\Sigma_{t}^{f}$ by the sample covariance matrix of the forecast ensemble, namely
    $\hat{\boldsymbol{\Sigma}}_{t}^{f}=(n-1)^{-1}\sum_{j=1}^{n}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}^{\mathrm{T}}\triangleq\mathbf{S}_{t}^{f},$ (2.10)
    and obtain an estimate of the Kalman gain matrix $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$. Here we note computationally efficient expressions
    $\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}$ $=(n-1)^{-1}\sum_{j=1}^{n}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}^{\mathrm{T}}\,\text{and}$ (2.11)
    $\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}$ $=(n-1)^{-1}\sum_{j=1}^{n}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}^{\mathrm{T}}$
    since the dimension $q$ of the observation is much smaller than that ($p$) of the state vector.
  • Step (iii):
    Calculate the perturbed residuals $\mathbf{d}_{t,j}=\mathbf{y}_{t}+\boldsymbol{\varepsilon}_{t,j}-{\mathbf{H}}_{t}\mathbf{x}_{t,j}^{f}$ where $\{\boldsymbol{\varepsilon}_{t,j}\}_{j=1}^{n}\stackrel{{\scriptstyle IID}}{{\sim}}{\mathrm{N}}\left(\mathbf{0},\mathbf{R}_{t}\right)$, and let $\mathbf{d}_{t}=n^{-1}\sum_{j=1}^{n}\mathbf{d}_{t,j}$. Compute the perturbed analysis states and their average, the assimilated (analysis) state,
    $\mathbf{x}_{t,j}^{a}$ $=\mathbf{x}_{t,j}^{f}+\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}\mathbf{d}_{t,j},$ (2.12)
    $\hat{\mathbf{x}}_{t}^{a}$ $=\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{a}=\hat{\mathbf{x}}_{t}^{f}+\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}\mathbf{d}_{t},$
    which will be used for the data assimilation at the next time step.
In practice, the algorithm is initialized at time $t=0$ by drawing $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ independently from ${\mathrm{N}}_{p}\left(\mathbf{x}_{0}^{a},\Sigma_{0}^{a}\right)$ where $\mathbf{x}_{0}^{a}$ and $\Sigma_{0}^{a}$ are either chosen based on the physical model or by pre-running the physical model for a period of time. Then, the ensemble is propagated forward through time, alternating between the forecast and the analysis steps, according to the three-step procedure above. If $\mathbf{R}_{t}$ and $\mathbf{Q}_{t}$ are unknown, their estimates can be entertained to substitute them in (2.9) and (2.12).
The EnKF does not require the expensive recursions (2.3) and (2.6)-(2.7) used in the Kalman Filter. Instead, it estimates the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance or the reduced dimension counterpart (2.11) via the observation operator $\mathbf{H}_{t}$.
For the case with non-linear model operator as in (2.8), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.6)-(2.7), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance defined by $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$. In fact, $\mathbf{x}_{t}^{a}$ defined by (2.6)-(2.7) is the solution to the minimizing problem (2.13) with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$. For the ease of reference, and to avoid differentiating between the linear and the non-linear cases, we also call $\mathbf{x}_{t}^{a}$ as the analysis state of the Kalman filter even in the non-linear model case. The next two sections will establish the convergence of the EnKF analysis state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman filter.
It is noted that the update formula (2.12) can be obtained by minimizing
$J(\mathbf{x})=\frac{1}{2}\left\{(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\right\}$ (2.13)
with respect to $\mathbf{x}$. To appreciate this, note that by setting $\partial J(\mathbf{x})/{\partial\mathbf{x}}=\mathbf{0}$ and by the Sherman-Morrison-Woodbury formula, we get
$\mathbf{x}$ =$=\left\{\hat{\boldsymbol{\Sigma}}_{t}^{f}-\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{R}_{t}+\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{)}^{-1}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\right\}\left\{(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}\mathbf{\hat{x}}_{t}^{f}+\mathbf{H}_{t}^{\mathrm{T}}\mathbf{R}_{t}^{-1}\mathbf{y}_{t}\right\}$
$=\mathbf{\hat{x}}_{t}^{f}+\mathbf{\hat{K}}_{t}\big{(}\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}_{t}^{f}\big{)},$
which is the data assimilation formula (2.12). The objective function (2.13) is related to that of the 3D-Var or 4D-Var, but not the same. However, over the same time interval, under the assumption of perfect model with the same inputs (initial mean and covariance matrix, observation error covariance), the 4D-Var analysis at the end is equal to the analysis of the Kalman filter; see Lorenc (2003), Hamill (2006) and Constantinescu et al. (2007) for details.
3 Consistency for fixed dimension case
We first provide the upper bound of the estimation error of the estimated Kalman gain matrix and then establish the convergence of the EnKF analysis states to the underlying Kalman Filter under both the linear and the non-linear model operators for the fixed dimensional state vectors with a convergence rate. Indeed, it appears that such results are lacking in the literature. Additionally, the error bounds could further be extended to the high dimension cases, which will be discussed later in Section 4.
[…]
Assumption 1.
(i) The model operator $\mathcal{M}_{t}$ is Lipschitz continuous, that is $\|\mathcal{M}_{t}(\mathbf{x})-\mathcal{M}_{t}(\mathbf{x}^{\prime})\|\leq\ell_{0}\|\mathbf{x}-\mathbf{x}^{\prime}\|$ for any $\mathbf{x},\mathbf{x}^{\prime}\in R^{p}$ for a positive constant $\ell_{0}$. (ii) The model operator $\mathcal{M}_{t}$ can be locally linearized and satisfies $\|\mathcal{M}_{t}(\mathbf{x})-\mathcal{M}_{t}(\mathbf{x}^{\prime})\|\leq\|\dot{M}_{t}\|\|\mathbf{x}-\mathbf{x}^{\prime}\|$ for $\mathbf{x},\mathbf{x}^{\prime}\in R^{p}$ and $\|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\|\mathbf{h}\|$ for a small positive $\mathbf{h}$, where $\dot{M}_{t}$ is a matrix related to the linear expansion of $\mathcal{M}_{t}$ at $\mathbf{x}$ and $\|\dot{M}_{t}\|\leq\ell_{t}$ for some positive constants $\ell_{t}$.
Assumption 2.
The model error covariance matrix $\mathbf{Q}_{t}$, the observation error covariance matrix $\mathbf{R}_{t}$ and the initial analysis covariance matrix $\boldsymbol{\Sigma}_{0}^{a}$ satisfy $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{Q}_{t}\right)\leq\lambda_{\max}\left(\mathbf{Q}_{t}\right)\leq 1/\varepsilon_{0}$, $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{R}_{t}\right)\leq\lambda_{\max}\left(\mathbf{R}_{t}\right)\leq 1/\varepsilon_{0},$ and $\lambda_{\max}\left(\boldsymbol{\Sigma}_{0}^{a}\right)\leq 1/\varepsilon_{0},$ for a positive $\varepsilon_{0}$, respectively, where $\lambda_{\max}$ and $\lambda_{\min}$ are the maximum and minimum eigenvalue operators, respectively.
Assumption 3.
The observation operator $\mathbf{H}_{t}$ satisfies $\|\mathbf{H}_{t}\|\leq C$ for a constant $C>0$.
Assumption 1 (i) and (ii) are two versions of the condition on the model operator $\mathcal{M}_{t}$. If model $\mathcal{M}_{t}$ is linear, it automatically satisfies both (i) and (ii) of Assumption 1. If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite interval, that is $\|\mathcal{M}_{t}(\mathbf{x})\|\leq\ell_{0}(1+\|\mathbf{x}\|)$ for any $\mathbf{x}\in R^{p}$ for a positive constant $\ell_{0}$. However, strongly nonlinear systems may not satisfy the Lipschitz condition. In that case, we assume $\mathcal{M}_{t}$ satisfies condition (ii), which to some extent is similar to the tangent linear hypothesis (Thepaut and Courtier, 1991; Bouttier and Courtier, 1999) in the 4D-Var, where $\mathcal{M}_{t}$ is determined in the linear approximation $\mathcal{M}_{t}(\mathbf{x}_{t}+\mathbf{h})\approx\mathcal{M}_{t}(\mathbf{x}_{t})+\dot{M}_{t}\mathbf{h}$ in the vicinity of an arbitrary state $\mathbf{x}_{t}$ for a perturbation $\mathbf{h}$, and the operator $\dot{M}_{t}$ is called the differential or the first derivative or tangent linear function of $\mathcal{M}_{t}$ at point $\mathbf{x}_{t}$. Assumption 2 assumes the eigenvalues of the three basic covariance matrices $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$ and $\boldsymbol{\Sigma}_{0}^{a}$ are bounded away from zero and infinity. As presented in Lemma 3.1, under Assumption 2, the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ satisfies $\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\geq\epsilon_{0}>0$, which is the condition for the thresholding estimator for the high dimensional case in next section. With Assumptions 1 and 2, Lemma 3.1 shows that the eigenvalues of $\boldsymbol{\Sigma}_{t}^{f}$ are bounded from both ends, which assures $\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})\leq C$ for some positive constant $C$ and is also the condition for the bandable or the circular bandable covariance matrix classes. Assumption 3 assumes $\|\mathbf{H}_{t}\|\leq C$ for some positive constant $C$, which is a mild condition on the observation operator $\mathbf{H}_{t}$.
Lemma 3.1.
Suppose Assumption 2 holds, then the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ satisfies $\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\geq\epsilon_{0}>0$ for a positive $\epsilon_{0}$. If we further assume that Assumption 1 holds, then $0<\epsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{t}^{f})\leq 1/\epsilon_{0}$ for a positive $\epsilon_{0}$.
The following proposition provides a basic result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$ is an estimator of the Kalman gain matrix assuming $\mathbf{R}_{t}$ is known.
Proposition 3.1.
Suppose Assumption 2 holds and $\lambda_{\min}(\mathbf{R}_{t})\geq\varepsilon_{0}$ for a positive $\varepsilon_{0}$,
$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|\leq\lambda_{\min}^{-1}\left(\mathbf{R}_{t}\right)\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|\leq{\lambda_{\min}^{-1}\left(\mathbf{R}_{t}\right)}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|.$
Proposition 3.1 suggests that the upper bound of the estimation error of the Kalman gain matrix estimation $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$. Proposition 3.1 holds under weak conditions and actually is valid for both fixed and high dimensions. Note that for the fixed dimension data, the sample covariance matrix is consistent with the underlying population covariance at the $\sqrt{n}$ rate. As $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ is a sample covariance estimator, this means that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$.
Remark 1.
The last inequality is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$. Let us take a close look at $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$. For the extreme case of $\mathbf{R}_{t}=\mathbf{0}$, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be 00 if $\mathbf{H}_{t}$ is invertible. Intuitively, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be close to 00 if $\mathbf{R}_{t}$ is close to $\mathbf{0}$. Suppose $p=q=1$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=(1+\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}/\sigma_{\mathbf{R}_{t}}^{2})^{-1}<1$. This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ decreases with the increase of $\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}$ or decrease of $\sigma_{\mathbf{R}_{t}}^{2}$. For $p,q>1$, if $\mathbf{H}_{t}=\mathbf{I}$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=\|\mathbf{I}-\boldsymbol{\Sigma}_{t}^{f}(\boldsymbol{\Sigma}_{t}^{f}+\mathbf{R}_{t})^{-1}\|=\varepsilon_{0}^{2}/(1+\varepsilon_{0}^{2})$ if we assume $\mathbf{R}_{t}=\varepsilon_{0}^{2}\boldsymbol{\Sigma}_{t}^{f}$. This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ will decrease with the decrease of $\varepsilon_{0}^{2}$, the ratio of $\mathbf{R}_{t}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
In practice, the true observation error covariance matrix $\mathbf{R}_{t}$ may be unknown. Let $\hat{\mathbf{R}}_{t}$ be an estimator of $\mathbf{R}_{t}$ based on either knowledge or historical data. Corollary 3.1 below provides an upper bound of the estimation error of the $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ when both the forecast and observation error covariance matrices $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
Corollary 3.1.
Under the assumptions of Proposition 3.1 and assume $\|\boldsymbol{\Sigma}_{t}^{f}\|\leq C$ for some positive constant $C$ and that $\hat{\mathbf{R}}_{t}$ is invertible with probability approaching $1$, then
$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ $\leq\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\|\mathbf{K}_{t}\|}\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|$
$\leq{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\lambda_{\min}^{-1}(\mathbf{R}_{t})\left\|\mathbf{H}_{t}\right\|\|\boldsymbol{\Sigma}_{t}^{f}\|\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|,$
where the Kalman gain matrix estimator takes the form of $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{\hat{R}}_{t}\big{)}^{-1}$.
Compared with the upper bound in Proposition 3.1, there is an extra term related to the bound of the estimation error of $\mathbf{R}_{t}$. Most of the results established in this paper are for the case of known $\mathbf{R}_{t}$ as is commonly assumed in the literature of the EnKF. Results for estimated $\mathbf{R}_{t}$ can be attained in the fashion shown in Corollary 3.1.
Recall that $\hat{\mathbf{x}}_{t}^{a}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{a}$ and $\hat{\mathbf{x}}_{t}^{f}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f}$ are, respectively, the analysis and forecast states of the EnKF. Based on the result in Proposition 3.1, we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.6)-(2.7). We also provide an upper bound for the difference between the forecast and the analysis states of the EnKF and the true state $\mathbf{x}_{t}$.
We first study the accuracy of the oracle $\mathbf{x}_{t}^{a}$ and $\mathbf{x}_{t}^{f}$ as “estimates” of the underlying true state $\mathbf{x}_{t}$. With (2.2), (2.3), (2.5), and (2.6), we have that $\mathbb{E}\big{\{}(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\big{\}}=\mathbf{0}$$\operatorname{Var}\big{\{}(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\big{\}}=\boldsymbol{\Sigma}_{t}^{f}$, $\mathbb{E}\left\{(\mathbf{x}_{t}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\right\}=\mathbf{0}$$\operatorname{Var}\left\{(\mathbf{x}_{t}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\right\}=\boldsymbol{\Sigma}_{t}^{a}$. Thus, $\mathbb{E}\|(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})\,\text{and }\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}\left(\boldsymbol{\Sigma}_{t}^{a}\right)=\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}\leq\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})$, and the discrepancies between the EnKF forecast and analysis states in comparison to the true state are, respectively,
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}$ $=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}\,\text{ and}$ (3.1)
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ $=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}.$
As mentioned before, the analysis state $\mathbf{x}_{t}^{a}$ of the Kalman Filter achieves optimality under the linear Gaussian state-space model (2.1). In fact, as shown in the following Proposition 3.2, under the linear model operator $\mathcal{M}_{t}$, Gaussian initial state and fixed dimensional $p$ and $q$, the asymptotic distributions of the forecast and analysis ensembles in the EnKF are the same as those of the Kalman Filter. Both the analysis state $\hat{\mathbf{x}}_{t}^{a}$ and the analysis covariance $\hat{\boldsymbol{\Sigma}}_{t}^{a}$ of the EnKF converge to the counterparts of the Kalman Filter $\mathbf{x}_{t}^{a}$ and $\boldsymbol{\Sigma}_{t}^{a}$, respectively.
Proposition 3.2.
Suppose Assumptions 2 and 3 hold, the model operator $\mathcal{M}_{t}$ is linear, $\mathbf{x}_{0}\mid\mathbf{y}_{0}\sim{\mathrm{N}}_{p}\left(\mathbf{x}_{0}^{a},\boldsymbol{\Sigma}_{0}^{a}\right)$, from which the initial analysis ensemble $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ are independently generated, then, for fixed dimensions $p$ and $q$, the forecast state of the EnKF obtained by (2.9) converges conditionally in probability or in distribution such that $\hat{\mathbf{x}}_{t}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbf{x}_{t}^{f}$, $\hat{\boldsymbol{\Sigma}}_{t}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Sigma}_{t}^{f}$, $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathrm{N}_{p}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$, and the analysis state of the EnKF in (2.12) satisfies $\hat{\mathbf{x}}_{t}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbf{x}_{t}^{a}$, $\hat{\boldsymbol{\Sigma}}_{t}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Sigma}_{t}^{a}$, and $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathrm{N}_{p}(\mathbf{x}_{t}^{a},\boldsymbol{\Sigma}_{t}^{a})$ as $n\rightarrow\infty$, where $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{a}$ are the sample covariance matrices of the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{i=1}^{n}$ and the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{i=1}^{n}$, respectively.
Proposition 3.2 implies that with linear model operator $\mathcal{M}_{t}$ and Gaussian initial state, the asymptotic conditional distributions of the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{j=1}^{n}$ given $\mathbf{y}_{1:t}$ and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ given $\mathbf{y}_{1:t-1}$ in the EnKF coincide with the distributions of $\mathbf{x}_{t}|y_{1:t}$ and $\mathbf{x}_{t}|y_{1:t-1}$ in the Kalman Filter, respectively.
The following Theorem 3.1 provides the upper bounds of the mean square error (MSE) between the analysis and forecasting states of the EnKF in comparison to those of the Kalman Filters according to the true underlying evolution given in (2.6)-(2.7) for the one-step assimilation for a general case where the model operator $\mathcal{M}_{t}$ is nonlinear.
To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
$\Omega_{t}=$ $\operatorname{tr}\big{(}{n}^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({{\mathbf{H}}_{t}\boldsymbol{\Sigma}_{t}^{f}{\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t}})\big{)}+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f})\|^{2}$
$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}/n\big{)}+2\operatorname{tr}\big{(}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}\big{)}.$ (3.2)
Theorem 3.1.
Suppose Assumptions 1(i) or 1(ii), 2, 3 hold, and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$ with $\mathbf{x}_{t,j}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$ at time $t$. Then, the forecast and analysis states of the EnKF at time $t$ satisfy
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}$ $={n}^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}{n}^{-1}\big{)},$
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $=\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+\Omega_{t}=\operatorname{tr}\big{(}n^{-1}\boldsymbol{\Sigma}_{t}^{a}\big{)}+\Omega_{t}$ (3.3)
$=O\big{(}{n}^{-1}+\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}=O\big{(}{n}^{-1}\big{)},$
where the results reported in the $O$-terms are valid for the fixed dimension case while the other results are valid for both fixed and high dimensional cases.
Note that the first term in (3.3) is the variance of the sample mean of the analysis ensemble using the true forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the second term $\Omega_{t}$ results from the estimation error of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$. Theorem 3.1 implies that $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=O\left(n^{-1}\right)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ for the one-step data assimilation under the fixed dimensional case, which means both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$. These are because the sample covariance matrix $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ is consistent with the underlying population covariance at the $\sqrt{n}$ rate, and it follows that $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}=O(n^{-1})$. The proof of these results can be found in the proof of Theorem 3.1 in the SM. It should also be noted that the above results are valid for the model operators $\mathcal{M}_{t}$ given in Assumption 1-(i) and (ii).
The discrepancies between $\hat{\mathbf{x}}_{t}^{f}$ or $\hat{\mathbf{x}}_{t}^{a}$ and the true state $\mathbf{x}_{t}$ are given by (3.1). As $\mathbf{x}_{t}$ is random with $(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$ in model (2.8), $\mathbb{E}\|(\mathbf{x}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})$ and $\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ are constants not related to the ensemble size $n$. This gives the result $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=O({n}^{-1}+1)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}$, which is not related to the ensemble size $n$ in the leading order and can not converge to zero. This is due to the fact that the underlying state is a random variable. In contrast, $\mathbf{x}_{t}^{a}$ is a conditional mean and is possible to be estimated consistently.
The above justification on the validity of the Kalman Filter and EnKF algorithms is made for the fixed or low dimensional state vectors $\mathbf{x}_{t}$ where $p$ is of a smaller order of the ensemble size $n$ such that $p/n\to 0$. As the modern measurement technology on the earth system improves, the spatial resolution of the state variables becomes much higher. This makes $\mathbf{x}_{t}$ ultra high dimensional and creates high dimensional challenges on the validity of the EnKF, which is treated in the next section.
4 Convergence for High Dimensional Case
Remote sensing technology has been advancing the measurement of massive amounts of data sets for many geophysical processes, which produces high resolution data and brings high dimensional challenges. When the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.10) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ as shown in Bai and Yin (1993). The inconsistency of the forecast error covariance estimator (2.10) to $\boldsymbol{\Sigma}_{t}^{f}$ is a major threat to the validity of the EnKF method. Although there are methods such as the localization and inflation approaches designed to rescue the EnKF in the high dimensional case, the statistical properties of these approaches remain largely unknown. The last 15 years have seen rapid development of high dimensional statistical inference, including efforts to construct consistent estimators for high dimensional covariance matrices which have proposed the banding, the thresholding and the tapering estimators (Bickel and Levina, 2008a, b; Cai, Zhang and Zhou, 2010).
In this section, we first introduce the covariance classes where consistent estimation of high dimensional covariance matrices can be made, and present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$. Then, we provide results on the consistent estimation of the Kalman gain matrix $\mathbf{K}_{t}$ and evaluate the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ from the EnKF to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.
4.1 Consistent estimation of high dimensional forecast error covariance matrix
(a) $h(k)$ of $\boldsymbol{\Sigma}\in\mathcal{U}_{b}$
(b) $h(k)$ of $\boldsymbol{\Sigma}\in\mathcal{U}_{bc}$
(c) Heatmap of $\boldsymbol{\Sigma}\in\mathcal{U}_{b}$
(d) Heatmap of $\boldsymbol{\Sigma}\in\mathcal{U}_{bc}$
Figure 1: Bandable and circular bandable covariance matrices: the average energy function $h(k)$ in Panels (a) and (b) and their heatmaps in Panels (c) and (d), where $h(k)$ is the average of the squares of the $k^{th}$ sub-diagonal entries.
The covariance matrix class suitable for the banding estimator is the following bandable covariance class (Bickel and Levina, 2008a)
$\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}$ $\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0$
[…]
for some positive $\alpha$, C and $\varepsilon_{0}$. The bandable class prescribes a decaying pattern of $|\sigma_{ij}|$ as $|i-j|$ increases with the decaying rate $\alpha$. This bandable pattern is generally consistent with the dependence between variables observed at sites $i$ and $j$ in atmospheric and geoscience studies, as the dependence becomes weaker as the distance between the two sites gets larger. Note that $\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|$ is the $L_{1,1}$ norm of $B_{k}\left(\boldsymbol{\Sigma}\right)-\boldsymbol{\Sigma}$, which bounds the spectral norm from above (Golub and Van Loan, 2013).
In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or the poles of the earth, the state vectors are highly dependent for the variable components between the two ends of the state vector. This leads to a circular form of the bandable structure as follows:
$\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=$ $(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }$
$0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$
Let $h(k)$ be the average of the squares of the $k^{th}$ subdiagonal entries of a covariance:
$h(k):=\frac{1}{2(p-k)}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=\frac{1}{p-k}\sum_{l=1}^{p-k}\sigma_{ll+k}^{2}.$
Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariance matrices that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$. Figure 1 (a) and (b) show that the $h(k)$ function of the covariance matrix in $\mathcal{U}_{b}$ decays rapidly as the subdiagonal index $k$ increases, while the $h(k)$ function of the covariance matrix in $\mathcal{U}_{bc}$ first decays rapidly, then remains almost unchanged and rises again.
The bandable matrix class requires that the state vector $\mathbf{x}_{t}$ has an ordering (or a permutation) such that the correlation decays as two components of $\mathbf{x}_{t}$ are further apart. This may be argued for climate and geospatial applications, where components of the state vectors can be arranged via the geo-distance of the observational sites where the state vectors are associated with. However, the presence of geo-spatial processes (weather front or hurricane) and geography (mountains or ocean) will interrupt the spatial dependence ordering prescribed by the banding covariance class. In such situations, the following thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008b) may be employed. The thresholding covariance class is a class of covariance matrices which are invariant under permutations of components of the state vector $\mathbf{x}_{t}$, defined by
$\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,$ $\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i$
[…]
We now present the estimators of the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding based on the sample covariance $\mathbf{S}_{t}^{f}$ of the EnKF ensembles. For any matrix $M=\left[m_{ij}\right]_{p\times p}$, and any $0\leq k<p$, the banding matrix operator with a bandwidth $k$ is
$B_{k}(M)=\left[m_{ij}\mathbf{1}(|i-j|\leq k)\right].$
The banding estimator of the forecast error covariance matrix with bandwidth $k$ is $B_{k}(\mathbf{S}_{t}^{f})$.
The banding estimator administrates a hard threshold as it sets all matrix elements beyond the $k$-th sub-diagonals to zero. Cai, Zhang and Zhou (2010) updated the banding estimator with a tapering alteration
$\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)=\mathbf{S}_{t}^{f}\circ\mathbf{W}(k)=\big{(}w_{ij}(k)s_{tij}^{f}\big{)}_{p\times p},$
where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\left\{(k-|i-j|)_{+}-\left(k/2-|i-j|\right)_{+}\right\},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$. It is noted that the tapering estimator is a softer version of the banding estimator by tapering down the level of the thresholding gradually.
In order to provide valid estimation for forecast error covariance belonging to the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, we define a mid-banding operator with bandwidths $k_{1}$ and $k_{2}$,
$B_{k_{1},k_{2}}(M)=\left[m_{ij}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))\right],$
for any matrix $M=\left[m_{ij}\right]_{p\times p}$ and $0\leq k_{1}<(p-k_{2})\leq p$. Then the mid-banding forecast error covariance matrix estimator is $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$.
Recall that both the banding and tapering estimators require the variables in $\mathbf{x}_{t}$ to have a natural ordering under a permutation such that the correlation decays as two variables are further apart. For random vectors that cannot be ordered in such a way but belong to the thresholding covariance class $\mathcal{U}_{\tau}\left(q,c_{0}(p),M\right)$, Bickel and Levina (2008b) proposed the thresholding operator
$T_{s}(M)=\left[m_{ij}1\left(\left|m_{ij}\right|\geq s\right)\right],$
at a threshold level $s$. Then the thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})$.
Note that all of the banding, mid-banding, tapering and thresholding estimators depend on either the bandwidth $k$ or the threshold $s$. The selection of these parameters can be carried out by methods in Qiu and Chen (2015). For the bandwidths $k_{1},k_{2}$ of the mid-banding estimator, we propose a similar selection method and details can be found in the SM.
To apply the high dimensional estimators of the forecasting error covariance $\boldsymbol{\Sigma}_{t}^{f}$, we need the following distributional assumption.
Assumption 4.
The random vectors $\mathbf{x}_{t,i}^{f}=(x_{t,i1}^{f},\cdots,x_{t,ip}^{f})^{\mathrm{T}}$ are IID and each $x_{t,1j}^{f}\stackrel{{\scriptstyle d}}{{\sim}}F_{t,j}$ and $(x_{t,1j}^{f})^{2}\stackrel{{\scriptstyle d}}{{\sim}}G_{t,j}$, where $G_{t,j}(u)=F_{t,j}(\sqrt{u})-F_{t,j}(-\sqrt{u})$ for $u>0$. There exists a $\lambda_{0}>0$ such that $\max_{1\leq j\leq p}\mathbb{E}(e^{\lambda(x_{t,1j}^{f})^{2}})<\infty\textrm{ for }\lambda\in(-\lambda_{0},\lambda_{0}).$
Assumption 4 effectively assume $(x_{t,1j}^{f})^{2}$ are sub-exponentially distributed, which include the Sub-Gaussian random variables as a special case. The following Proposition 4.1 summarizes the convergence rates of the forecast error covariance matrix estimators by the banding, mid-banding, tapering and thresholding methods.
Proposition 4.1.
Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally IID Gaussian or sub-exponentially distributed as in Assumption 4.
  • (i)
    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$,
    $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
  • (ii)
    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$,
    $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
  • (iii)
    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$, suppose $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are sub-Gaussian, then for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$ the optimal rate of convergence of the tapering estimator is
    $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
  • (iv)
    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$ and $n^{-1}{\log p}=o(1)$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$,
    $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}.$
The results in Proposition 4.1 (i), (iii) and (iv) are, respectively, from Bickel and Levina (2008a), Cai, Zhang and Zhou (2010) and Bickel and Levina (2008b). And that for the mid-banding estimator in Proposition 4.1 (ii) is a new result whose proof is given in the SM. They imply that the high dimensional covariance estimators are consistent to $\boldsymbol{\Sigma}_{t}^{f}$ under the spectral norm if $\log(p)=o(n)$ or $p=o(e^{cn})$ for any positive constant $c$.
According to Proposition 3.1, the convergence results in Proposition 4.1 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$. Specifically, if $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$ or $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$,
$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}\,\text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k}(\mathbf{S}_{t}^{f})\text{ or }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f});\,$ (4.1)
[…]
$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}\text{ for }\,\hat{\boldsymbol{\Sigma}}_{t}^{f}=\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k);$ (4.2)
[…]
$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}\ \text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=T_{s}(\mathbf{S}_{t}^{f}).$ (4.3)
These suggest that under the conditions of Proposition 3.1, $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ has the same convergence rate as $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$ for any of the four high dimensional estimators.
As will be shown in the following, the estimation error of the Kalman gain matrix $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ plays a crucial role in quantifying the effects of the high dimensional  banding, the mid-banding, the tapering and the thresholding estimators for $\boldsymbol{\Sigma}_{t}^{f}$.
4.2 Convergence of High Dimensional Ensemble Kalman Filter
When the dimension $p$ of the state vector is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 4.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$. Based on the consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman gain matrix $\mathbf{K}_{t}$, we propose the high dimensional EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$. It is noted that for nonlinear model operator, $\mathbf{x}_{t}^{a}$ is defined as (2.6)-(2.7) as indicated before (2.13).
Similar to Theorem 3.1, the following Theorem 4.1 presents the upper bounds of the discrepancy between the EnKF and the Kalman Filters for the one-step data assimilation.
Theorem 4.1.
Suppose Assumptions 1(i) or 1(ii), 2, 3 hold and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the conditional distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t,j}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$ at time $t$. Then, when $p,q\rightarrow\infty$, the forecast and analysis states of the EnKF at time $t$ satisfy
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}$ $=$ ${n}^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}n^{-1}\big{)},$
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $=$ $\operatorname{tr}\big{(}n^{-1}p^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+p^{-1}\Omega_{t}$
$=$ $O\big{(}n^{-1}+p^{-1}\Delta_{\mathbf{K},t}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}.$
Theorem 4.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (4.1)-(4.3). Specifically, with the banding or mid-banding estimators, the tapering estimator and the thresholding estimator, the one-step EnKF data assimilation MSE are, respectively,
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\},$ (4.5)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}q(n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p})\big{\}},$ (4.6)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}qc_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}\big{\}}.$ (4.7)
To show the impact of the dimensionality on the MSE, we take the banding and the mid-banding estimators as an example. If $q\asymp p$, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O((n^{-1}{\log p})^{2\alpha/(2(\alpha+1))})$, which implies that the MSE between the analysis states of the EnKF $\hat{\mathbf{x}}_{t}^{a}$ and the $\mathbf{x}_{t}^{a}$ mainly depends on the estimation error of the Kalman gain matrix. Commonly in high resolution data assimilation, the dimension of the observation vector $\mathbf{y}_{t}$ is of much smaller order than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$. Assume $q=p^{\beta}$ for a $\beta\in(0,1)$, then the one-step assimilation MSE can be written as
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}n^{-1}+n^{-1}p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\big{)}.$
If $\beta$ is large so that $p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\rightarrow\infty$, the estimation error of the Kalman gain matrix will dominate the MSE, that is, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$. If $\beta$ is small such that $p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\rightarrow 0$, then $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O(n^{-1})$. For the impact of the observation dimension $q$ on the MSE, a study reported in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease as the first term in (4.1) will decrease as $q$ increases.
5 Imperfect Models
[…]
Assumption 5.
[…]
Theorem 5.1 presents the bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter for the one-step assimilation when $\mathcal{M}_{t}$ is unknown. Denote
$V_{t}$ $=\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\Xi_{\xi_{t}}\big{)}+\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2}$
$+\operatorname{tr}\big{(}n^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({\mathbf{H}}_{t}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}}){\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t})\big{)}$
$+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}-{\mathbf{H}}_{t}\mu_{\xi_{t}})\|^{2}+2\operatorname{tr}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}$ (5.1)
$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\{\mathbf{H}_{t}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}})/n+(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}-{\mathbf{H}}_{t}\mu_{\xi_{t}})\mu_{\xi_{t}}^{\mathrm{T}}\}\big{)}.$
Theorem 5.1.
Suppose Assumptions 1(i) or 1(ii), 2, 3, 5 hold and the analysis ensemble $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t-1}^{a}$ and covariance $\boldsymbol{\Sigma}_{t-1}^{a}$ at time $t-1$. Then, the forecast and analysis states of the EnKF satisfy
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=n^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}})+p^{-1}\|\mu_{\xi_{t}}\|^{2},$ (5.2)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=p^{-1}\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+V_{t}$ (5.3)
$=O\Big{(}n^{-1}+p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2}+p^{-1}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{(}\|\Xi_{\xi_{t}}\|^{2}+\tilde{\Delta}_{\mathbf{K},t}\big{)}\Big{)}$
where $\tilde{\Delta}_{\mathbf{K},t}=\mathbb{E}\|\tilde{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ and $\tilde{\mathbf{K}}_{t}$ is the estimate of $\mathbf{K}_{t}$ with true model $\mathcal{M}_{t}$.
[…]
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}n^{-1}+\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}(\|\Xi_{\xi_{t}}\|^{2}+n^{-1})\big{)}=O\big{(}n^{-1}+\|\Xi_{\xi_{t}}\|^{2})\big{)}.$ (5.4)
Thus, if $\|\Xi_{\xi_{t}}\|^{2}=O(n^{-1})$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=O\left(n^{-1}\right)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$, which means that the model error does not alter the order of magnitude of the assimilation and forecasting MSEs of the EnKF. However, if $n\|\Xi_{\xi_{t}}\|^{2}\rightarrow\infty$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O(\|\Xi_{\xi_{t}}\|^{2})$, which means the modelling error dominates the upper bounds of the MSE in the analysis states of the EnKF. When the dimensions $p$ and $q\rightarrow\infty$,
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}\asymp\max\{n^{-1},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$ (5.5)
[…]
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}\text{ and }$ (5.6)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
The above analysis shows that for both the fixed and the diverging dimension, and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the MSE of the assimilated $\hat{\mathbf{x}}_{t}^{a}$ by the EnKF, which highlights the benefits of having more accurate model operators. While this may be obvious, Theorem 5.1 describes exactly how it happens. In summary, the approximation error of the model operator $\mathcal{M}_{t}$ has a strong influence on the MSE of the data assimilation of the EnKF. This suggests that a good approximation of the model operator is a significant aspect of data assimilation. We also introduce an iterative EnKF update algorithm in S.1.2 to account for the model approximation error to some extent since most geophysical models are imperfect in practice.
6 Simulation
This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed high dimensional EnKF algorithms. These two models are often used for testing new approaches in atmospheric sciences for both data assimilation and the spatial-temporal forecasts. The proposed high dimensional EnKF methods are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models. Details of the proposed high dimensional EnKF algorithms are given in Section S.1 of the SM, including the consistent EnKF update in S.1.1 and the iterative EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators.
6.1 Lorenz-96 model
Lorenz-96 model (L96) (Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere. It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation
${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),$ (6.1)
for $j=1,\ldots,p$, with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denote the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $-\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is the force term. The $F$ can be set as other constants to generate imperfect models, the context discussed in Section 5 .
L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016), which approximates the model as $\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)=\mathbf{x}_{t}+h(k_{1}+2k_{2}+2k_{3}+k_{4})/{6},$ where $k_{1}=\mathbf{f}\left(\mathbf{x}_{t},t\right)$, $k_{2}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{1}/2,t+h/2\right)$, $k_{3}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{2}/2,t+h/2\right)$, $k_{4}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{3},t+h\right)$ and $\mathbf{f}=\left(f_{1},f_{2},\dots,f_{p}\right)^{\mathrm{T}}$. An approximation to the true state was simulated by either (i) implementing the RK4 method with a finer time step, say $h=0.05$, which corresponds to $6$ hours in the real-time assuming that the time-scale of the dissipation in the atmosphere is $5$ days (Lorenz, 1996); or (ii) by adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$. In our experiment, $\mathbf{w}_{t}\sim\mathrm{N}(0,\sigma_{0}\mathbf{I}_{p})$ and $\sigma_{0}$ were set as $0.1,0.25$ and $0.5$, respectively.
Furthermore, we experimented $p=40$ and $100$, under which the system behaves chaotically, since the L96 model behaves like chaos for $p\geq 7$ (Bedrossian, Blumenthal and Punshon-Smith, 2022). The true forcing was $F=8$. Similar to Liang et al. (2012), we also considered the misspecified model setting by making $F=6,7,9,10$ in the assimilation processes, corresponding to the imperfect models in Section 5. The initial state $\mathbf{x}_{1}$ was such that $x_{1}^{(j)}=F$ for $j\neq 20$ and $x_{1}^{(20)}=F+0.001$, and the initial forecast ensemble was generated by $\mathbf{x}_{1,j}^{f}=\mathbf{x}_{1}+\mathrm{N}_{p}(0,0.1\mathbf{I}_{p})$ for $j=1,\ldots,n$.
The observations $\mathbf{y}_{t}$ were made at $q$ randomly selected components of the state vector $\mathbf{x}_{t}$ with $q=30,40$ and $100$ while $p=40$ and $100$, respectively, followed by adding observation errors $\varepsilon_{t}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{q}(0,\mathbf{R}_{t})$ to the true states. The $(i,j)-$element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,p-|i-j|\right\}}$, which prescribes spatially correlated observation errors to suit geophysical data. The observation operator $\mathbf{H}_{t}$ was a binary $q\times p$ matrix with the $(j,i)$ element being $1$ if and only if the $j-$th component of $\mathbf{y}_{t}$ was the observed $i-$th dimension of the state $\mathbf{x}_{t}$. We simulated observations every $4$ time steps for a total of $2000$ steps and reported the results for the last $1000$ steps to allow the system to settle down. The ensemble size $n$ was $20,30,60$ and $90$. All the simulation experiments were repeated 500 times.
We considered the high dimensional EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance to estimate the forecast error covariance without the proposed high dimensional treatment and the inflation methods (Liang et al., 2012). The performances of the assimilation methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$, averaged over the last 1000 time steps.
However, the underlying analysis states $\mathbf{x}_{t}^{a}$ were unknown in the simulation of L96 due to its nonlinear nature. Indeed, for a nonlinear $\mathcal{M}_{t}$, the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ could not be readily updated by (2.3) in the Kalman Filter. To gain an accurate estimate of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance matrices to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman filter update formula to obtain $\mathbf{x}_{t}^{a}$, since the large ensemble size would make the sample covariance a high quality substitute for $\boldsymbol{\Sigma}_{t}^{f}$.
Figure 2: Average RMSEs to the oracle analysis state and their $25-75\%$ quantile bands (color bands) as a function of forcing $F$ used in ${\cal{M}}_{t}$ (true value is 8) by the standard EnKF (blue), the modified EnKF methods: the inflation (red), banding (green), tapering (cyan) and thresholding (purple) for $p=q=40$ and $100$, $n=20,30,60,90$, $\sigma_{0}=0.1,0.25$ and $0.5$, respectively. The light-colored bars in the lower part of each panel denote the divergence rates of the assimilation schemes.
Figure 2 displays the average RMSEs to the oracle analysis state of the five assimilation methods, namely the standard EnKF, the inflation method (Liang et al., 2012) and the three high dimensional EnKF equipped with the banding and tapering (both respect the circular bandable structures), and the thresholding estimators, with respect to different forcing level $F$ used in the model at moderate dimension $p=40$ and high dimension $p=100$ under different $\sigma_{0}$ and $n$. It shows that the EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) model settings. Even when $F$ became more deviated from the underlying value $8$, the analysis states obtained by the high dimensional EnKFs were still much closer to the oracle analysis states than the inflation method although both strains of methods showed increased RMSEs with respect to $|F-8|$. The standard EnKF with the sample covariance estimator to $\boldsymbol{\Sigma}_{t}^{f}$ had much larger assimilation error in all cases.
Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), we see that the advantages of the proposed high dimensional EnKF schemes over the inflation method were more apparent for the higher dimension setting $p=100$. The analysis RMSEs of the high dimensional EnKF methods kept at a lower level especially when the force $F$ was equal or close to the true value ($F=8$), while the analysis RMSE of the inflation method ranked behind but better than the standard EnKF. This means that the proposed high dimensional EnKF methods are quite attractive and useful as the geophysical models in earth science are of ultra high dimension. The figure also displays the percentages of divergence of the five methods, which shows that the inflation method endured quite high divergences, even higher than the standard EnKF. In contrast, the three high dimensional EnKF methods had quite low divergence rates for all cases. The divergence rate of the inflation method increases substantially along with the increase of the model error, which, for instance, reached 50% when $\sigma=0.5$.
Panels (d), (g)-(f) of Figure 2 demonstrate the influence of the ensemble size on the analysis RMSEs, which show that the analysis RMSEs of the inflation and the high dimensional EnKF methods became smaller with the increase of the ensemble size $n$. The proposed high dimensional EnKF methods performed much better than the inflation method with smaller RMSEs, especially when the ensemble size $n$ was small (panels (d) and (g) with $n=20$). This suggests that the proposed high dimensional EnKF methods could provide more accurate analysis states when the ensemble size was not big enough. Ensemble size is a significant restriction for the running of large geophysical models as they are extremely costly to run.
Table 1: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes for Lorenz-96 model with $p=q=40,100$, $n=30$, $\sigma_{0}=0$ and $F=8$ for the correct and other $F$ values for the misspecified models. The numbers inside the parentheses are the $25\%$ and $75\%$ quantiles of the RMSEs.
p type method F=6 F=7 F=8 F=9 F=10
$40$ oracle standard 4.27 (4.22,4.33) 4.37 (4.30,4.44) 4.62 (4.54,4.69) 4.92 (4.85,4.99) 5.23 (5.17,5.30)
inflation 1.08 (1.06,1.10) 0.83 (0.81,0.84) 0.60 (0.58,0.62) 0.93 (0.91,0.94) 1.21 (1.18,1.23)
banding 0.88 (0.87,0.88) 0.75 (0.74,0.76) 0.69 (0.68,0.70) 0.79 (0.78,0.80) 0.96 (0.95,0.97)
tapering 0.88 (0.87,0.88) 0.75 (0.74,0.75) 0.68 (0.67,0.69) 0.80 (0.79,0.81) 0.97 (0.96,0.99)
threshold 0.91 (0.90,0.92) 0.78 (0.77,0.79) 0.51 (0.38,0.57) 0.82 (0.81,0.83) 0.98 (0.97,0.99)
true standard 4.31 (4.25,4.36) 4.39 (4.32,4.47) 4.62 (4.54,4.69) 4.91 (4.84,4.98) 5.22 (5.16,5.29)
inflation 1.18 (1.16,1.20) 0.88 (0.86,0.89) 0.59 (0.57,0.62) 0.96 (0.94,0.98) 1.27 (1.25,1.29)
banding 0.98 (0.97,0.99) 0.80 (0.79,0.81) 0.71 (0.70,0.72) 0.83 (0.82,0.84) 1.04 (1.03,1.06)
tapering 0.98 (0.97,0.99) 0.80 (0.79,0.81) 0.70 (0.69,0.71) 0.84 (0.83,0.85) 1.06 (1.05,1.07)
threshold 1.02 (1.01,1.03) 0.83 (0.82,0.84) 0.57 (0.45,0.62) 0.87 (0.86,0.88) 1.07 (1.06,1.08)
$100$ oracle standard 4.40 (4.37,4.44) 4.58 (4.54,4.62) 4.84 (4.80,4.88) 5.14 (5.10,5.18) 5.45 (5.41,5.49)
inflation 2.91 (2.88,2.95) 2.83 (2.79,2.88) 2.88 (2.83,2.92) 3.11 (3.05,3.15) 3.40 (3.36,3.44)
banding 1.15 (1.13,1.16) 0.74 (0.73,0.75) 0.50 (0.49,0.50) 0.89 (0.87,0.91) 1.68 (1.63,1.72)
tapering 1.27 (1.24,1.30) 0.75 (0.74,0.76) 0.46 (0.45,0.47) 1.08 (1.05,1.11) 2.56 (2.46,2.65)
threshold 1.72 (1.65,1.79) 1.04 (1.02,1.06) 0.83 (0.82,0.85) 1.15 (1.12,1.18) 1.89 (1.82,1.96)
true standard 4.43 (4.40,4.46) 4.59 (4.55,4.63) 4.84 (4.80,4.88) 5.13 (5.10,5.17) 5.44 (5.40,5.47)
inflation 2.94 (2.91,2.98) 2.85 (2.80,2.89) 2.87 (2.83,2.92) 3.10 (3.05,3.15) 3.41 (3.37,3.45)
banding 1.26 (1.25,1.28) 0.84 (0.83,0.85) 0.60 (0.59,0.60) 0.97 (0.95,0.99) 1.75 (1.71,1.79)
tapering 1.38 (1.35,1.41) 0.85 (0.84,0.86) 0.57 (0.56,0.58) 1.15 (1.12,1.18) 2.60 (2.51,2.69)
threshold 1.76 (1.70,1.84) 1.06 (1.04,1.08) 0.82 (0.80,0.84) 1.15 (1.12,1.19) 1.92 (1.85,1.99)
Table 1 reports the average analysis RMSEs to the oracle analysis state and true state and their $25\%$ and $75\%$ quantiles by the standard EnKF, the inflation and the high dimensional EnKF methods using the banding and tapering and the thresholding estimators when $\sigma_{0}=0$ which is the same setting as Liang et al. (2012). One can see that the RMSEs to the oracle analysis state and true state of the proposed high dimensional methods were much smaller than the standard EnKF and the inflation method under both the correct and misspecified models. In the case of a higher dimensional setting with $p=100$, the effectiveness of the proposed high-dimensional EnKF schemes became even more pronounced. In contrast, the inflation method demonstrated significantly higher analysis RMSEs, even when applied to the correct model ($F=8$). For instance, when $p=100,F=8$, the banding method achieved an RMSE of 0.5, while the inflation method resulted in a much higher RMSE of 2.88. Besides, the high dimensional EnKF with banding estimator performed the best with the smallest RMSEs in almost all considered settings. Consistent with the theoretical results, the RMSEs to the true state were largely higher than those to the oracle analysis state.
Table 2: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes and their divergence rates (in the parentheses) for Lorenz-96 model for $p=40$, $q=30$, $n=30$, $\sigma_{0}=0.1$, and $F=8$ for the correct and other $F$ values for the misspecified models.
method F=6 F=7 F=8 F=9 F=10
standard 4.26 (0.01) 4.43 (0.06) 4.69 (0.15) 5.05 (0.29) 5.41 (0.54)
inflation 1.32 (0.80) 0.96 (0.65) 0.77 (0.63) 1.03 (0.69) 1.38 (0.86)
oracle banding 1.10 (0.11) 0.90 (0.08) 0.83 (0.09) 0.99 (0.12) 1.27 (0.29)
tapering 1.08 (0.13) 0.87 (0.07) 0.79 (0.06) 0.99 (0.13) 1.31 (0.25)
threshold 1.12 (0.06) 0.90 (0.02) 0.84 (0.04) 1.00 (0.04) 1.35 (0.16)
standard 4.28 (0.01) 4.44 (0.06) 4.70 (0.15) 5.05 (0.29) 5.40 (0.54)
inflation 1.42 (0.80) 1.04 (0.65) 0.84 (0.63) 1.11 (0.69) 1.47 (0.86)
true banding 1.23 (0.11) 1.01 (0.08) 0.93 (0.09) 1.09 (0.12) 1.37 (0.29)
tapering 1.21 (0.13) 0.98 (0.07) 0.90 (0.06) 1.09 (0.13) 1.42 (0.25)
threshold 1.24 (0.06) 1.00 (0.02) 0.93 (0.04) 1.09 (0.04) 1.44 (0.16)
Table 2 displays the analysis RMSEs of different assimilation methods and their divergence rates (in the parentheses) with sparser observations $q=30$ while $p=40$, which illustrates the influence of the sparsity of the observations on the assimilation accuracy. We see that the analysis RMSEs of different assimilation methods were slightly higher than those under more observations $p=q=40$. The RMSEs of the high dimensional EnKF methods were smaller than those of the inflation method and the standard for imperfect models ($F\neq 8$) and were comparable to the inflation method for the correct model ($F=8$). However, it is noted that the inflation method had a higher divergence rate which was 63% even under the correct model ($F=8$) and much higher (86%) for misspecified models ($F=10$). In contrast, the proposed high dimensional EnKF methods had low divergence rates, showing their robustness to the sparsity of the observations.
6.2 Shallow Water Equation model
The second experiment was conducted on the Shallow Water Equation model (SWE). The SWE is a system of hyperbolic/parabolic partial differential equations governing fluid flow, which is widely used in river flow simulations, as well as in other settings where the water depth is significantly lower than the horizontal length scale of motion. The barotropic nonlinear SWE takes the following form (Lei, Stauffer and Deng, 2012), for $0\leq z_{1}\leq L,0\leq z_{2}\leq D,$
$\begin{gathered}\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial z_{1}}+v\frac{\partial u}{\partial z_{2}}-fv=-g\frac{\partial h}{\partial z_{1}}+k\nabla^{2}u,\\ \frac{\partial v}{\partial t}+u\frac{\partial v}{\partial z_{1}}+v\frac{\partial v}{\partial z_{2}}+fu=-g\frac{\partial h}{\partial z_{2}}+k\nabla^{2}v,\\ \frac{\partial h}{\partial t}+u\frac{\partial h}{\partial z_{1}}+v\frac{\partial h}{\partial z_{2}}=-h\left(\frac{\partial u}{\partial z_{1}}+\frac{\partial v}{\partial z_{2}}\right)+k\nabla^{2}h,\end{gathered}$ (6.2)
where the fluid velocity components $u$ in $z_{1}$ direction and $v$ in $z_{2}$ direction as well as the fluid depth $h$ are the model variables, $g$ is the gravity acceleration $9.8\mathrm{~{}m}\mathrm{~{}s}^{-2}$, $f$ is the Coriolis parameter defined as constant $10^{-4}\mathrm{~{}s}^{-1}$ using the $f$-plane approximation, $k$ is the diffusion coefficient which was specified as $5\times 10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$, and $L$ and $D$ are the domain limits which were set to $500\mathrm{~{}km}$ and $300\mathrm{~{}km}$, respectively. The SWE model was discretized with a uniform grid spacing of $10\mathrm{~{}km}$ in the $z_{1}$ and $z_{2}$ directions and was solved using the Lax-Wendroff method with a time step of $30$ seconds. Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ having dimension $p=3p^{\prime}=4650$. The initial height (depth) field was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~{}m},5.5\mathrm{~{}m}$ and $3.325\mathrm{~{}m}$, respectively. The initial velocity field was derived from the initial height field with the geostrophic relation.
The SWE was simulated for $48$ hours with a total of 5760 time steps at 30 second intervals after a half hour (60-steps) pre-running. The observations of $u,v$ and $h$ were made at $q^{\prime}=310$ grid points from $10$ randomly selected rows in $z_{1}$ direction every 3 hours, namely 360 time points. The whole observation vector was $\mathbf{y}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ with dimension $q=3q^{\prime}=930$, which means the observation data were quite sparse with $q=0.2p$. To generate the imperfect models, the diffusion coefficient $k$ was also considered as $10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$ in the assimilation process. The observation error variances were specified as $\sigma_{\mathrm{u}}^{2}=\sigma_{\mathrm{v}}^{2}=0.5,\sigma_{\mathrm{h}}^{2}=1.0$ for the three dimensions and the observation error covariance matrix was $\mathbf{R}_{t}=diag(\sigma_{\mathrm{u}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{v}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{h}}^{2}\mathbf{I}_{p^{\prime}})$. The ensemble size $n$ was 100, and the initial ensemble was generated by adding random noises from $\mathrm{N}_{p}(0,\mathbf{I}_{p})$ to the initial state $\mathbf{x}_{1}$.
Figure 3: RMSEs to the oracle analysis state (the left panels) and the true state (the right panels) at each assimilation time step by the standard, inflation, banding, tapering and threshold under the correct model ($k=5\times 10^{4}$) and imperfect model ($k=10^{4}$) with $n=100$, $p=4650$ and $q=930$ on the SWE model.
Table 3: Average RMSEs to the oracle analysis state and true state over the last 24 hours (2880 steps) of the simulation experiments and their $25\%$ and $75\%$ quantiles (in parentheses) by the five EnKF schemes for the SWE model under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models with $n=100$, and $p=4650$ and $q=930$.
method k = 50000 (correct model) k = 10000 (imperfect model)
oracle true oracle true
standard 0.61 (0.45, 0.76) 0.62 (0.47, 0.78) 1.16 (1.05, 1.26) 1.18 (1.07, 1.28)
inflation 0.34 (0.27, 0.40) 0.38 (0.31, 0.44) 0.75 (0.68, 0.82) 0.77 (0.70, 0.84)
banding 0.20 (0.17, 0.22) 0.26 (0.23, 0.28) 0.42 (0.40, 0.45) 0.46 (0.43, 0.48)
tapering 0.25 (0.23, 0.28) 0.30 (0.28, 0.33) 0.44 (0.41, 0.46) 0.47 (0.44, 0.50)
threshold 0.28 (0.26, 0.30) 0.33 (0.31, 0.35) 0.44 (0.42, 0.46) 0.48 (0.46, 0.50)
Figure 3 reports the RMSEs at each assimilation time step of the five EnKF methods ( the standard EnKF, the inflation and the three high dimensional EnKF) to the oracle analysis states $\mathbf{x}_{t}^{a}$ and the true states $\mathbf{x}_{t}$ under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models. From Figure 3, the RMSEs at each assimilation time step of the three proposed high dimensional EnKF schemes became lower and lower along the time for both the correct and misspecified models. While the RMSEs of the standard EnKF increased over time and the inflation method also saw a slight increase in the RMSE for the true model $k=5\times 10^{4}$. Figure 3 shows that the three proposed high dimensional EnKF methods outperformed the inflation method for both the correct and imperfect models with much smaller RMSEs after the system settled down. The three proposed EnKF schemes saw much larger decreases in assimilation RMSEs in comparison to the standard EnKF and the inflation methods and the differences became larger over time, especially under the misspecified model. Among the three high dimensional EnKF schemes, the one utilizing the banding estimators again showed superior performance as in the simulations of the L96 model.
Table 3 summarizes the average analysis RMSEs to the oracle analysis state and true state over the last 24 hours, together with their $25\%$ and $75\%$ quantiles, by the standard EnKF, the inflation and the three high dimensional EnKF methods for both the correct and misspecified SWE models with $p=4650,q=930$. The results of Table 3 reveal that the high dimensional EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands when compared to the standard EnKF and the inflation-based EnKF. These results were particularly evident in cases where the model was imperfect. Specifically, the proposed high dimensional EnKF schemes were able to reduce the RMSE to the oracle analysis states by 41% in comparison with the inflation method. This provides evidence to support the notion that the high dimensional EnKF methods could attain more accurate and robust assimilated outcomes for SWE models even when the observations were extremely sparse and the state variables were ultra-high dimensional. This also suggests that the proposed high dimensional EnKF methods are highly appealing, especially considering that geophysical models in earth science often possess an ultra-high dimension.
7 Sea Temperature Assimilation in Northwest Pacific
Ocean is a complex nonlinear system containing motions in various spatial and temporal scales. The multiscale dynamic processes of the ocean and its cross-scale and cross-sphere interactions determine the material and energy cycles of the ocean, which profoundly affect the ocean’s heat absorption and carbon storage capacity, and regulate the evolution of global climate and marine ecosystems.
In this section, we demonstrate how the proposed method is applied to sea temperature assimilation based on the oceanography model named Regional Oceanic Modeling System (ROMS) with satellite data. Satellites provide a valuable tool for environmental and ocean monitoring. While the satellites produce high-resolution sea surface temperature data, they can not provide ocean measurement beneath the ocean surface where the physical model ROMS plays a key role. ROMS is a three-dimensional, free-surface, terrain-following ocean general circulation model that solves the Reynolds-averaged Navier–Stokes equations using the hydrostatic and Boussinesq assumptions (Haidvogel et al., 2000; Shchepetkin and McWilliams, 2005; Haidvogel et al., 2008). Terrain-following vertical coordinates are employed which allow for greater vertical resolution in regions with complex bathymetry.
The data assimilation was conducted from January 1st, 2008 to March 31st, 2008 over a study area 25.5N-34.7N, 144.4E-154E. The spatial resolution of the ROMS was $0.56^{\circ}\times 0.49^{\circ}$ with $18\times 20$ horizontal grids and $18$ uneven vertical layers over the water column. These made the dimension $p$ for the state variable of sea temperature ($\mathbf{x}_{t}$) to be $6480=18\times 20\times 18$. To scale down the computation, we did not include the other ocean variables namely salinity, horizontal velocity and vertical velocity of ocean current as the assimilated variables, and let them be internally determined by ROMS at each time step with the sea temperature as the assimilated variables.
In our study, the integration step (finer time discretization step) for ROMS was 5 minutes and the assimilation time length was daily which was also the frequency of the observations. The observational data were the daily Optimum Interpolation Sea Surface Temperature (OISST) from satellites provided by National Oceanic and Atmospheric Administration (NOAA), which can be downloaded at https://www.ncei.noaa.gov/products/optimum-interpolation-sst. OISST is a long-term data records for Sea Surface Temperature (SST) that incorporates observations from satellites, ships and buoys into regular global grids, where ships and buoys observations are used to correct satellite SST biases. Descriptions for OISST data can be found in (Reynolds et al., 2007; Banzon and Reynolds, 2013). As the OISST only provides sea surface temperature, the observation dimension was $q=360$.
To run ROMS, we need to input the initial state of sea temperature $\mathbf{x}_{1}$, the boundary conditions and smoothed bathymetry data (sea bottom topography). The bathymetry data came from ETOPO2, which is a global map of the earth’s seafloor and land elevations based on satellite altimetry, shipboard echo-sounding, and other data sources and can be downloaded from https://www.ncei.noaa.gov/products/etopo-global-relief-model, as shown in Figure S.1 of the SM. The initial state ($\mathbf{x}_{1}$) was from the Global Ocean Reanalysis and Simulations (GLORYS) (https://data.marine.copernicus.eu/products) on January 1st. The boundary conditions of ROMS can be found in Marchesiello, McWilliams and Shchepetkin (2001).
As the ocean temperature, salinity and current fields have different spatial scales and resolution in the three dimension study region both in the ROMS and the physical ocean, we transform the study region to a scale down space to facilitate the bandwidth selection in the banding estimator. Specifically, the grid points in the study region are denoted by coordinates $\{(u,v,h),u=1,2,\cdots 18,v=1,2,\cdots 20,h=1,2,\cdots 18\}$ in the transformed space of unit rectangles. The banding estimator is performed in the standardized grids. Specifically, the temperatures on the physical grids are labelled as $\{x_{u,v,h},u=1,2,\cdots 18,v=1,2,\cdots 20,h=1,2,\cdots 18\}$ in the standardized grids and the whole state vector in the physical grids can be labelled by $\mathbf{x}=(x_{1,1,1},\cdots,x_{18,20,18})^{\mathrm{T}}$ with the subscript $t$ hidden. The banding estimator with bandwidth $k$ means that we keep the covariances between $x_{u,v,h}$ and $x_{u^{\prime},{v^{\prime}},h^{\prime}}$ on the standardized grids if they satisfy $\max\{|u^{\prime}-u|,|v^{\prime}-v|,|h^{\prime}-h|\}\leq k$.
Considering that ocean is comparatively stable over a short period and to reduce the computation cost, we selected the bandwidth $k$ in the banding estimator by minimizing the out-of-sample RMSE over the first month of the study period between assimilated temperatures and temperatures recorded by Argo in different depths for four candidate levels $k\in\{1,2,3,4\}$. It was found that the RMSE increased with $k$ and was the smallest at $k=1$, which led to the use of $k=1$ as the bandwidth. The inflation factor for the inflation method was similarly selected by minimizing the out-of-sample RMSE over the first month which was $1.1$.
We performed the high dimensional EnKF scheme using the banding estimator with an ensemble size of $n=30$ to obtain the assimilated three-dimensional sea temperature for the study period and region. We also compared the proposed method with the standard EnKF and the inflation method. The out-of-sample performances of these methods were assessed on data collected under the ocean Argo project over the study period in the study region, which can be accessed at https://argo.ucsd.edu/data/. The Argo project collects information from inside the ocean using floats that initially drift with the ocean currents and then surface up and down between the surface and a mid-sea to collect measurements. As the observations provided by Argo are in-situ observations at various depths, they are more precise than the remote sensing data by the satellites, and hence, are used as substitutes for the ground truth.
Figure 4 displays both the in-sample RMSEs of the assimilated sea surface temperature against the OISST observations, and the out-of-sample RMSEs of the assimilated sea water temperature against the temperatures recorded by Argos at various depths. It also presents the correlations (Corrs) between the assimilated sea surface temperatures and those from OISST, and the Corrs between the assimilated temperatures with those recorded by Argo floats. Figure 4 shows that the assimilation RMSEs to the OISST data of the high dimensional EnKF kept stably at a much lower level during the whole assimilation period compared with those for the standard EnKF and the inflation methods. At the same time, the correlations between the OISST and the assimilated sea surface temperatures by the high dimensional EnKF were much higher and more stable than the two other methods which fluctuated quite substantially over the study period. On the out-of-sample Argo data, the high dimensional EnKF with the banding method produced assimilated sea temperatures which were closer to Argo data as the assimilation RMSEs were slightly smaller and the correlations to Argo data were slightly higher for most days during the assimilation period. These implies that the proposed high dimensional EnKF algorithm could achieve more accurate and robust assimilation results in practice.
Table 4 summarizes the average assimilation RMSEs and the correlations of the assimilated sea temperatures $\hat{\mathbf{x}}_{t}^{a}$ to the OISST data and the Argo data over the whole assimilation period by different methods together with their standard errors (in the parentheses). Obviously, the proposed high dimensional EnKF with banding estimator outperformed the standard EnKF and the inflation methods with much smaller RMSEs and much higher correlations to both the OISST and the Argo data. The reductions in the RMSEs and the increases in the correlations against both the OISST and the Argo data of the proposed high dimensional EnKF algorithm were significant compared with the standard EnKF and the inflation methods and they were much more apparent against the OISST data, suggesting that the high dimensional EnKF could make better use of the observation information. Besides, the standard errors of the RMSEs and correlations of the high dimensional EnKF were also smaller compared to the standard EnKF and the inflation methods, which was apparent from Figure 4. This implies that high dimensional EnKF provided more robust assimilation results.
Figure 4: Assimilation RMSEs and correlations to the OISST data (left panels) and the Argo data (right panels) from Jan. 1st 2008 to Mar. 31st 2008 using the standard EnKF (blue), the modified EnKF methods: inflation (red) and banding (green), respectively. The dashed line represents the average assimilation RMSEs and correlations to the OISST data and the Argo data over the whole assimilation period.
Table 4: The average assimilation RMSEs and correlations and their standard errors (in the parentheses) to the OISST data and the Argo data over the whole assimilation period.
Method OISST Argo
RMSE Corr RMSE Corr
standard 0.677(0.0067) 0.943(0.0012) 1.242(0.0421) 0.984(0.0013)
inflation 0.647(0.0071) 0.948(0.0011) 1.244(0.0421) 0.983(0.0012)
banding 0.277(0.0043) 0.990(0.0003) 1.134(0.0425) 0.986(0.0014)
Figure 5: The spatial distribution of the assimilated sea surface temperature on January 1st, February 15th and March 31st in 2008 over the study area, with the fill color from red to blue representing sea surface temperature from high to low.
Figure 5 depicts the spatial distribution of the assimilated sea surface temperature over the study area on January 1st, February 15th and March 31st in 2008. From the picture, we can see the spatial distribution of the assimilated sea surface temperature varied over time and that there was an overall downward trend in the study region during the study period.
8 Discussion
In this paper, we propose high dimensional EnKF algorithms to suit the needs of high resolution data assimilation. We take the opportunity to study the theoretical properties of the EnKF under both the fixed and the high dimensional Gaussian state-space model (2.8). The proposed high dimensional EnKF methods and algorithms can provide consistent data assimilation results in the high dimensional setting. The investigation of the high dimensional EnKF when the physical model $\mathcal{M}_{t}$ is misspecified gives a description of the performance of the EnKF under misspecified models.
The study shows that the proposed high dimensional EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations. One is that the established theoretical analysis on the consistency and the error bounds makes the proposed high dimensional methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies and the empirical study on sea temperature assimilation. The latter means that a smaller ensemble size for the EnKF can be used at a given level of accuracy which is particularly attractive as the running of the geophysical models is generally very expensive. The proposed high dimensional EnKF offers significant potential for application in earth science where the geophysical models, the state variables or the observations are ultra-high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.
{acks}
[Acknowledgments] The authors would like to thank Professor Xiaogu Zheng for insightful discussions and comments.
{funding}
The research was partially supported by a National Natural Science Foundation of China grant Nos. 12292980 and 12292983. This research was also partially supported by the National Key Scientific and Technological Infrastructure project “Earth System Science Numerical Simulator Facility” (EarthLab).
[…]

### v2

@@
-High Dimensional Ensemble Kalman Filter
+aoas/sm_enkf
 Abstract
 The ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of earth science, engineering and beyond.
-However, there are several unknown theoretical aspects of the EnKF, especially when the state variable is of high dimensional accompanied with high resolution observation and physical models.
-This paper first proposes several high dimensional EnKF methods which provide consistent estimators for the important forecast error covariance and the Kalman gain matrix.
-It then studies the theoretical properties of the EnKF under both the fixed and high dimensional state variables, which provides the mean square errors of the analysis states to the underlying oracle states offered by the Kalman filter and gives the much needed insight into the roles played by forecast error covariance on the accuracy of the EnKF.
+However, several theoretical aspects of the EnKF remain unknown, especially when the state variable is of high dimensional accompanied with high resolution observation and physical models.
+This paper first proposes several high dimensional EnKF (HD-EnKF) methods that provide consistent estimators for the important forecast error covariance and the Kalman gain matrix.
+It then studies the theoretical properties of the EnKF under both the fixed and high dimensional state variables, which provides the mean square errors of the analysis states to the underlying oracle states offered by the Kalman filter and gives the much needed insight into the roles played by the forecast error covariance on the accuracy of the EnKF.
 The accuracy of the data assimilation under the misspecified physical model is also considered.
-Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed high dimensional EnKF algorithms perform better than the standard EnKF methods as they provide more robust and accurate assimilated results.
-The high dimensional EnKF is applied to assimilate sea temperature in Northwest Pacific, which showed more accurate out-sample performance than the existing methods.
-Data Assimilation,
-Ensemble Kalman Filter,
-High Dimension Imputation,
-keywords:
-[…]
-[…]
-, and
+Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed HD-EnKF algorithms perform better than the standard EnKF methods as they provide more robust and accurate assimilated results.
+The HD-EnKF was applied to assimilate sea temperature in the Northwest Pacific, which showed more accurate out-sample performances than the existing methods.
+Keywords: Data Assimilation; Dynamic System; Ensemble Kalman Filter; High Dimensional Imputation.
 1 Introduction
 Kalman Filter (Kalman, 1960) is a fundamental approach for forecasting and data assimilation for a partially observed dynamic system.
-Kalman Filter and its later developments (Ensembled Kalman Filter, three and four–dimensional variational assimilation (3D-Var, 4D-Var) (Pailleux, 1990; Courtier et al., 1998) are based on the optimal estimation theory under the multivariate normal distribution.
-Data assimilation is a statistical procedure that revises a forecast based on the physical model to the state variables of the dynamic system upon receiving new observations.
-It combines information from numerical physical model forecasts with observed data to obtain the best possible description of a dynamical system and its uncertainty (Evensen, Vossepoel and van Leeuwen, 2022).
+Data assimilation is a statistical procedure that combines information from numerical physical model forecasts with observed data to obtain the best possible description of a dynamical system and its uncertainty (Evensen et al., 2022).
 The assimilated state variables at a time are used to generate the forecast for the next time step.
 For chaotic systems sensitive to initial values, data assimilation can significantly improve the accuracy of the forecast by using optimized state variables as initial values (Lorenz, 1963).
-Although much applied in numerical prediction for weather and oceanography, Kalman Filter and its later versions are the methodological core for scientific algorithms over a range of fields, including hydrology, remote sensing data utilization, air quality prediction, carbon data assimilation, robotic control and automatic driving.
+Although much applied in numerical weather and oceanography prediction, Kalman Filter and its later developments (Ensemble Kalman Filter, three and four-dimensional variational assimilation (3D-Var, 4D-Var; Pailleux, 1990; Courtier et al., 1998)) are the methodological core for scientific algorithms over a range of fields, including hydrology, remote sensing data utilization, carbon data assimilation, robotic control and automatic driving.
 Data assimilation methods also provide the underlying algorithms and predictive techniques for building digital twins for social and economic systems.
-The limitations of the Kalman Filter are (i) it assumes a linear model system which is known to be restrictive for many non-linear dynamic systems; and (ii) the forecast error covariance matrix and Kalman Gain matrix are expensive to store and update when the dimension of the state vector is large.
-The Ensemble Kalman Filter (EnKF; Evensen, 1994) was proposed to overcome these limitations, where the forecast error covariance matrix is estimated by the sample covariance based on generated forecast ensembles.
-The EnKF only requires storing and operating on reduced size state vectors and the ensembles can be generated parallelly.
-These advantages have made the EnKF an effective data assimilation scheme and has been widely used in atmospheric and oceanic sciences, and engineering.
-Estimating the forecast error covariance matrices is a key in the Kalman Filter and the EnKF schemes, which is made in the EnKF by the sample covariance matrix of the forecast ensembles.
-Past works on the EnKF found that the sampling error can lead to underestimation of the forecast error covariance and can eventually cause the filter to diverge (Anderson and Anderson, 1999; Constantinescu et al., 2007).
-The filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation.
-One reason for the filter divergence is the variance of the ensemble forecast error becomes too small when there are unaccounted model errors (Lorenc, 2003); the other is due to the substantial mismatch between the estimated and true error covariance matrices (Houtekamer and Mitchell, 1998; Constantinescu et al., 2007).
-Along with advances in remote sensing technology in measuring the earth system, high spatial resolution observations of the earth system are increasingly available, which leads to high dimensional geophysical models with high dimensional state vectors (Stroud et al., 2010; Guinehut et al., 2012; Lellouche et al., 2013).
-High resolution data assimilation and forecasting techniques are much needed in modern meteorology and atmospheric environment systems.
-For example, in global CO2 calculation, the dimension of the state variable can reach tens of millions.
+The limitations of the Kalman Filter are (i) it assumes a linear model which is restrictive for many non-linear dynamic systems; and (ii) the forecast error covariance and Kalman gain matrices are expensive to store and update when the dimension of the state vector is large.
+To overcome these limitations, Evensen (1994) proposed the Ensemble Kalman Filter(EnKF), where the forecast error covariance is estimated by the sample covariance based on generated forecast ensembles.
+The EnKF only requires storing and operating on much reduced-size state vectors and the ensembles can be generated in parallel.
+These advantages have made the EnKF an effective and widely used data assimilation scheme.
+Estimating the forecast error covariance is a key in the EnKF, which is made in the EnKF by the sample covariance of the forecast ensembles.
+Past works on the EnKF found that the sampling error can lead to underestimation of the forecast error covariance and can eventually cause filter divergence(Anderson and Anderson, 1999; Constantinescu et al., 2007).
+Filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation.
+High resolution data assimilation and forecasting techniques are much needed in modern meteorology and atmospheric systems like global carbon calculation and ocean data assimilation, which leads to high dimensional geophysical models with high dimensional state vectors.
+Besides, along with advances in remote sensing technology in measuring the earth system, there is a growing availability of high spatial resolution satellite observations (Stroud et al., 2010; Guinehut et al., 2012).
 The high dimensional state vectors lead to high dimensional  forecast error covariance and Kalman gain matrices.
-The high dimensionality is known to cause the non-convergence of the forecast error covariance matrix (Bai and Yin, 1993).
-In practice, the high dimensional effect is felt in the form of the filter divergence and unreliable data assimilation.
+The high dimensionality is known to cause the non-convergence of the sample forecast error covariance (Bai and Yin, 1993), which is felt in the form of filter divergence and unreliable data assimilation in practice.
 Geophysicists have proposed two methods to counter the problems.
-One is the localization method (Furrer, Genton and Nychka, 2006; Furrer and Bengtsson, 2007) that sets covariances to zero if the geographical distances between two sites are bigger than a threshold; and the other is the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term to avoid filter divergence.
-However, these methods lack support of statistical theory.
-The last two decades have witnessed rapid development in high dimensional statistical inference.
-Consistent high dimensional  covariance estimators have been proposed, which include the banding, tapering and thresholding estimators (Bickel and Levina, 2008a, b; Cai, Zhang and Zhou, 2010).
-Both the banding and tapering estimators require a bandwidth parameter, and the thresholding estimator needs the threshold level.
-Qiu and Chen (2015) proposed bandwidth and thresholding level selection methods by minimizing an empirical estimate of the expected squared Frobenius norms of the estimation error matrix.
-This paper proposes several high dimensional EnKF methods that utilize the high dimensional covariance estimators, and conducts a theoretical analysis of the proposed high dimensional EnKF.
-We also provide a theoretical analysis of the performance of the EnKF under the fixed dimensional setting since such a study appears lacking and is important.
-It is shown that with the use of the high dimensional covariance matrix estimators, consistent estimation of the Kalman Gain matrix and the underlying analysis states can be attained by the proposed high dimensional EnKF procedure.
-Specifically, we establish the mean square error (MSE) bounds of the analysis states from the EnKF methods to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models.
-Our analysis shows that consistent estimator of the forecast error covariance leads to the consistency of the Kalman Gain matrix by establishing an inequality that bounds the error of estimation of the Kalman Gain matrix by that of the forecast error covariance, which implies that, even at the high dimensional setting, the MSEs of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators.
-Thus, consistent high dimensional covariance estimators would make the high dimensional EnKF consistent under certain conditions.
-Indeed, our theoretical studies show that both the forecast and analysis states of the EnKF converge to their underlying counterparts of the Kalman Filter if the forecast error covariance can be estimated consistently.
-The investigation of the high dimensional EnKF under the misspecified model gives a more realistic description of the performance of the EnKF.
-Under the misspecified model, the MSE not only depends on the estimation error of the forecast error covariance matrices but also the discrepancy between the misspecified and the true model.
-Simulation studies on the Lorenz-96 model and the Shallow Water Equation model suggest that the proposed high dimensional EnKF algorithm could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method.
-The application to the sea temperature data assimilation in Northwest Pacific provides a comprehensive and accurate description of the spatiotemporal distribution of the three-dimensional sea temperature field by combining the ocean model with satellite data, which also illustrates the superiority and usefulness of the proposed high dimensional EnKF.
-This means the proposed high dimensional EnKF has wide application prospects in earth science like global carbon calculation, weather forecasting and the production of high-resolution data sets.
+One is the localization method (Furrer et al., 2006; Furrer and Bengtsson, 2007) that sets covariances to zero if the geographical distances are bigger than a threshold; and the other is the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term (Wang and Bishop, 2003; Liang et al., 2012; Constantinescu et al., 2007).
+However, the statistical properties of these EnKF methods are generally lacking, despite their widespread applications.
+Katzfuss et al.
+(2020) proposed Gibbs Ensemble Kalman Filter and Smoother, and particle Ensemble Kalman Filter and Smoother algorithms for high dimensional hierarchical dynamic models from a Bayesian perspective.
+This paper proposes several high dimensional EnKF (HD-EnKF) methods based on consistent high dimensional covariance estimators and conducts a theoretical analysis of assimilated state vectors.
+It is shown that consistent estimation of the forecast error covariance leads to the consistency of the Kalman gain matrix by establishing an inequality that bounds the error of estimation of the Kalman Gain matrix by that of the forecast error covariance.
+This implies that, even at the high dimensional setting, the mean square errors (MSEs) of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators.
+Thus, consistent recovery of the underlying analysis states can be attained by the proposed HD-EnKF procedure.
+We establish the mean square error bounds and weak convergence of the analysis states from the HD-EnKF to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models.
+The MSE bounds for the multiple-step forecasting and data assimilation are also provided.
+Simulation studies suggest that the proposed HD-EnKF algorithm could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method.
+We apply the proposed HD-EnKF to assimilate sea temperature by combining a physical ocean model with satellite observations and use the ARGO data to validate the quality of the assimilated sea temperature.
+Both the simulation and the empirical study showed advantages of the HD-EnKF approach.
 The paper is structured as follows.
 Section 2 reviews the Kalman Filter and the EnKF.
-Theoretical properties of the EnKF for the fixed dimensional case are given in Section 3.
-Section 4 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman gain matrix, and establishes the theoretical properties of the EnKF for high dimensional case.
+Section 3 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman gain matrix, and establishes the theoretical properties of the EnKF.
 The effects of imperfect or misspecified models on the accuracy of the EnKF are provided in Section 4.
-Results of empirical data assimilation based on two well known geophysical models are reported in Section 6 and the application to the sea temperature data assimilation is presented in Section 7 , followed by a conclusion in Section 8.
-Details of the proposed high dimensional EnKF algorithms, proofs of the theoretical results and additional results are given in the Supplementary Material (SM).
+Results of empirical data assimilation based on two well-known geophysical models are reported in Section 5 and the application to the sea temperature data assimilation is presented in Section 6, followed by a conclusion in Section 7.
+Details of the proposed HD-EnKF algorithms, additional theoretical and simulation results, and technical details are given in the Supplementary Material (SM).
 2 Review on Kalman and Ensemble Kalman Filters
@@
 $\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathbf{M}_{t}\mathbf{x}_{t-1}+\mathbf{w}_{t},\quad\ \quad\mathbf{w}_{t}\sim\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\,\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}(\mathbf{0},\mathbf{R}_{t}),\end{array}\right.$ (2.1)
-where $\mathbf{x}_{t}$ is the $p-$dimensional not fully observed state vector of a dynamic system at time $t$, $\mathbf{M}_{t}$ is the linear model operator (also called evolution or forecast operator), $\mathbf{w}_{t}$ is the error term that follows a zero-mean multivariate normal distribution with $\mathbf{Q}_{t}$ as the covariance, $\mathbf{y}_{t}$ is the $q-$dimensional observation vector at time $t$, $\mathbf{H}_{t}$ is a linear observation operator which relates the state $\mathbf{x}_{t}$ to the observation $\mathbf{y}_{t}$, and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix.
+where $\mathbf{x}_{t}$ is the $p-$dimensional not fully observed state vector of a dynamic system at time $t$, $\mathbf{M}_{t}$ is the linear model operator (also called evolution or forecast operator), $\mathbf{w}_{t}$ is the model error term with $\mathbf{Q}_{t}$ as its covariance, $\mathbf{y}_{t}$ is the $q-$dimensional observation vector at time $t$, $\mathbf{H}_{t}$ is a linear observation operator which relates the state $\mathbf{x}_{t}$ to the observation $\mathbf{y}_{t}$, and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix.
 Assuming $\mathbf{M}_{t}$, $\mathbf{Q}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are known, the Kalman Filter provides a two-step alternative updating framework based on the theory of multivariate normal distribution.
@@
 In the update (assimilation) step, the goal is to update $\mathbf{x}_{t}^{f}$ according to the new observation $\mathbf{y}_{t}$ and to reduce the difference between the true state $\mathbf{x}_{t}$ and the forecast state $\mathbf{x}_{t}^{f}$.
-The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is
-$\left(\begin{array}[]{c}\mathbf{x}_{t}\\ \mathbf{y}_{t}\end{array}\right)\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p+q}\left(\left(\begin{array}[]{c}\mathbf{x}_{t}^{f}\\ \mathbf{H}_{t}\mathbf{x}_{t}^{f}\end{array}\right),\left(\begin{array}[]{cc}\boldsymbol{\Sigma}_{t}^{f}&\boldsymbol{\Sigma}_{t}^{f}\mathbf{H}_{t}^{T}\\ \mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}&\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}\mathbf{H}_{t}^{T}+\mathbf{R}_{t}\end{array}\right)\right),$ (2.4)
-which implies that
-$\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right)$ (2.5)
+The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is Gaussian, which implies that
+$\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right),$ (2.4)
 where, by the conditional Gaussian distribution formulae,
@@
 which are, respectively, the analysis (assimilation) state and the associated analysis covariance matrix, and the Kalman gain matrix.
-It is noted that the analysis state $\mathbf{x}_{t}^{a}$ is the minimum variance unbiased estimator for the mean of $\mathbf{x}_{t}$ under the linear state-space model (2.1).
-However, models in earth science or engineering are often nonlinear, which implies that the evolution model in (2.1) may not be linear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$, where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2).
-Besides, for large dynamic systems, the dimension $p$ of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store.
-To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error variance matrix is no longer updated as (2.3) but by the sample covariance based on $n$ forecast ensembles.
-It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing, and thus improves the computing efficiency and reduces the storage cost.
+Note that the analysis state $\mathbf{x}_{t}^{a}$ is the minimum variance unbiased estimator for the mean of $\mathbf{x}_{t}$ under the linear state-space model (2.1).
+However, models in earth science or engineering are often nonlinear, that is the evolution model in (2.1) may not be linear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$, where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2).
+Besides, for large dynamic systems, the dimension $p$ of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store.
+To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error covariance is no longer updated as (2.3) but by the sample covariance based on $n$ forecast ensembles.
+It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing.
 The EnKF assumes possibly nonlinear model system with linear observation operator:
 $\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1}\right)+\mathbf{w}_{t},\quad\quad\ \mathbf{w}_{t}\sim\mathrm{N}_{p}(\mathbf{0},\mathbf{Q}_{t}),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\quad\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}\left(\mathbf{0},\mathbf{R}_{t}\right),\end{array}\right.$ (2.7)
-where $\mathcal{M}_{t}$ is the possibly nonlinear model operator, $\mathbf{H}_{t}$ is a linear observation operator and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix.
+where $\mathcal{M}_{t}$ is the possibly nonlinear model operator, $\mathbf{H}_{t}$ is a linear observation operator.
 It further assumes that the forecast error $\boldsymbol{\eta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$ follows a Gaussian distribution given $\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$, where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)$.
@@
   • Step (ii):
-    Estimate the forecast error covariance matrix $\Sigma_{t}^{f}$ by the sample covariance matrix of the forecast ensemble, namely
+    Estimate the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance matrix of the forecast ensemble, namely
     $\hat{\boldsymbol{\Sigma}}_{t}^{f}=(n-1)^{-1}\sum_{j=1}^{n}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}^{\mathrm{T}}\triangleq\mathbf{S}_{t}^{f},$ (2.9)
     and obtain an estimate of the Kalman gain matrix $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$.
-Here we note computationally efficient expressions
-    $\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}$ $=(n-1)^{-1}\sum_{j=1}^{n}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}^{\mathrm{T}}\,\text{and}$ (2.11)
-    $\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}$ $=(n-1)^{-1}\sum_{j=1}^{n}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}^{\mathrm{T}}$
-    since the dimension $q$ of the observation is much smaller than that ($p$) of the state vector.
   • Step (iii):
@@
     which will be used for the data assimilation at the next time step.
-In practice, the algorithm is initialized at time $t=0$ by drawing $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ independently from ${\mathrm{N}}_{p}\left(\mathbf{x}_{0}^{a},\Sigma_{0}^{a}\right)$ where $\mathbf{x}_{0}^{a}$ and $\Sigma_{0}^{a}$ are either chosen based on the physical model or by pre-running the physical model for a period of time.
-Then, the ensemble is propagated forward through time, alternating between the forecast and the analysis steps, according to the three-step procedure above.
-If $\mathbf{R}_{t}$ and $\mathbf{Q}_{t}$ are unknown, their estimates can be entertained to substitute them in (2.9) and (2.12).
 The EnKF does not require the expensive recursions (2.3) and (2.5)-(2.6) used in the Kalman Filter.
-Instead, it estimates the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance or the reduced dimension counterpart (2.11) via the observation operator $\mathbf{H}_{t}$.
+Instead, it estimates the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance.
 For the case with non-linear model operator as in (2.7), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.5)-(2.6), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance defined by $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$.
-In fact, $\mathbf{x}_{t}^{a}$ defined by (2.6)-(2.7) is the solution to the minimizing problem (2.13) with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$.
+In fact, $\mathbf{x}_{t}^{a}$ defined by (2.5)-(2.6) is the solution to the minimizing problem (2.11)
+$J(\mathbf{x})=\frac{1}{2}\big{\{}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\big{\}}$ (2.11)
+with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$.
 For the ease of reference, and to avoid differentiating between the linear and the non-linear cases, we also call $\mathbf{x}_{t}^{a}$ as the analysis state of the Kalman Filter even in the non-linear model case.
 The next two sections will establish the convergence of the EnKF analysis state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.
-It is noted that the update formula (2.12) can be obtained by minimizing
-$J(\mathbf{x})=\frac{1}{2}\left\{(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\right\}$ (2.13)
-with respect to $\mathbf{x}$. To appreciate this, note that by setting $\partial J(\mathbf{x})/{\partial\mathbf{x}}=\mathbf{0}$ and by the Sherman-Morrison-Woodbury formula, we get
-$\mathbf{x}$ =$=\left\{\hat{\boldsymbol{\Sigma}}_{t}^{f}-\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{R}_{t}+\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{)}^{-1}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\right\}\left\{(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}\mathbf{\hat{x}}_{t}^{f}+\mathbf{H}_{t}^{\mathrm{T}}\mathbf{R}_{t}^{-1}\mathbf{y}_{t}\right\}$
-$=\mathbf{\hat{x}}_{t}^{f}+\mathbf{\hat{K}}_{t}\big{(}\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}_{t}^{f}\big{)},$
-which is the data assimilation formula (2.12).
-The objective function (2.13) is related to that of the 3D-Var or 4D-Var, but not the same.
-However, over the same time interval, under the assumption of perfect model with the same inputs (initial mean and covariance matrix, observation error covariance), the 4D-Var analysis at the end is equal to the analysis of the Kalman filter; see Lorenc (2003), Hamill (2006) and Constantinescu et al.
-(2007) for details.
-3 Consistency for fixed dimension case
-We first provide the upper bound of the estimation error of the estimated Kalman gain matrix and then establish the convergence of the EnKF analysis states to the underlying Kalman Filter under both the linear and the non-linear model operators for the fixed dimensional state vectors with a convergence rate.
-Indeed, it appears that such results are lacking in the literature.
-Additionally, the error bounds could further be extended to the high dimension cases, which will be discussed later in Section 4.
+3 Convergence of Ensemble Kalman Filter
 […]
@@
 Assumption 2.
-The model error covariance matrix $\mathbf{Q}_{t}$, the observation error covariance matrix $\mathbf{R}_{t}$ and the initial analysis covariance matrix $\boldsymbol{\Sigma}_{0}^{a}$ satisfy $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{Q}_{t}\right)\leq\lambda_{\max}\left(\mathbf{Q}_{t}\right)\leq 1/\varepsilon_{0}$, $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{R}_{t}\right)\leq\lambda_{\max}\left(\mathbf{R}_{t}\right)\leq 1/\varepsilon_{0},$ and $\lambda_{\max}\left(\boldsymbol{\Sigma}_{0}^{a}\right)\leq 1/\varepsilon_{0},$ for a positive $\varepsilon_{0}$, respectively, where $\lambda_{\max}$ and $\lambda_{\min}$ are the maximum and minimum eigenvalue operators, respectively.
+The model error covariance $\mathbf{Q}_{t}$, the observation error covariance $\mathbf{R}_{t}$ and the initial analysis covariance $\boldsymbol{\Sigma}_{0}^{a}$ satisfy $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{Q}_{t}\right)\leq\lambda_{\max}\left(\mathbf{Q}_{t}\right)\leq 1/\varepsilon_{0}$, $0<\varepsilon_{0}\leq\lambda_{\min}\left(\mathbf{R}_{t}\right)\leq\lambda_{\max}\left(\mathbf{R}_{t}\right)\leq 1/\varepsilon_{0},$ and $\lambda_{\max}\left(\boldsymbol{\Sigma}_{0}^{a}\right)\leq 1/\varepsilon_{0},$ for a positive $\varepsilon_{0}$, respectively, where $\lambda_{\max}$ and $\lambda_{\min}$ are the maximum and minimum eigenvalue operators, respectively.
 Assumption 3.
@@
 If model $\mathcal{M}_{t}$ is linear, it automatically satisfies both (i) and (ii) of Assumption 1.
-If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite interval, that is $\|\mathcal{M}_{t}(\mathbf{x})\|\leq\ell_{0}(1+\|\mathbf{x}\|)$ for any $\mathbf{x}\in R^{p}$ for a positive constant $\ell_{0}$.
+If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite interval.
 However, strongly nonlinear systems may not satisfy the Lipschitz condition.
-In that case, we assume $\mathcal{M}_{t}$ satisfies condition (ii), which to some extent is similar to the tangent linear hypothesis (Thepaut and Courtier, 1991; Bouttier and Courtier, 1999) in the 4D-Var, where $\mathcal{M}_{t}$ is determined in the linear approximation $\mathcal{M}_{t}(\mathbf{x}_{t}+\mathbf{h})\approx\mathcal{M}_{t}(\mathbf{x}_{t})+\dot{M}_{t}\mathbf{h}$ in the vicinity of an arbitrary state $\mathbf{x}_{t}$ for a perturbation $\mathbf{h}$, and the operator $\dot{M}_{t}$ is called the differential or the first derivative or tangent linear function of $\mathcal{M}_{t}$ at point $\mathbf{x}_{t}$.
-Assumption 2 assumes the eigenvalues of the three basic covariance matrices $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$ and $\boldsymbol{\Sigma}_{0}^{a}$ are bounded away from zero and infinity.
-As presented in Lemma 3.1, under Assumption 2, the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ satisfies $\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\geq\epsilon_{0}>0$, which is the condition for the thresholding estimator for the high dimensional case in next section.
-With Assumptions 1 and 2, Lemma 3.1 shows that the eigenvalues of $\boldsymbol{\Sigma}_{t}^{f}$ are bounded from both ends, which assures $\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})\leq C$ for some positive constant $C$ and is also the condition for the bandable or the circular bandable covariance matrix classes.
-Assumption 3 assumes $\|\mathbf{H}_{t}\|\leq C$ for some positive constant $C$, which is a mild condition on the observation operator $\mathbf{H}_{t}$.
-Lemma 3.1.
-Suppose Assumption 2 holds, then the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ satisfies $\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\geq\epsilon_{0}>0$ for a positive $\epsilon_{0}$.
-If we further assume that Assumption 1 holds, then $0<\epsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{t}^{f})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{t}^{f})\leq 1/\epsilon_{0}$ for a positive $\epsilon_{0}$.
+In that case, we assume $\mathcal{M}_{t}$ satisfies condition (ii), which is related to the tangent linear hypothesis (Thepaut and Courtier, 1991; Bouttier and Courtier, 1999) in the 4D-Var, where $\mathcal{M}_{t}$ is determined in the linear approximation $\mathcal{M}_{t}(\mathbf{x}_{t}+\mathbf{h})\approx\mathcal{M}_{t}(\mathbf{x}_{t})+\dot{M}_{t}\mathbf{h}$ in the vicinity of $\mathbf{x}_{t}$ for a perturbation $\mathbf{h}$, and the operator $\dot{M}_{t}$ is called the differential or tangent linear function of $\mathcal{M}_{t}$ at point $\mathbf{x}_{t}$.
+Assumption 2 assumes the eigenvalues of the three basic covariances $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$ and $\boldsymbol{\Sigma}_{0}^{a}$ are bounded away from zero and infinity.
+From Lemma S.3.1 of the SM, the eigenvalues of $\boldsymbol{\Sigma}_{t}^{f}$ satisfy the conditions for the thresholding and bandable(or circular bandable) covariance classes under Assumption 2 and Assumptions 1-2, respectively.
+Assumption 3 puts a mild condition on the observation operator $\mathbf{H}_{t}$.
 The following proposition provides a basic result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$ is an estimator of the Kalman gain matrix assuming $\mathbf{R}_{t}$ is known.
@@
 $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|\leq\lambda_{\min}^{-1}\left(\mathbf{R}_{t}\right)\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|\leq{\lambda_{\min}^{-1}\left(\mathbf{R}_{t}\right)}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|.$
-Proposition 3.1 suggests that the upper bound of the estimation error of the Kalman gain matrix estimation $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$.
+The last inequality is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$.
+Proposition 3.1 suggests that the upper bound of the estimation error in $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$.
 Proposition 3.1 holds under weak conditions and actually is valid for both fixed and high dimensions.
-Note that for the fixed dimension data, the sample covariance matrix is consistent with the underlying population covariance at the $\sqrt{n}$ rate.
-As $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ is a sample covariance estimator, this means that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$.
-Remark 1.
-The last inequality is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$.
-Let us take a close look at $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$.
-For the extreme case of $\mathbf{R}_{t}=\mathbf{0}$, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be 00 if $\mathbf{H}_{t}$ is invertible.
-Intuitively, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be close to 00 if $\mathbf{R}_{t}$ is close to $\mathbf{0}$.
-Suppose $p=q=1$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=(1+\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}/\sigma_{\mathbf{R}_{t}}^{2})^{-1}<1$.
-This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ decreases with the increase of $\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}$ or decrease of $\sigma_{\mathbf{R}_{t}}^{2}$.
-For $p,q>1$, if $\mathbf{H}_{t}=\mathbf{I}$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=\|\mathbf{I}-\boldsymbol{\Sigma}_{t}^{f}(\boldsymbol{\Sigma}_{t}^{f}+\mathbf{R}_{t})^{-1}\|=\varepsilon_{0}^{2}/(1+\varepsilon_{0}^{2})$ if we assume $\mathbf{R}_{t}=\varepsilon_{0}^{2}\boldsymbol{\Sigma}_{t}^{f}$.
-This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ will decrease with the decrease of $\varepsilon_{0}^{2}$, the ratio of $\mathbf{R}_{t}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
+Note that under fixed dimension, the sample covariance is consistent with the underlying population covariance at the $\sqrt{n}$ rate, that is $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$.
+The results for the high dimensions are shown later.
 In practice, the true observation error covariance matrix $\mathbf{R}_{t}$ may be unknown.
-Let $\hat{\mathbf{R}}_{t}$ be an estimator of $\mathbf{R}_{t}$ based on either knowledge or historical data.
-Corollary 3.1 below provides an upper bound of the estimation error of the $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ when both the forecast and observation error covariance matrices $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
-Corollary 3.1.
-Under the assumptions of Proposition 3.1 and assume $\|\boldsymbol{\Sigma}_{t}^{f}\|\leq C$ for some positive constant $C$ and that $\hat{\mathbf{R}}_{t}$ is invertible with probability approaching $1$, then
-$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ $\leq\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\|\mathbf{K}_{t}\|}\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|$
-$\leq{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\lambda_{\min}^{-1}(\mathbf{R}_{t})\left\|\mathbf{H}_{t}\right\|\|\boldsymbol{\Sigma}_{t}^{f}\|\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|,$
-where the Kalman gain matrix estimator takes the form of $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{\hat{R}}_{t}\big{)}^{-1}$.
-Compared with the upper bound in Proposition 3.1, there is an extra term related to the bound of the estimation error of $\mathbf{R}_{t}$.
-Most of the results established in this paper are for the case of known $\mathbf{R}_{t}$ as is commonly assumed in the literature of the EnKF.
-Results for estimated $\mathbf{R}_{t}$ can be attained in the fashion shown in Corollary 3.1.
-Recall that $\hat{\mathbf{x}}_{t}^{a}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{a}$ and $\hat{\mathbf{x}}_{t}^{f}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f}$ are, respectively, the analysis and forecast states of the EnKF.
-Based on the result in Proposition 3.1, we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.6)-(2.7).
-We also provide an upper bound for the difference between the forecast and the analysis states of the EnKF and the true state $\mathbf{x}_{t}$.
-We first study the accuracy of the oracle $\mathbf{x}_{t}^{a}$ and $\mathbf{x}_{t}^{f}$ as “estimates” of the underlying true state $\mathbf{x}_{t}$.
-With (2.2), (2.3), (2.5), and (2.6), we have that $\mathbb{E}\big{\{}(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\big{\}}=\mathbf{0}$$\operatorname{Var}\big{\{}(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\big{\}}=\boldsymbol{\Sigma}_{t}^{f}$, $\mathbb{E}\left\{(\mathbf{x}_{t}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\right\}=\mathbf{0}$$\operatorname{Var}\left\{(\mathbf{x}_{t}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\right\}=\boldsymbol{\Sigma}_{t}^{a}$.
-Thus, $\mathbb{E}\|(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})\,\text{and }\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}\left(\boldsymbol{\Sigma}_{t}^{a}\right)=\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}\leq\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})$, and the discrepancies between the EnKF forecast and analysis states in comparison to the true state are, respectively,
-$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}$ $=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}\,\text{ and}$ (3.1)
-$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ $=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}.$
-As mentioned before, the analysis state $\mathbf{x}_{t}^{a}$ of the Kalman Filter achieves optimality under the linear Gaussian state-space model (2.1).
-In fact, as shown in the following Proposition 3.2, under the linear model operator $\mathcal{M}_{t}$, Gaussian initial state and fixed dimensional $p$ and $q$, the asymptotic distributions of the forecast and analysis ensembles in the EnKF are the same as those of the Kalman Filter.
-Both the analysis state $\hat{\mathbf{x}}_{t}^{a}$ and the analysis covariance $\hat{\boldsymbol{\Sigma}}_{t}^{a}$ of the EnKF converge to the counterparts of the Kalman Filter $\mathbf{x}_{t}^{a}$ and $\boldsymbol{\Sigma}_{t}^{a}$, respectively.
-Proposition 3.2.
-Suppose Assumptions 2 and 3 hold, the model operator $\mathcal{M}_{t}$ is linear, $\mathbf{x}_{0}\mid\mathbf{y}_{0}\sim{\mathrm{N}}_{p}\left(\mathbf{x}_{0}^{a},\boldsymbol{\Sigma}_{0}^{a}\right)$, from which the initial analysis ensemble $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ are independently generated, then, for fixed dimensions $p$ and $q$, the forecast state of the EnKF obtained by (2.9) converges conditionally in probability or in distribution such that $\hat{\mathbf{x}}_{t}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbf{x}_{t}^{f}$, $\hat{\boldsymbol{\Sigma}}_{t}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Sigma}_{t}^{f}$, $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathrm{N}_{p}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$, and the analysis state of the EnKF in (2.12) satisfies $\hat{\mathbf{x}}_{t}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbf{x}_{t}^{a}$, $\hat{\boldsymbol{\Sigma}}_{t}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Sigma}_{t}^{a}$, and $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathrm{N}_{p}(\mathbf{x}_{t}^{a},\boldsymbol{\Sigma}_{t}^{a})$ as $n\rightarrow\infty$, where $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{a}$ are the sample covariance matrices of the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{i=1}^{n}$ and the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{i=1}^{n}$, respectively.
-Proposition 3.2 implies that with linear model operator $\mathcal{M}_{t}$ and Gaussian initial state, the asymptotic conditional distributions of the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{j=1}^{n}$ given $\mathbf{y}_{1:t}$ and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ given $\mathbf{y}_{1:t-1}$ in the EnKF coincide with the distributions of $\mathbf{x}_{t}|y_{1:t}$ and $\mathbf{x}_{t}|y_{1:t-1}$ in the Kalman Filter, respectively.
-The following Theorem 3.1 provides the upper bounds of the mean square error (MSE) between the analysis and forecasting states of the EnKF in comparison to those of the Kalman Filters according to the true underlying evolution given in (2.6)-(2.7) for the one-step assimilation for a general case where the model operator $\mathcal{M}_{t}$ is nonlinear.
-To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
-$\Omega_{t}=$ $\operatorname{tr}\big{(}{n}^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({{\mathbf{H}}_{t}\boldsymbol{\Sigma}_{t}^{f}{\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t}})\big{)}+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f})\|^{2}$
-$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}/n\big{)}+2\operatorname{tr}\big{(}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}\big{)}.$ (3.2)
-Theorem 3.1.
-Suppose Assumptions 1(i) or 1(ii), 2, 3 hold, and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$ with $\mathbf{x}_{t,j}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$ at time $t$.
-Then, the forecast and analysis states of the EnKF at time $t$ satisfy
-$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}$ $={n}^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}{n}^{-1}\big{)},$
-$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $=\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+\Omega_{t}=\operatorname{tr}\big{(}n^{-1}\boldsymbol{\Sigma}_{t}^{a}\big{)}+\Omega_{t}$ (3.3)
-$=O\big{(}{n}^{-1}+\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}=O\big{(}{n}^{-1}\big{)},$
-where the results reported in the $O$-terms are valid for the fixed dimension case while the other results are valid for both fixed and high dimensional cases.
-Note that the first term in (3.3) is the variance of the sample mean of the analysis ensemble using the true forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the second term $\Omega_{t}$ results from the estimation error of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$.
-Theorem 3.1 implies that $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=O\left(n^{-1}\right)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ for the one-step data assimilation under the fixed dimensional case, which means both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$.
-These are because the sample covariance matrix $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ is consistent with the underlying population covariance at the $\sqrt{n}$ rate, and it follows that $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}=O(n^{-1})$.
-The proof of these results can be found in the proof of Theorem 3.1 in the SM.
-It should also be noted that the above results are valid for the model operators $\mathcal{M}_{t}$ given in Assumption 1-(i) and (ii).
-The discrepancies between $\hat{\mathbf{x}}_{t}^{f}$ or $\hat{\mathbf{x}}_{t}^{a}$ and the true state $\mathbf{x}_{t}$ are given by (3.1).
-As $\mathbf{x}_{t}$ is random with $(\mathbf{x}_{t}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$ in model (2.8), $\mathbb{E}\|(\mathbf{x}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})$ and $\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ are constants not related to the ensemble size $n$.
-This gives the result $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=O({n}^{-1}+1)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t})|\mathbf{y}_{1:t-1}\|^{2}$, which is not related to the ensemble size $n$ in the leading order and can not converge to zero.
-This is due to the fact that the underlying state is a random variable.
-In contrast, $\mathbf{x}_{t}^{a}$ is a conditional mean and is possible to be estimated consistently.
-The above justification on the validity of the Kalman Filter and EnKF algorithms is made for the fixed or low dimensional state vectors $\mathbf{x}_{t}$ where $p$ is of a smaller order of the ensemble size $n$ such that $p/n\to 0$.
-As the modern measurement technology on the earth system improves, the spatial resolution of the state variables becomes much higher.
-This makes $\mathbf{x}_{t}$ ultra high dimensional and creates high dimensional challenges on the validity of the EnKF, which is treated in the next section.
-4 Convergence for High Dimensional Case
-Remote sensing technology has been advancing the measurement of massive amounts of data sets for many geophysical processes, which produces high resolution data and brings high dimensional challenges.
-When the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.10) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ as shown in Bai and Yin (1993).
-The inconsistency of the forecast error covariance estimator (2.10) to $\boldsymbol{\Sigma}_{t}^{f}$ is a major threat to the validity of the EnKF method.
-Although there are methods such as the localization and inflation approaches designed to rescue the EnKF in the high dimensional case, the statistical properties of these approaches remain largely unknown.
-The last 15 years have seen rapid development of high dimensional statistical inference, including efforts to construct consistent estimators for high dimensional covariance matrices which have proposed the banding, the thresholding and the tapering estimators (Bickel and Levina, 2008a, b; Cai, Zhang and Zhou, 2010).
-In this section, we first introduce the covariance classes where consistent estimation of high dimensional covariance matrices can be made, and present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$.
-Then, we provide results on the consistent estimation of the Kalman gain matrix $\mathbf{K}_{t}$ and evaluate the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ from the EnKF to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.
-4.1 Consistent estimation of high dimensional forecast error covariance matrix
+Corollary S.3.1 of the SM provides an upper bound of the estimation error of the $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ when both $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
+Based on the result in Proposition 3.1, we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.5)-(2.6) by firstly establishing the convergence of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
+For the fixed dimension case, the sample covariance matrix is consistent such that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$.
+However, when the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.9) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ (Bai and Yin, 1993), which is a major threat to the validity of the EnKF method.
+Next, we present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and their covariance classes before we give the convergence of the EnKF.
+3.1 Consistent estimation of high dimensional $\boldsymbol{\Sigma}_{t}^{f}$
+The banding estimator (Bickel and Levina, 2008b) of the forecast error covariance matrix with bandwidth $k$ is $B_{k}(\mathbf{S}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|i-j|\leq k)\big{]}$, where $0\leq k<p$ and $\mathbf{S}_{t}^{f}=[s_{tij}^{f}]_{p\times p}$ is the sample covariance of the EnKF forecast ensembles.
+The covariance matrix class suitable for the banding estimator is the following bandable covariance class, $\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}$ for some positive $\alpha$, C and $\varepsilon_{0}$.
+This bandable pattern is generally consistent with the dependence between variables observed at sites $i$ and $j$ in geoscience studies, as the dependence becomes weaker as the distance between the two sites gets larger.
+The banding estimator administrates a hard threshold as it sets all matrix elements beyond the $k$-th sub-diagonals to zero.
+Cai et al.
+(2010) updated the banding estimator with a tapering alteration which tapers down the level of the thresholding gradually
+$\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)=\mathbf{S}_{t}^{f}\circ\mathbf{W}(k)=\big{(}w_{ij}(k)s_{tij}^{f}\big{)}_{p\times p},$
+where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\left\{(k-|i-j|)_{+}-\left(k/2-|i-j|\right)_{+}\right\},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$.
 (a) $h(k)$ of $\boldsymbol{\Sigma}\in\mathcal{U}_{b}$
@@
 Figure 1: Bandable and circular bandable covariance matrices: the average energy function $h(k)$ in Panels (a) and (b) and their heatmaps in Panels (c) and (d), where $h(k)$ is the average of the squares of the $k^{th}$ sub-diagonal entries.
-The covariance matrix class suitable for the banding estimator is the following bandable covariance class (Bickel and Levina, 2008a)
-$\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}$ $\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0$
-[…]
-for some positive $\alpha$, C and $\varepsilon_{0}$.
-The bandable class prescribes a decaying pattern of $|\sigma_{ij}|$ as $|i-j|$ increases with the decaying rate $\alpha$.
-This bandable pattern is generally consistent with the dependence between variables observed at sites $i$ and $j$ in atmospheric and geoscience studies, as the dependence becomes weaker as the distance between the two sites gets larger.
-Note that $\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|$ is the $L_{1,1}$ norm of $B_{k}\left(\boldsymbol{\Sigma}\right)-\boldsymbol{\Sigma}$, which bounds the spectral norm from above (Golub and Van Loan, 2013).
-In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or the poles of the earth, the state vectors are highly dependent for the variable components between the two ends of the state vector.
-This leads to a circular form of the bandable structure as follows:
-$\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=$ $(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }$
-$0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$
-Let $h(k)$ be the average of the squares of the $k^{th}$ subdiagonal entries of a covariance:
-$h(k):=\frac{1}{2(p-k)}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=\frac{1}{p-k}\sum_{l=1}^{p-k}\sigma_{ll+k}^{2}.$
-Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariance matrices that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$.
-Figure 1 (a) and (b) show that the $h(k)$ function of the covariance matrix in $\mathcal{U}_{b}$ decays rapidly as the subdiagonal index $k$ increases, while the $h(k)$ function of the covariance matrix in $\mathcal{U}_{bc}$ first decays rapidly, then remains almost unchanged and rises again.
+In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or its poles, the state vectors are highly dependent for the variable components between the two ends of the state vector.
+Thus we propose a circular form of the bandable structure, $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$ Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariance matrices that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$, where $h(k)$ is the average of the squares of the $k^{th}$ subdiagonal entries of a covariance defined by $h(k):=\frac{1}{2(p-k)}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=\frac{1}{p-k}\sum_{l=1}^{p-k}\sigma_{ll+k}^{2}.$ Figures 1 (a)-(b) show that the $h(k)$ function of covariances in $\mathcal{U}_{b}$ is a $L$-type curve as the subdiagonal index $k$ increases, while it is an $U$-type curve for covariances in $\mathcal{U}_{bc}$.
+For covariances in the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, the proposed mid-banding forecast error covariance estimator with bandwidths $k_{1}$ and $k_{2}$ is $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})=[s_{tij}^{f}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))]$, where $0\leq k_{1}<(p-k_{2})\leq p$.
 The bandable matrix class requires that the state vector $\mathbf{x}_{t}$ has an ordering (or a permutation) such that the correlation decays as two components of $\mathbf{x}_{t}$ are further apart.
-This may be argued for climate and geospatial applications, where components of the state vectors can be arranged via the geo-distance of the observational sites where the state vectors are associated with.
 However, the presence of geo-spatial processes (weather front or hurricane) and geography (mountains or ocean) will interrupt the spatial dependence ordering prescribed by the banding covariance class.
-In such situations, the following thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008b) may be employed.
-The thresholding covariance class is a class of covariance matrices which are invariant under permutations of components of the state vector $\mathbf{x}_{t}$, defined by
-$\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,$ $\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i$
-[…]
-We now present the estimators of the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding based on the sample covariance $\mathbf{S}_{t}^{f}$ of the EnKF ensembles.
-For any matrix $M=\left[m_{ij}\right]_{p\times p}$, and any $0\leq k<p$, the banding matrix operator with a bandwidth $k$ is
-$B_{k}(M)=\left[m_{ij}\mathbf{1}(|i-j|\leq k)\right].$
-The banding estimator of the forecast error covariance matrix with bandwidth $k$ is $B_{k}(\mathbf{S}_{t}^{f})$.
-The banding estimator administrates a hard threshold as it sets all matrix elements beyond the $k$-th sub-diagonals to zero.
-Cai, Zhang and Zhou (2010) updated the banding estimator with a tapering alteration
-$\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)=\mathbf{S}_{t}^{f}\circ\mathbf{W}(k)=\big{(}w_{ij}(k)s_{tij}^{f}\big{)}_{p\times p},$
-where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\left\{(k-|i-j|)_{+}-\left(k/2-|i-j|\right)_{+}\right\},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$.
-It is noted that the tapering estimator is a softer version of the banding estimator by tapering down the level of the thresholding gradually.
-In order to provide valid estimation for forecast error covariance belonging to the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, we define a mid-banding operator with bandwidths $k_{1}$ and $k_{2}$,
-$B_{k_{1},k_{2}}(M)=\left[m_{ij}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))\right],$
-for any matrix $M=\left[m_{ij}\right]_{p\times p}$ and $0\leq k_{1}<(p-k_{2})\leq p$. Then the mid-banding forecast error covariance matrix estimator is $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$.
-Recall that both the banding and tapering estimators require the variables in $\mathbf{x}_{t}$ to have a natural ordering under a permutation such that the correlation decays as two variables are further apart.
-For random vectors that cannot be ordered in such a way but belong to the thresholding covariance class $\mathcal{U}_{\tau}\left(q,c_{0}(p),M\right)$, Bickel and Levina (2008b) proposed the thresholding operator
-$T_{s}(M)=\left[m_{ij}1\left(\left|m_{ij}\right|\geq s\right)\right],$
-at a threshold level $s$. Then the thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})$.
-Note that all of the banding, mid-banding, tapering and thresholding estimators depend on either the bandwidth $k$ or the threshold $s$.
-The selection of these parameters can be carried out by methods in Qiu and Chen (2015).
-For the bandwidths $k_{1},k_{2}$ of the mid-banding estimator, we propose a similar selection method and details can be found in the SM.
-To apply the high dimensional estimators of the forecasting error covariance $\boldsymbol{\Sigma}_{t}^{f}$, we need the following distributional assumption.
-Assumption 4.
-The random vectors $\mathbf{x}_{t,i}^{f}=(x_{t,i1}^{f},\cdots,x_{t,ip}^{f})^{\mathrm{T}}$ are IID and each $x_{t,1j}^{f}\stackrel{{\scriptstyle d}}{{\sim}}F_{t,j}$ and $(x_{t,1j}^{f})^{2}\stackrel{{\scriptstyle d}}{{\sim}}G_{t,j}$, where $G_{t,j}(u)=F_{t,j}(\sqrt{u})-F_{t,j}(-\sqrt{u})$ for $u>0$.
-There exists a $\lambda_{0}>0$ such that $\max_{1\leq j\leq p}\mathbb{E}(e^{\lambda(x_{t,1j}^{f})^{2}})<\infty\textrm{ for }\lambda\in(-\lambda_{0},\lambda_{0}).$
-Assumption 4 effectively assume $(x_{t,1j}^{f})^{2}$ are sub-exponentially distributed, which include the Sub-Gaussian random variables as a special case.
-The following Proposition 4.1 summarizes the convergence rates of the forecast error covariance matrix estimators by the banding, mid-banding, tapering and thresholding methods.
+In such situations, the thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008a) may be employed, where $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i\text{ and }\lambda_{\min}(\boldsymbol{\Sigma})\geq\varepsilon_{0}>0\big{\}},\text{ for }0\leq\gamma<1.$ The thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|s_{tij}^{f}|\geq s)\big{]}$.
+Note that all of the banding, mid-banding, tapering and thresholding estimators depend on either the bandwidth $k$ or the threshold $s$ which can be selected by methods in Qiu and Chen (2015).
+For the bandwidths $k_{1},k_{2}$ of the mid-banding estimator, we propose a similar selection method and details can be found in Section S.2 of the SM.
+The following Proposition 3.2 summarizes the convergence rates of the forecast error covariance estimators by the banding, mid-banding, tapering and thresholding methods.
 Proposition 3.2.
-Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally IID Gaussian or sub-exponentially distributed as in Assumption 4.
+Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally IID Gaussian or sub-exponentially distributed as in Assumption S.3.1 of the SM.
   • (i)
-    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$,
-    $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
+    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
   • (ii)
-    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$,
-    $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
+    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
   • (iii)
-    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$, suppose $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are sub-Gaussian, then for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$ the optimal rate of convergence of the tapering estimator is
-    $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
+    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$, suppose $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are sub-Gaussian, then for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$, $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
   • (iv)
-    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$ and $n^{-1}{\log p}=o(1)$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$,
-    $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}.$
-The results in Proposition 4.1 (i), (iii) and (iv) are, respectively, from Bickel and Levina (2008a), Cai, Zhang and Zhou (2010) and Bickel and Levina (2008b).
+    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$, $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}.$
+The results in Proposition 3.2 (i), (iii) and (iv) are, respectively, from Bickel and Levina (2008b), Cai et al.
+(2010) and Bickel and Levina (2008a).
 And that for the mid-banding estimator in Proposition 3.2 (ii) is a new result whose proof is given in the SM.
@@
 According to Proposition 3.1, the convergence results in Proposition 3.2 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$.
-Specifically, if $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$ or $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$,
-$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}\,\text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k}(\mathbf{S}_{t}^{f})\text{ or }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f});\,$ (4.1)
+Specifically,
 […]
-$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}\text{ for }\,\hat{\boldsymbol{\Sigma}}_{t}^{f}=\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k);$ (4.2)
 […]
-$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}\ \text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=T_{s}(\mathbf{S}_{t}^{f}).$ (4.3)
-These suggest that under the conditions of Proposition 3.1, $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ has the same convergence rate as $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$ for any of the four high dimensional estimators.
+[…]
 As will be shown in the following, the estimation error of the Kalman gain matrix $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ plays a crucial role in quantifying the effects of the high dimensional  banding, the mid-banding, the tapering and the thresholding estimators for $\boldsymbol{\Sigma}_{t}^{f}$.
 3.2 Convergence of High Dimensional Ensemble Kalman Filter
-When the dimension $p$ of the state vector is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 4.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$.
-Based on the consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman gain matrix $\mathbf{K}_{t}$, we propose the high dimensional EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$.
-It is noted that for nonlinear model operator, $\mathbf{x}_{t}^{a}$ is defined as (2.6)-(2.7) as indicated before (2.13).
-Similar to Theorem 3.1, the following Theorem 4.1 presents the upper bounds of the discrepancy between the EnKF and the Kalman Filters for the one-step data assimilation.
+When the dimension $p$ of $\mathbf{x}_{t}$ is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 3.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$.
+Based on the consistent estimators of $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{K}_{t}$, we propose the HD-EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$.
+To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
+$\Omega_{t}=$ $\operatorname{tr}\big{(}{n}^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({{\mathbf{H}}_{t}\boldsymbol{\Sigma}_{t}^{f}{\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t}})\big{)}+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f})\|^{2}$
+$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}/n\big{)}+2\operatorname{tr}\big{(}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}\big{)}.$ (3.2)
+The following Theorem 3.1 presents the upper bounds of the discrepancy between the EnKF and the Kalman Filters for the one-step data assimilation.
 Theorem 3.1.
-Suppose Assumptions 1(i) or 1(ii), 2, 3 hold and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the conditional distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t,j}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$ at time $t$.
-Then, when $p,q\rightarrow\infty$, the forecast and analysis states of the EnKF at time $t$ satisfy
+Suppose Assumptions 1(i) or 1(ii), 2, 3 hold and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t}^{f}$ and covariance $\boldsymbol{\Sigma}_{t}^{f}$.
+Then, the forecast and analysis states of the EnKF at time $t$ satisfy
 $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}$ $=$ ${n}^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}n^{-1}\big{)},$
@@
 $=$ $O\big{(}n^{-1}+p^{-1}\Delta_{\mathbf{K},t}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}.$
-Theorem 4.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (4.1)-(4.3).
+Theorem 3.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (3.1).
 Specifically, with the banding or mid-banding estimators, the tapering estimator and the thresholding estimator, the one-step EnKF data assimilation MSE are, respectively,
@@
 $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}qc_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}\big{\}}.$ (3.6)
-To show the impact of the dimensionality on the MSE, we take the banding and the mid-banding estimators as an example.
-If $q\asymp p$, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O((n^{-1}{\log p})^{2\alpha/(2(\alpha+1))})$, which implies that the MSE between the analysis states of the EnKF $\hat{\mathbf{x}}_{t}^{a}$ and the $\mathbf{x}_{t}^{a}$ mainly depends on the estimation error of the Kalman gain matrix.
-Commonly in high resolution data assimilation, the dimension of the observation vector $\mathbf{y}_{t}$ is of much smaller order than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$.
-Assume $q=p^{\beta}$ for a $\beta\in(0,1)$, then the one-step assimilation MSE can be written as
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}n^{-1}+n^{-1}p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\big{)}.$
-If $\beta$ is large so that $p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\rightarrow\infty$, the estimation error of the Kalman gain matrix will dominate the MSE, that is, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$.
-If $\beta$ is small such that $p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\rightarrow 0$, then $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O(n^{-1})$.
+For fixed $p,q$, both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$ as $\Delta_{\mathbf{K},t}=O(n^{-1})$, as also shown in Theorem S.3.1 of the SM.
+Furthermore, Proposition S.3.1 implies that with linear model operator $\mathcal{M}_{t}$ and Gaussian initial state, the asymptotic distributions of $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}$ in the EnKF are the same as those of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t}|\mathbf{y}_{1:t-1}$ in the Kalman Filter, respectively.
+To show the impact of the dimensionality on the MSE, we take the banding and mid-banding estimators as examples.
+If $q\asymp p$, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O((n^{-1}{\log p})^{2\alpha/(2(\alpha+1))})$, which implies that the one-step assimilation MSE mainly depends on the estimation error of the Kalman gain matrix.
+Commonly in high resolution data assimilation, the dimension of the observation $\mathbf{y}_{t}$ is much smaller than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$.
+If $q$ is relatively large such that $p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\}\gg n^{-1}$, then the estimation error of the Kalman gain matrix will dominate, that is, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$.
+Otherwise, the first term $n^{-1}$ dominates.
 For the impact of the observation dimension $q$ on the MSE, a study reported in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease as the first term in (3.1) will decrease as $q$ increases.
+For the discrepancies between the EnKF analysis state and the true state, as $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ and $\mathbf{x}_{t}$ is a random variable with $\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ which is a constant not related to the ensemble size $n$ for a fixed $p$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ can not converge to zero.
+In practice, the data assimilation process can be repeated for multiple time steps ahead to make longer-time predictions.
+The following Theorem 3.2 gives the $h$-step MSE bounds between the forecast/analysis states of the EnKF in comparison to the underlying Kalman Filters.
+To simplify writing, define two non-stochastic constants
+$C_{t,h}^{A}=\frac{1-L_{t,h}^{h}A_{t,h}^{h}}{1-L_{t,h}A_{t,h}}\,\text{ and }C_{t,h}^{B}=\frac{1-L_{t,h}^{h}B_{t,h}^{h}}{1-L_{t,h}B_{t,h}},$ (3.7)
+where $A_{t,h}=\max_{t<k\leq t+h}\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|$, $B_{t,h}=\max_{t<k\leq t+h}\{\mathbb{E}\|\mathbf{I}-\hat{\mathbf{K}}_{k}\mathbf{H}_{k}\|\}$, and $L_{t,h}=\max_{0<k\leq h}\{\ell_{0},\ell_{t+k}\}$ for a fixed $t$, with $\ell_{0}$ and $\ell_{t}$ defined in Assumption 1.
+We further denote
+$U_{k}=\big{\{}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\Delta_{\mathbf{K},k}+\operatorname{tr}\big{(}\mathbb{E}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})^{\mathrm{T}}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})({\mathbf{H}}_{k}\boldsymbol{\Sigma}_{k}^{f}{\mathbf{H}}_{k}^{\mathrm{T}}+\mathbf{R}_{k})\big{)}\big{\}}^{1/2}.$ (3.8)
+Theorem 3.2.
+Assumptions 1(i) or 1(ii), 2, 3 hold and the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t}$ with mean $\mathbf{x}_{t}^{a}$ and covariance $\boldsymbol{\Sigma}_{t}^{a}$ at time $t$.
+By repeatedly performing the ensemble forecast-analysis process $h$ steps, the forecast and analysis states of the EnKF at the time $t+h$ satisfy
+$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f})|\mathbf{y}_{1:t+h-1}\|^{2}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{f})\}^{1/2}+L_{t,h}C_{t,h-1}^{B}\max_{t<k\leq t+h-1}U_{k}\Big{)}^{2}$
+$=O\Big{(}n^{-1}+p^{-1}L_{t,h}^{2}(C_{t,h-1}^{A})^{2}\max_{t<k\leq t+h-1}\big{\{}\Delta_{\mathbf{K},k}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\big{\}}\Big{)},$
+$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{a})\}^{1/2}+C_{t,h}^{B}\max_{t<k\leq t+h}U_{k}\Big{)}^{2}$
+$=O\Big{(}n^{-1}+p^{-1}(C_{t,h}^{A})^{2}\max_{t<k\leq t+h}\big{\{}\Delta_{\mathbf{K},k}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\big{\}}\Big{)}.$
+Compared with the one-step results given in Theorem 3.1, the term related to the estimation error of $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the $h$-step analysis MSE is inflated with constants $(C_{t,h}^{A})^{2}$ which are related to $L_{t,h}$ and $A_{t,h}$ defined before (3.7).
+However, for any fixed $h$, $\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}\asymp\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}.$
+Sometimes, the assimilation process proceeds every $k$ time steps not every step as one may not have observations at each step or the assimilation in every step is too computationally consuming.
+Theorem S.3.2 of the SM gives the MSE bounds for multiple-step forecasting and assimilation where the assimilation process proceeds every $k$ time steps.
 4 Imperfect Models
@@
 […]
-Theorem 5.1 presents the bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter for the one-step assimilation when $\mathcal{M}_{t}$ is unknown. Denote
+Theorem 4.1 presents the one-step assimilation bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter when $\mathcal{M}_{t}$ is unknown. Denote
 $V_{t}$ $=\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\Xi_{\xi_{t}}\big{)}+\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2}$
@@
 Theorem 4.1.
-Suppose Assumptions 1(i) or 1(ii), 2, 3, 5 hold and the analysis ensemble $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t-1}^{a}$ and covariance $\boldsymbol{\Sigma}_{t-1}^{a}$ at time $t-1$.
+Suppose Assumptions 1(i) or 1(ii), 2, 3, 4 hold and $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t-1}^{a}$ and covariance $\boldsymbol{\Sigma}_{t-1}^{a}$ at time $t-1$.
 Then, the forecast and analysis states of the EnKF satisfy
@@
 where $\tilde{\Delta}_{\mathbf{K},t}=\mathbb{E}\|\tilde{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ and $\tilde{\mathbf{K}}_{t}$ is the estimate of $\mathbf{K}_{t}$ with true model $\mathcal{M}_{t}$.
-[…]
-$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}n^{-1}+\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}(\|\Xi_{\xi_{t}}\|^{2}+n^{-1})\big{)}=O\big{(}n^{-1}+\|\Xi_{\xi_{t}}\|^{2})\big{)}.$ (5.4)
-Thus, if $\|\Xi_{\xi_{t}}\|^{2}=O(n^{-1})$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=O\left(n^{-1}\right)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$, which means that the model error does not alter the order of magnitude of the assimilation and forecasting MSEs of the EnKF.
-However, if $n\|\Xi_{\xi_{t}}\|^{2}\rightarrow\infty$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O(\|\Xi_{\xi_{t}}\|^{2})$, which means the modelling error dominates the upper bounds of the MSE in the analysis states of the EnKF.
-When the dimensions $p$ and $q\rightarrow\infty$,
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}\asymp\max\{n^{-1},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$ (5.5)
-[…]
+The theorem suggests that the model approximation error creates extra terms related to $\mu_{\xi_{t}}$ and $\|\boldsymbol{\Xi}_{\xi_{t}}\|$, which are absent from the results in Theorem 3.1 where accurate model is assumed.
+Specifically, the assimilation MSEs for fixed $p,q$ and $p,q\rightarrow\infty$ are, respectively,
 $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}\text{ and }$ (4.4)
 $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
-The above analysis shows that for both the fixed and the diverging dimension, and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the MSE of the assimilated $\hat{\mathbf{x}}_{t}^{a}$ by the EnKF, which highlights the benefits of having more accurate model operators.
+[…]
+The above analysis shows that for both the fixed and diverging dimensions and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the assimilation MSE, which suggests that a good approximation of the model operator is a significant aspect of data assimilation.
 While this may be obvious, Theorem 4.1 describes exactly how it happens.
-In summary, the approximation error of the model operator $\mathcal{M}_{t}$ has a strong influence on the MSE of the data assimilation of the EnKF.
-This suggests that a good approximation of the model operator is a significant aspect of data assimilation.
-We also introduce an iterative EnKF update algorithm in S.1.2 to account for the model approximation error to some extent since most geophysical models are imperfect in practice.
+The accumulation of the model approximation error is more pronounced in the $h-$step assimilation MSEs as shown in Theorem S.3.3 of the SM.
+Thus, an iterative HD-EnKF algorithm is introduced in S.1.2 to account for the model approximation error to some extent since most geophysical models are imperfect in practice, of which the main idea is to substitute $\hat{\mathbf{x}}_{t}^{f}$ in the original forecast sample covariance matrix by the analysis state $\hat{\mathbf{x}}_{t}^{a}$ in the iteration.
 5 Simulation
-This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed high dimensional EnKF algorithms.
+This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed HD-EnKF algorithms.
 These two models are often used for testing new approaches in atmospheric sciences for both data assimilation and the spatial-temporal forecasts.
-The proposed high dimensional EnKF methods are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models.
-Details of the proposed high dimensional EnKF algorithms are given in Section S.1 of the SM, including the consistent EnKF update in S.1.1 and the iterative EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators.
+The proposed HD-EnKF methods are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models.
+Details of the proposed HD-EnKF algorithms are given in Section S.1 of the SM, including the consistent HD-EnKF update in S.1.1 and the iterative HD-EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators.
 5.1 Lorenz-96 model
-Lorenz-96 model (L96) (Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere.
+Lorenz-96 model (L96; Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere.
 It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation
-${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),$ (6.1)
-for $j=1,\ldots,p$, with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denote the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $-\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is the force term.
-The $F$ can be set as other constants to generate imperfect models, the context discussed in Section 5 .
-L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016), which approximates the model as $\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)=\mathbf{x}_{t}+h(k_{1}+2k_{2}+2k_{3}+k_{4})/{6},$ where $k_{1}=\mathbf{f}\left(\mathbf{x}_{t},t\right)$, $k_{2}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{1}/2,t+h/2\right)$, $k_{3}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{2}/2,t+h/2\right)$, $k_{4}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{3},t+h\right)$ and $\mathbf{f}=\left(f_{1},f_{2},\dots,f_{p}\right)^{\mathrm{T}}$.
-An approximation to the true state was simulated by either (i) implementing the RK4 method with a finer time step, say $h=0.05$, which corresponds to $6$ hours in the real-time assuming that the time-scale of the dissipation in the atmosphere is $5$ days (Lorenz, 1996); or (ii) by adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$.
+${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),\,\text{ for }j=1,\ldots,p,$ (5.1)
+with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denotes the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $-\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is a forcing term.
+L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016).
+An approximation to the true state was simulated by (i) implementing the RK4 method with a finer time step, say $h=0.05$; and (ii) adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$.
 In our experiment, $\mathbf{w}_{t}\sim\mathrm{N}(0,\sigma_{0}\mathbf{I}_{p})$ and $\sigma_{0}$ were set as $0.1,0.25$ and $0.5$, respectively.
-Furthermore, we experimented $p=40$ and $100$, under which the system behaves chaotically, since the L96 model behaves like chaos for $p\geq 7$ (Bedrossian, Blumenthal and Punshon-Smith, 2022).
+We experimented $p=40$ and $100$, under which the system behaves chaotically, since the L96 model behaves like chaos for $p\geq 7$ (Bedrossian et al., 2022).
 The true forcing was $F=8$.
-Similar to Liang et al.
-(2012), we also considered the misspecified model setting by making $F=6,7,9,10$ in the assimilation processes, corresponding to the imperfect models in Section 5.
-The initial state $\mathbf{x}_{1}$ was such that $x_{1}^{(j)}=F$ for $j\neq 20$ and $x_{1}^{(20)}=F+0.001$, and the initial forecast ensemble was generated by $\mathbf{x}_{1,j}^{f}=\mathbf{x}_{1}+\mathrm{N}_{p}(0,0.1\mathbf{I}_{p})$ for $j=1,\ldots,n$.
-The observations $\mathbf{y}_{t}$ were made at $q$ randomly selected components of the state vector $\mathbf{x}_{t}$ with $q=30,40$ and $100$ while $p=40$ and $100$, respectively, followed by adding observation errors $\varepsilon_{t}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{q}(0,\mathbf{R}_{t})$ to the true states.
-The $(i,j)-$element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,p-|i-j|\right\}}$, which prescribes spatially correlated observation errors to suit geophysical data.
-The observation operator $\mathbf{H}_{t}$ was a binary $q\times p$ matrix with the $(j,i)$ element being $1$ if and only if the $j-$th component of $\mathbf{y}_{t}$ was the observed $i-$th dimension of the state $\mathbf{x}_{t}$.
+We also considered the misspecified model setting by making $F=6,7,9,10$ in the assimilation processes, corresponding to the imperfect models in Section 4.
+The initial state was $x_{1}^{(j)}=F$ for $j\neq 20$ and $x_{1}^{(20)}=F+0.001$, and the initial forecast ensemble was generated by $\mathbf{x}_{1,j}^{f}=\mathbf{x}_{1}+\mathrm{N}_{p}(0,0.1\mathbf{I}_{p})$ for $j=1,\ldots,n$.
+The observations $\mathbf{y}_{t}$ were made at $q$ randomly selected components of the state vector $\mathbf{x}_{t}$ with $q=30,40$ and $100$ while $p=40$ and $100$, respectively, followed by adding observation errors $\boldsymbol{\varepsilon}_{t}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{q}(0,\mathbf{R}_{t})$ to the true states.
+The $(i,j)-$element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,q-|i-j|\right\}}$, which prescribes spatially correlated observation errors.
+The observation operator $\mathbf{H}_{t}$ was a $q\times p$ matrix with the $(j,i)$ element being $1$ if and only if the $j-$th component of $\mathbf{y}_{t}$ was the observed $i-$th dimension of the state $\mathbf{x}_{t}$.
 We simulated observations every $4$ time steps for a total of $2000$ steps and reported the results for the last $1000$ steps to allow the system to settle down.
 The ensemble size $n$ was $20,30,60$ and $90$.
-All the simulation experiments were repeated 500 times.
-We considered the high dimensional EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance to estimate the forecast error covariance without the proposed high dimensional treatment and the inflation methods (Liang et al., 2012).
-The performances of the assimilation methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$, averaged over the last 1000 time steps.
-However, the underlying analysis states $\mathbf{x}_{t}^{a}$ were unknown in the simulation of L96 due to its nonlinear nature.
-Indeed, for a nonlinear $\mathcal{M}_{t}$, the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ could not be readily updated by (2.3) in the Kalman Filter.
-To gain an accurate estimate of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance matrices to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman filter update formula to obtain $\mathbf{x}_{t}^{a}$, since the large ensemble size would make the sample covariance a high quality substitute for $\boldsymbol{\Sigma}_{t}^{f}$.
+All the experiments were repeated 500 times.
+We considered the HD-EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance to estimate the forecast error covariance and the inflation methods (Liang et al., 2012).
+The performances of these methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$.
+To gain accurate estimations of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance matrices to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman filter update formula to obtain $\mathbf{x}_{t}^{a}$.
 Figure 2: Average RMSEs to the oracle analysis state and their $25-75\%$ quantile bands (color bands) as a function of forcing $F$ used in ${\cal{M}}_{t}$ (true value is 8) by the standard EnKF (blue), the modified EnKF methods: the inflation (red), banding (green), tapering (cyan) and thresholding (purple) for $p=q=40$ and $100$, $n=20,30,60,90$, $\sigma_{0}=0.1,0.25$ and $0.5$, respectively.
 The light-colored bars in the lower part of each panel denote the divergence rates of the assimilation schemes.
-Figure 2 displays the average RMSEs to the oracle analysis state of the five assimilation methods, namely the standard EnKF, the inflation method (Liang et al., 2012) and the three high dimensional EnKF equipped with the banding and tapering (both respect the circular bandable structures), and the thresholding estimators, with respect to different forcing level $F$ used in the model at moderate dimension $p=40$ and high dimension $p=100$ under different $\sigma_{0}$ and $n$.
-It shows that the EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) model settings.
-Even when $F$ became more deviated from the underlying value $8$, the analysis states obtained by the high dimensional EnKFs were still much closer to the oracle analysis states than the inflation method although both strains of methods showed increased RMSEs with respect to $|F-8|$.
-The standard EnKF with the sample covariance estimator to $\boldsymbol{\Sigma}_{t}^{f}$ had much larger assimilation error in all cases.
-Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), we see that the advantages of the proposed high dimensional EnKF schemes over the inflation method were more apparent for the higher dimension setting $p=100$.
-The analysis RMSEs of the high dimensional EnKF methods kept at a lower level especially when the force $F$ was equal or close to the true value ($F=8$), while the analysis RMSE of the inflation method ranked behind but better than the standard EnKF.
-This means that the proposed high dimensional EnKF methods are quite attractive and useful as the geophysical models in earth science are of ultra high dimension.
-The figure also displays the percentages of divergence of the five methods, which shows that the inflation method endured quite high divergences, even higher than the standard EnKF.
-In contrast, the three high dimensional EnKF methods had quite low divergence rates for all cases.
-The divergence rate of the inflation method increases substantially along with the increase of the model error, which, for instance, reached 50% when $\sigma=0.5$.
-Panels (d), (g)-(f) of Figure 2 demonstrate the influence of the ensemble size on the analysis RMSEs, which show that the analysis RMSEs of the inflation and the high dimensional EnKF methods became smaller with the increase of the ensemble size $n$.
-The proposed high dimensional EnKF methods performed much better than the inflation method with smaller RMSEs, especially when the ensemble size $n$ was small (panels (d) and (g) with $n=20$).
-This suggests that the proposed high dimensional EnKF methods could provide more accurate analysis states when the ensemble size was not big enough.
-Ensemble size is a significant restriction for the running of large geophysical models as they are extremely costly to run.
-Table 1: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes for Lorenz-96 model with $p=q=40,100$, $n=30$, $\sigma_{0}=0$ and $F=8$ for the correct and other $F$ values for the misspecified models.
-The numbers inside the parentheses are the $25\%$ and $75\%$ quantiles of the RMSEs.
-p type method F=6 F=7 F=8 F=9 F=10
-$40$ oracle standard 4.27 (4.22,4.33) 4.37 (4.30,4.44) 4.62 (4.54,4.69) 4.92 (4.85,4.99) 5.23 (5.17,5.30)
-inflation 1.08 (1.06,1.10) 0.83 (0.81,0.84) 0.60 (0.58,0.62) 0.93 (0.91,0.94) 1.21 (1.18,1.23)
-banding 0.88 (0.87,0.88) 0.75 (0.74,0.76) 0.69 (0.68,0.70) 0.79 (0.78,0.80) 0.96 (0.95,0.97)
-tapering 0.88 (0.87,0.88) 0.75 (0.74,0.75) 0.68 (0.67,0.69) 0.80 (0.79,0.81) 0.97 (0.96,0.99)
-threshold 0.91 (0.90,0.92) 0.78 (0.77,0.79) 0.51 (0.38,0.57) 0.82 (0.81,0.83) 0.98 (0.97,0.99)
-true standard 4.31 (4.25,4.36) 4.39 (4.32,4.47) 4.62 (4.54,4.69) 4.91 (4.84,4.98) 5.22 (5.16,5.29)
-inflation 1.18 (1.16,1.20) 0.88 (0.86,0.89) 0.59 (0.57,0.62) 0.96 (0.94,0.98) 1.27 (1.25,1.29)
-banding 0.98 (0.97,0.99) 0.80 (0.79,0.81) 0.71 (0.70,0.72) 0.83 (0.82,0.84) 1.04 (1.03,1.06)
-tapering 0.98 (0.97,0.99) 0.80 (0.79,0.81) 0.70 (0.69,0.71) 0.84 (0.83,0.85) 1.06 (1.05,1.07)
-threshold 1.02 (1.01,1.03) 0.83 (0.82,0.84) 0.57 (0.45,0.62) 0.87 (0.86,0.88) 1.07 (1.06,1.08)
-$100$ oracle standard 4.40 (4.37,4.44) 4.58 (4.54,4.62) 4.84 (4.80,4.88) 5.14 (5.10,5.18) 5.45 (5.41,5.49)
-inflation 2.91 (2.88,2.95) 2.83 (2.79,2.88) 2.88 (2.83,2.92) 3.11 (3.05,3.15) 3.40 (3.36,3.44)
-banding 1.15 (1.13,1.16) 0.74 (0.73,0.75) 0.50 (0.49,0.50) 0.89 (0.87,0.91) 1.68 (1.63,1.72)
-tapering 1.27 (1.24,1.30) 0.75 (0.74,0.76) 0.46 (0.45,0.47) 1.08 (1.05,1.11) 2.56 (2.46,2.65)
-threshold 1.72 (1.65,1.79) 1.04 (1.02,1.06) 0.83 (0.82,0.85) 1.15 (1.12,1.18) 1.89 (1.82,1.96)
-true standard 4.43 (4.40,4.46) 4.59 (4.55,4.63) 4.84 (4.80,4.88) 5.13 (5.10,5.17) 5.44 (5.40,5.47)
-inflation 2.94 (2.91,2.98) 2.85 (2.80,2.89) 2.87 (2.83,2.92) 3.10 (3.05,3.15) 3.41 (3.37,3.45)
-banding 1.26 (1.25,1.28) 0.84 (0.83,0.85) 0.60 (0.59,0.60) 0.97 (0.95,0.99) 1.75 (1.71,1.79)
-tapering 1.38 (1.35,1.41) 0.85 (0.84,0.86) 0.57 (0.56,0.58) 1.15 (1.12,1.18) 2.60 (2.51,2.69)
-threshold 1.76 (1.70,1.84) 1.06 (1.04,1.08) 0.82 (0.80,0.84) 1.15 (1.12,1.19) 1.92 (1.85,1.99)
-Table 1 reports the average analysis RMSEs to the oracle analysis state and true state and their $25\%$ and $75\%$ quantiles by the standard EnKF, the inflation and the high dimensional EnKF methods using the banding and tapering and the thresholding estimators when $\sigma_{0}=0$ which is the same setting as Liang et al.
-(2012).
-One can see that the RMSEs to the oracle analysis state and true state of the proposed high dimensional methods were much smaller than the standard EnKF and the inflation method under both the correct and misspecified models.
-In the case of a higher dimensional setting with $p=100$, the effectiveness of the proposed high-dimensional EnKF schemes became even more pronounced.
-In contrast, the inflation method demonstrated significantly higher analysis RMSEs, even when applied to the correct model ($F=8$).
-For instance, when $p=100,F=8$, the banding method achieved an RMSE of 0.5, while the inflation method resulted in a much higher RMSE of 2.88.
-Besides, the high dimensional EnKF with banding estimator performed the best with the smallest RMSEs in almost all considered settings.
-Consistent with the theoretical results, the RMSEs to the true state were largely higher than those to the oracle analysis state.
-Table 2: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes and their divergence rates (in the parentheses) for Lorenz-96 model for $p=40$, $q=30$, $n=30$, $\sigma_{0}=0.1$, and $F=8$ for the correct and other $F$ values for the misspecified models.
+Figure 2 displays the average RMSEs to the oracle analysis state of the five assimilation methods, namely the standard EnKF, the inflation method and the three HD-EnKFs equipped with the banding and tapering (both respect the circular bandable structures), and the thresholding estimators, under different forcing $F$ at moderate dimension $p=40$ and high dimension $p=100$ with different $\sigma_{0}$ and $n$.
+It shows that the HD-EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) models.
+Even when $F$ deviated more from the true value $8$, the HD-EnKFs still achieved analysis states that were much closer to the oracle than the inflation method although both strains of methods showed increased RMSEs with respect to $|F-8|$.
+The standard EnKF with the sample covariance estimator to $\boldsymbol{\Sigma}_{t}^{f}$ had much larger assimilation errors in all cases.
+Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), the advantages of the proposed HD-EnKFs over the inflation method were more apparent for the higher dimension $p=100$.
+The analysis RMSEs of the HD-EnKFs kept at a lower level especially when the force $F$ was equal or close to the true value ($F=8$), while the analysis RMSE of the inflation method ranked behind but better than the standard EnKF.
+This means that the HD-EnKFs are quite attractive as the geophysical models in earth science are of ultra high dimension.
+The figure also displays the percentages of divergence of the five methods.
+It shows that the inflation method endured quite high divergences which increased substantially with the model error (e.g.
+reached 50% when $\sigma=0.5$), even higher than the standard EnKF.
+In contrast, the three HD-EnKF methods had quite low divergence rates for all cases.
+Panels (d), (g)-(f) of Figure 2 demonstrate the influence of the ensemble size on the analysis RMSEs.
+Although the analysis RMSEs of the inflation and the HD-EnKF methods all became smaller with the increase of the ensemble size $n$, the proposed HD-EnKFs outperformed the inflation method, especially when the ensemble size $n$ was small (panels (d) and (g) with $n=20$).
+This suggests that the proposed methods could provide more accurate analysis states when the ensemble size was not big enough.
+Ensemble size is a significant restriction for the large geophysical models as they are extremely costly to run.
+Table 1: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes and their divergence rates (in parentheses) for Lorenz-96 model with $p=40$, $q=30$, $n=30$, $\sigma_{0}=0.1$, and $F=8$ for the correct and other $F$ values for misspecified models.
 method F=6 F=7 F=8 F=9 F=10
@@
 threshold 1.24 (0.06) 1.00 (0.02) 0.93 (0.04) 1.09 (0.04) 1.44 (0.16)
-Table 2 displays the analysis RMSEs of different assimilation methods and their divergence rates (in the parentheses) with sparser observations $q=30$ while $p=40$, which illustrates the influence of the sparsity of the observations on the assimilation accuracy.
-We see that the analysis RMSEs of different assimilation methods were slightly higher than those under more observations $p=q=40$.
-The RMSEs of the high dimensional EnKF methods were smaller than those of the inflation method and the standard for imperfect models ($F\neq 8$) and were comparable to the inflation method for the correct model ($F=8$).
-However, it is noted that the inflation method had a higher divergence rate which was 63% even under the correct model ($F=8$) and much higher (86%) for misspecified models ($F=10$).
-In contrast, the proposed high dimensional EnKF methods had low divergence rates, showing their robustness to the sparsity of the observations.
+Table 1 displays the analysis RMSEs of different methods and their divergence rates under sparser observations $q=30$ while $p=40$, which illustrates the influence of the sparsity in the observations on the assimilation accuracy.
+Although the analysis RMSEs of different assimilation methods were slightly higher than those under more observations $p=q=40$, the RMSEs of the HD-EnKF methods were smaller than those of the other methods for imperfect models ($F\neq 8$) and were comparable to the inflation method for the correct model ($F=8$).
+However, it is noted that the inflation method had a higher divergence rate which reached 63% even under the correct model ($F=8$) while the proposed methods had low divergence rates, showing their robustness to the sparsity of the observations.
 5.2 Shallow Water Equation model
@@
 The SWE is a system of hyperbolic/parabolic partial differential equations governing fluid flow, which is widely used in river flow simulations, as well as in other settings where the water depth is significantly lower than the horizontal length scale of motion.
-The barotropic nonlinear SWE takes the following form (Lei, Stauffer and Deng, 2012), for $0\leq z_{1}\leq L,0\leq z_{2}\leq D,$
+The barotropic nonlinear SWE takes the following form (Lei et al., 2012), for $0\leq z_{1}\leq L,0\leq z_{2}\leq D,$
 $\begin{gathered}\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial z_{1}}+v\frac{\partial u}{\partial z_{2}}-fv=-g\frac{\partial h}{\partial z_{1}}+k\nabla^{2}u,\\ \frac{\partial v}{\partial t}+u\frac{\partial v}{\partial z_{1}}+v\frac{\partial v}{\partial z_{2}}+fu=-g\frac{\partial h}{\partial z_{2}}+k\nabla^{2}v,\\ \frac{\partial h}{\partial t}+u\frac{\partial h}{\partial z_{1}}+v\frac{\partial h}{\partial z_{2}}=-h\left(\frac{\partial u}{\partial z_{1}}+\frac{\partial v}{\partial z_{2}}\right)+k\nabla^{2}h,\end{gathered}$ (5.2)
-where the fluid velocity components $u$ in $z_{1}$ direction and $v$ in $z_{2}$ direction as well as the fluid depth $h$ are the model variables, $g$ is the gravity acceleration $9.8\mathrm{~{}m}\mathrm{~{}s}^{-2}$, $f$ is the Coriolis parameter defined as constant $10^{-4}\mathrm{~{}s}^{-1}$ using the $f$-plane approximation, $k$ is the diffusion coefficient which was specified as $5\times 10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$, and $L$ and $D$ are the domain limits which were set to $500\mathrm{~{}km}$ and $300\mathrm{~{}km}$, respectively.
-The SWE model was discretized with a uniform grid spacing of $10\mathrm{~{}km}$ in the $z_{1}$ and $z_{2}$ directions and was solved using the Lax-Wendroff method with a time step of $30$ seconds.
-Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ having dimension $p=3p^{\prime}=4650$.
+where the fluid velocity components $u$ in $z_{1}$ direction and $v$ in $z_{2}$ direction as well as the fluid depth $h$ are the model variables, $g$ is the gravity acceleration $9.8\mathrm{~{}m}\mathrm{~{}s}^{-2}$, $f$ is the Coriolis parameter defined as constant $10^{-4}\mathrm{~{}s}^{-1}$, $k$ is the diffusion coefficient specified as $5\times 10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$, and $L$ and $D$ are the domain limits set as $500\mathrm{~{}km}$ and $300\mathrm{~{}km}$, respectively.
+The SWE was discretized with a uniform grid spacing of $10\mathrm{~{}km}$ in the $z_{1}$ and $z_{2}$ directions and solved by the Lax-Wendroff method with a time step of $30$ seconds.
+Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ of dimension $p=3p^{\prime}=4650$.
 The initial height (depth) field was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~{}m},5.5\mathrm{~{}m}$ and $3.325\mathrm{~{}m}$, respectively.
@@
 The SWE was simulated for $48$ hours with a total of 5760 time steps at 30 second intervals after a half hour (60-steps) pre-running.
-The observations of $u,v$ and $h$ were made at $q^{\prime}=310$ grid points from $10$ randomly selected rows in $z_{1}$ direction every 3 hours, namely 360 time points.
-The whole observation vector was $\mathbf{y}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ with dimension $q=3q^{\prime}=930$, which means the observation data were quite sparse with $q=0.2p$.
+The observations of $u,v$ and $h$ were made at $q^{\prime}=310$ grid points from $10$ randomly selected rows in each direction every 3 hours, making the whole observation vector $\mathbf{y}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ be of dimension $q=3q^{\prime}=930$, which means the observations were quite sparse with $q=0.2p$.
 To generate the imperfect models, the diffusion coefficient $k$ was also considered as $10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$ in the assimilation process.
-The observation error variances were specified as $\sigma_{\mathrm{u}}^{2}=\sigma_{\mathrm{v}}^{2}=0.5,\sigma_{\mathrm{h}}^{2}=1.0$ for the three dimensions and the observation error covariance matrix was $\mathbf{R}_{t}=diag(\sigma_{\mathrm{u}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{v}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{h}}^{2}\mathbf{I}_{p^{\prime}})$.
+The observation error variances were $\sigma_{\mathrm{u}}^{2}=\sigma_{\mathrm{v}}^{2}=0.5,\sigma_{\mathrm{h}}^{2}=1.0$ for the three dimensions and $\mathbf{R}_{t}=diag(\sigma_{\mathrm{u}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{v}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{h}}^{2}\mathbf{I}_{p^{\prime}})$.
 The ensemble size $n$ was 100, and the initial ensemble was generated by adding random noises from $\mathrm{N}_{p}(0,\mathbf{I}_{p})$ to the initial state $\mathbf{x}_{1}$.
@@
 threshold 0.28 (0.26, 0.30) 0.33 (0.31, 0.35) 0.44 (0.42, 0.46) 0.48 (0.46, 0.50)
-Figure 3 reports the RMSEs at each assimilation time step of the five EnKF methods ( the standard EnKF, the inflation and the three high dimensional EnKF) to the oracle analysis states $\mathbf{x}_{t}^{a}$ and the true states $\mathbf{x}_{t}$ under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models.
-From Figure 3, the RMSEs at each assimilation time step of the three proposed high dimensional EnKF schemes became lower and lower along the time for both the correct and misspecified models.
-While the RMSEs of the standard EnKF increased over time and the inflation method also saw a slight increase in the RMSE for the true model $k=5\times 10^{4}$.
-Figure 3 shows that the three proposed high dimensional EnKF methods outperformed the inflation method for both the correct and imperfect models with much smaller RMSEs after the system settled down.
-The three proposed EnKF schemes saw much larger decreases in assimilation RMSEs in comparison to the standard EnKF and the inflation methods and the differences became larger over time, especially under the misspecified model.
-Among the three high dimensional EnKF schemes, the one utilizing the banding estimators again showed superior performance as in the simulations of the L96 model.
-Table 3 summarizes the average analysis RMSEs to the oracle analysis state and true state over the last 24 hours, together with their $25\%$ and $75\%$ quantiles, by the standard EnKF, the inflation and the three high dimensional EnKF methods for both the correct and misspecified SWE models with $p=4650,q=930$.
-The results of Table 3 reveal that the high dimensional EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands when compared to the standard EnKF and the inflation-based EnKF.
-These results were particularly evident in cases where the model was imperfect.
-Specifically, the proposed high dimensional EnKF schemes were able to reduce the RMSE to the oracle analysis states by 41% in comparison with the inflation method.
-This provides evidence to support the notion that the high dimensional EnKF methods could attain more accurate and robust assimilated outcomes for SWE models even when the observations were extremely sparse and the state variables were ultra-high dimensional.
-This also suggests that the proposed high dimensional EnKF methods are highly appealing, especially considering that geophysical models in earth science often possess an ultra-high dimension.
+Figure 3 reports the RMSEs at each assimilation time step of the five EnKF methods (the standard EnKF, inflation and three HD-EnKF) to the oracle analysis states $\mathbf{x}_{t}^{a}$ and the true states $\mathbf{x}_{t}$ under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models while Table 2 summarizes the average analysis RMSEs over the last 24 hours (2880 steps), together with their $25\%$ and $75\%$ quantiles.
+Figure 3 shows the three proposed HD-EnKF methods outperformed the inflation and the standard EnKF methods for both the correct and imperfect models with much smaller RMSEs after the system settled down and that the superiority became more apparent over time, especially under the misspecified model.
+Specifically, the RMSEs at each assimilation time step of the three proposed HD-EnKF schemes became lower and lower over time compared to increased RMSEs for the standard EnKF for all cases and slight increases RMSEs for the inflation method under true models.
+It is obvious from Table 2 that the HD-EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands with more evident superiority under imperfect models.
+Specifically, the proposed HD-EnKFs were able to reduce the RMSE by 41% compared to the inflation method, implying that the proposed methods could attain more accurate and robust assimilated outcomes even under extremely sparse observations and ultra-high dimensional state variables.
+Thus, the proposed methods are highly appealing, especially considering that geophysical models in earth science often possess ultra-high dimensions.
 6 Sea Temperature Assimilation in Northwest Pacific
-Ocean is a complex nonlinear system containing motions in various spatial and temporal scales.
-The multiscale dynamic processes of the ocean and its cross-scale and cross-sphere interactions determine the material and energy cycles of the ocean, which profoundly affect the ocean’s heat absorption and carbon storage capacity, and regulate the evolution of global climate and marine ecosystems.
-In this section, we demonstrate how the proposed method is applied to sea temperature assimilation based on the oceanography model named Regional Oceanic Modeling System (ROMS) with satellite data.
-Satellites provide a valuable tool for environmental and ocean monitoring.
+This section demonstrates how the proposed method is applied to sea temperature assimilation based on the oceanography model named Regional Oceanic Modeling System (ROMS) with satellite data.
+ROMS is a three-dimensional, free-surface, terrain-following ocean general circulation model that solves the Reynolds-averaged Navier–Stokes equations using the hydrostatic and Boussinesq assumptions (Shchepetkin and McWilliams, 2005; Haidvogel et al., 2008).
 While the satellites produce high-resolution sea surface temperature data, they can not provide ocean measurement beneath the ocean surface where the physical model ROMS plays a key role.
-ROMS is a three-dimensional, free-surface, terrain-following ocean general circulation model that solves the Reynolds-averaged Navier–Stokes equations using the hydrostatic and Boussinesq assumptions (Haidvogel et al., 2000; Shchepetkin and McWilliams, 2005; Haidvogel et al., 2008).
-Terrain-following vertical coordinates are employed which allow for greater vertical resolution in regions with complex bathymetry.
 The data assimilation was conducted from January 1st, 2008 to March 31st, 2008 over a study area 25.5N-34.7N, 144.4E-154E.
-The spatial resolution of the ROMS was $0.56^{\circ}\times 0.49^{\circ}$ with $18\times 20$ horizontal grids and $18$ uneven vertical layers over the water column.
-These made the dimension $p$ for the state variable of sea temperature ($\mathbf{x}_{t}$) to be $6480=18\times 20\times 18$.
-To scale down the computation, we did not include the other ocean variables namely salinity, horizontal velocity and vertical velocity of ocean current as the assimilated variables, and let them be internally determined by ROMS at each time step with the sea temperature as the assimilated variables.
-In our study, the integration step (finer time discretization step) for ROMS was 5 minutes and the assimilation time length was daily which was also the frequency of the observations.
-The observational data were the daily Optimum Interpolation Sea Surface Temperature (OISST) from satellites provided by National Oceanic and Atmospheric Administration (NOAA), which can be downloaded at https://www.ncei.noaa.gov/products/optimum-interpolation-sst.
-OISST is a long-term data records for Sea Surface Temperature (SST) that incorporates observations from satellites, ships and buoys into regular global grids, where ships and buoys observations are used to correct satellite SST biases.
-Descriptions for OISST data can be found in (Reynolds et al., 2007; Banzon and Reynolds, 2013).
+The spatial resolution of the ROMS was $0.56^{\circ}\times 0.49^{\circ}$ with $18\times 20$ horizontal grids and $18$ uneven vertical layers over the water column, making the dimension $p$ for the state variable of sea temperature ($\mathbf{x}_{t}$) to be $6480=18\times 20\times 18$.
+To scale down the computation, we did not include the other ocean variables namely salinity and current velocity as the assimilated variables, and let them be internally determined by ROMS at each step with the sea temperature as the assimilated variables.
+The assimilation time length was daily which was also the frequency of the observations.
+The satellite data were the daily Optimum Interpolation Sea Surface Temperature (OISST; https://www.ncei.noaa.gov/products).
+OISST is a long-term data records for Sea Surface Temperature (SST) that incorporate satellite observations into regular global grids using ships and buoys observations correct satellite SST biases (Banzon and Reynolds, 2013).
 As the OISST only provides sea surface temperature, the observation dimension was $q=360$.
 To run ROMS, we need to input the initial state of sea temperature $\mathbf{x}_{1}$, the boundary conditions and smoothed bathymetry data (sea bottom topography).
-The bathymetry data came from ETOPO2, which is a global map of the earth’s seafloor and land elevations based on satellite altimetry, shipboard echo-sounding, and other data sources and can be downloaded from https://www.ncei.noaa.gov/products/etopo-global-relief-model, as shown in Figure S.1 of the SM.
-The initial state ($\mathbf{x}_{1}$) was from the Global Ocean Reanalysis and Simulations (GLORYS) (https://data.marine.copernicus.eu/products) on January 1st.
-The boundary conditions of ROMS can be found in Marchesiello, McWilliams and Shchepetkin (2001).
+The bathymetry data came from ETOPO2 (https://www.ncei.noaa.gov/products/etopo-global-relief-model), which is a global map of the earth’s seafloor and land elevations as shown in Figure S.1 of the SM.
+The initial state ($\mathbf{x}_{1}$) was from the Global Ocean Reanalysis and Simulations (GLORYS; https://data.marine.copernicus.eu/products) on January 1st.
+Thorough descriptions of boundary conditions of ROMS can be found in Marchesiello et al.
+(2001).
 As the ocean temperature, salinity and current fields have different spatial scales and resolution in the three dimension study region both in the ROMS and the physical ocean, we transform the study region to a scale down space to facilitate the bandwidth selection in the banding estimator.
-Specifically, the grid points in the study region are denoted by coordinates $\{(u,v,h),u=1,2,\cdots 18,v=1,2,\cdots 20,h=1,2,\cdots 18\}$ in the transformed space of unit rectangles.
-The banding estimator is performed in the standardized grids.
-Specifically, the temperatures on the physical grids are labelled as $\{x_{u,v,h},u=1,2,\cdots 18,v=1,2,\cdots 20,h=1,2,\cdots 18\}$ in the standardized grids and the whole state vector in the physical grids can be labelled by $\mathbf{x}=(x_{1,1,1},\cdots,x_{18,20,18})^{\mathrm{T}}$ with the subscript $t$ hidden.
+Specifically, the grid points in the study region are denoted by coordinates $\{(u,v,h),u=1,\cdots 18,v=1,\cdots 20,h=1,\cdots 18\}$ in the transformed space of unit rectangles, where the banding estimator is performed.
+The temperatures on the physical grids are labelled as $\{x_{u,v,h}\}$ in the standardized grids and the whole state vector can be labelled by $\mathbf{x}=(x_{1,1,1},\cdots,x_{18,20,18})^{\mathrm{T}}$ with the subscript $t$ hidden.
 The banding estimator with bandwidth $k$ means that we keep the covariances between $x_{u,v,h}$ and $x_{u^{\prime},{v^{\prime}},h^{\prime}}$ on the standardized grids if they satisfy $\max\{|u^{\prime}-u|,|v^{\prime}-v|,|h^{\prime}-h|\}\leq k$.
-Considering that ocean is comparatively stable over a short period and to reduce the computation cost, we selected the bandwidth $k$ in the banding estimator by minimizing the out-of-sample RMSE over the first month of the study period between assimilated temperatures and temperatures recorded by Argo in different depths for four candidate levels $k\in\{1,2,3,4\}$.
-It was found that the RMSE increased with $k$ and was the smallest at $k=1$, which led to the use of $k=1$ as the bandwidth.
-The inflation factor for the inflation method was similarly selected by minimizing the out-of-sample RMSE over the first month which was $1.1$.
-We performed the high dimensional EnKF scheme using the banding estimator with an ensemble size of $n=30$ to obtain the assimilated three-dimensional sea temperature for the study period and region.
+We performed the HD-EnKF using the banding estimator with an ensemble size of $n=30$ to obtain the assimilated three-dimensional sea temperature for the study period and region.
 We also compared the proposed method with the standard EnKF and the inflation method.
-The out-of-sample performances of these methods were assessed on data collected under the ocean Argo project over the study period in the study region, which can be accessed at https://argo.ucsd.edu/data/.
-The Argo project collects information from inside the ocean using floats that initially drift with the ocean currents and then surface up and down between the surface and a mid-sea to collect measurements.
-As the observations provided by Argo are in-situ observations at various depths, they are more precise than the remote sensing data by the satellites, and hence, are used as substitutes for the ground truth.
-Figure 4 displays both the in-sample RMSEs of the assimilated sea surface temperature against the OISST observations, and the out-of-sample RMSEs of the assimilated sea water temperature against the temperatures recorded by Argos at various depths.
-It also presents the correlations (Corrs) between the assimilated sea surface temperatures and those from OISST, and the Corrs between the assimilated temperatures with those recorded by Argo floats.
-Figure 4 shows that the assimilation RMSEs to the OISST data of the high dimensional EnKF kept stably at a much lower level during the whole assimilation period compared with those for the standard EnKF and the inflation methods.
-At the same time, the correlations between the OISST and the assimilated sea surface temperatures by the high dimensional EnKF were much higher and more stable than the two other methods which fluctuated quite substantially over the study period.
-On the out-of-sample Argo data, the high dimensional EnKF with the banding method produced assimilated sea temperatures which were closer to Argo data as the assimilation RMSEs were slightly smaller and the correlations to Argo data were slightly higher for most days during the assimilation period.
-These implies that the proposed high dimensional EnKF algorithm could achieve more accurate and robust assimilation results in practice.
-Table 4 summarizes the average assimilation RMSEs and the correlations of the assimilated sea temperatures $\hat{\mathbf{x}}_{t}^{a}$ to the OISST data and the Argo data over the whole assimilation period by different methods together with their standard errors (in the parentheses).
-Obviously, the proposed high dimensional EnKF with banding estimator outperformed the standard EnKF and the inflation methods with much smaller RMSEs and much higher correlations to both the OISST and the Argo data.
-The reductions in the RMSEs and the increases in the correlations against both the OISST and the Argo data of the proposed high dimensional EnKF algorithm were significant compared with the standard EnKF and the inflation methods and they were much more apparent against the OISST data, suggesting that the high dimensional EnKF could make better use of the observation information.
-Besides, the standard errors of the RMSEs and correlations of the high dimensional EnKF were also smaller compared to the standard EnKF and the inflation methods, which was apparent from Figure 4.
-This implies that high dimensional EnKF provided more robust assimilation results.
-Figure 4: Assimilation RMSEs and correlations to the OISST data (left panels) and the Argo data (right panels) from Jan.
+The out-of-sample performances of these methods were assessed on data collected under the ocean Argo project over the study period and region, accessed at https://argo.ucsd.edu/data/.
+The Argo project collects information from inside the ocean using floats that initially drift with the ocean currents and then surface up and down between the surface and mid-sea to collect measurements.
+As the observations provided by Argo are in-situ observations at various depths, they are more precise than the remote sensing data by the satellites, and hence, are used as substitutes for the ground truth for the out-of-sample assessment.
+The bandwidth $k$ used in the banding estimator was $k=1$, selected by minimizing the out-of-sample RMSE over the first month of the study period between assimilated temperatures and temperatures recorded by Argo at different depths.
+The inflation factor for the inflation method was similarly selected, which was $1.1$.
+Figure 4: Daily assimilation RMSEs and correlations to the OISST(left panels) and the Argo data (right panels) from Jan.
 1st 2008 to Mar.
-31st 2008 using the standard EnKF (blue), the modified EnKF methods: inflation (red) and banding (green), respectively.
-The dashed line represents the average assimilation RMSEs and correlations to the OISST data and the Argo data over the whole assimilation period.
+31st 2008 using the standard EnKF (blue), inflation (red) and HD-EnKF with banding (green), respectively.
+The dashed horizontal line represents the average assimilation RMSEs and correlations over the whole assimilation period.
 Table 3: The average assimilation RMSEs and correlations and their standard errors (in the parentheses) to the OISST data and the Argo data over the whole assimilation period.
@@
 Figure 5: The spatial distribution of the assimilated sea surface temperature on January 1st, February 15th and March 31st in 2008 over the study area, with the fill color from red to blue representing sea surface temperature from high to low.
-Figure 5 depicts the spatial distribution of the assimilated sea surface temperature over the study area on January 1st, February 15th and March 31st in 2008.
-From the picture, we can see the spatial distribution of the assimilated sea surface temperature varied over time and that there was an overall downward trend in the study region during the study period.
+Figure 4 displays both the in-sample RMSEs and correlations between the assimilated sea surface temperature and the OISST observations, and the out-of-sample RMSEs and correlation between the assimilated sea water temperature and the temperatures recorded by Argos at various depths.
+From Figure 4, the HD-EnKF provided much lower and more stable assimilation RMSEs and much higher correlations in the entire assimilation period than those for the standard EnKF and the inflation methods.
+On the out-of-sample performance on the Argo data, the HD-EnKF produced assimilated sea temperatures with smaller RMSEs and higher correlations both on average and for most days during the assimilation period.
+These imply that the proposed HD-EnKF algorithm could offer more accurate and robust assimilation results in practice.
+Detailed numerical results of these methods are reported in Table 3, which summarizes the average assimilation RMSEs and correlations of the assimilated $\hat{\mathbf{x}}_{t}^{a}$ to the OISST and the Argo data over the whole assimilation period.
+The spatial distribution of the assimilated sea surface temperature over the study area on January 1st, February 15th and March 31st in 2008 is depicted in Figure 5, which varied over time and saw an overall downward trend during the study period.
 7 Discussion
-In this paper, we propose high dimensional EnKF algorithms to suit the needs of high resolution data assimilation.
-We take the opportunity to study the theoretical properties of the EnKF under both the fixed and the high dimensional Gaussian state-space model (2.8).
-The proposed high dimensional EnKF methods and algorithms can provide consistent data assimilation results in the high dimensional setting.
-The investigation of the high dimensional EnKF when the physical model $\mathcal{M}_{t}$ is misspecified gives a description of the performance of the EnKF under misspecified models.
-The study shows that the proposed high dimensional EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations.
-One is that the established theoretical analysis on the consistency and the error bounds makes the proposed high dimensional methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies and the empirical study on sea temperature assimilation.
+In this paper, we propose HD-EnKF algorithms to suit the needs of high dimensional data assimilation.
+We take the opportunity to study the theoretical properties of the EnKF under both the fixed and the high dimensional cases for both the correct and misspecified models.
+The study shows that the proposed HD-EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations.
+One is that the established theoretical analysis on the consistency and the error bounds makes the proposed HD-EnKF methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies and the empirical study on sea temperature assimilation.
 The latter means that a smaller ensemble size for the EnKF can be used at a given level of accuracy which is particularly attractive as the running of the geophysical models is generally very expensive.
-The proposed high dimensional EnKF offers significant potential for application in earth science where the geophysical models, the state variables or the observations are ultra-high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.
-{acks}
-[Acknowledgments] The authors would like to thank Professor Xiaogu Zheng for insightful discussions and comments.
-{funding}
-The research was partially supported by a National Natural Science Foundation of China grant Nos.
-12292980 and 12292983.
-This research was also partially supported by the National Key Scientific and Technological Infrastructure project “Earth System Science Numerical Simulator Facility” (EarthLab).
-[…]
-[…]
+The proposed HD-EnKF has great potential for applications in earth science where the geophysical models, the state variables or the observations are ultra-high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.
+SUPPLEMENTARY MATERIAL
+We provide detailed steps of the proposed HD-EnKF algorithm, bandwidth selection of the mid-banding estimator, additional theoretical results and simulation results, and the details for the proofs of the theoretical results in the Supplementary.

### v3

@@
-aoas/sm_enkf
+High Dimensional Ensemble Kalman Filter
 Abstract
-The ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of earth science, engineering and beyond.
-However, several theoretical aspects of the EnKF remain unknown, especially when the state variable is of high dimensional accompanied with high resolution observation and physical models.
-This paper first proposes several high dimensional EnKF (HD-EnKF) methods that provide consistent estimators for the important forecast error covariance and the Kalman gain matrix.
-It then studies the theoretical properties of the EnKF under both the fixed and high dimensional state variables, which provides the mean square errors of the analysis states to the underlying oracle states offered by the Kalman filter and gives the much needed insight into the roles played by the forecast error covariance on the accuracy of the EnKF.
+The Ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of the sciences and engineering.
+When the state variable is of high dimensional accompanied with high resolution observations of physical models, some key theoretical aspects of the EnKF are open for investigation.
+This paper proposes several high dimensional EnKF (HD-EnKF) methods equipped with consistent estimators for the important forecast error covariance and Kalman Gain matrices.
+It then studies the theoretical properties of the EnKF under both fixed and high dimensional state variables, which provides one-step and multiple-step mean square errors of the analysis states to the underlying oracle states offered by the Kalman Filter and gives the much needed insight to the roles played by the forecast error covariance on the accuracy of the EnKF.
 The accuracy of the data assimilation under the misspecified physical model is also considered.
-Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed HD-EnKF algorithms perform better than the standard EnKF methods as they provide more robust and accurate assimilated results.
-The HD-EnKF was applied to assimilate sea temperature in the Northwest Pacific, which showed more accurate out-sample performances than the existing methods.
+Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed HD-EnKF algorithms outperform the standard EnKF and widely used inflation methods.
 Keywords: Data Assimilation; Dynamic System; Ensemble Kalman Filter; High Dimensional Imputation.
@@
 For chaotic systems sensitive to initial values, data assimilation can significantly improve the accuracy of the forecast by using optimized state variables as initial values (Lorenz, 1963).
-Although much applied in numerical weather and oceanography prediction, Kalman Filter and its later developments (Ensemble Kalman Filter, three and four-dimensional variational assimilation (3D-Var, 4D-Var; Pailleux, 1990; Courtier et al., 1998)) are the methodological core for scientific algorithms over a range of fields, including hydrology, remote sensing data utilization, carbon data assimilation, robotic control and automatic driving.
+Although much applied in numerical weather and oceanography prediction, Kalman Filter and its later developments (EnKF, three and four-dimensional variational assimilation (3D-Var, 4D-Var; Pailleux, 1990; Courtier et al., 1998)) are the methodological core for scientific algorithms over various fields, including hydrology, remote sensing data inversion, carbon data assimilation, robotic control and automatic driving.
 Data assimilation methods also provide the underlying algorithms and predictive techniques for building digital twins for social and economic systems.
@@
 Filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation.
-High resolution data assimilation and forecasting techniques are much needed in modern meteorology and atmospheric systems like global carbon calculation and ocean data assimilation, which leads to high dimensional geophysical models with high dimensional state vectors.
-Besides, along with advances in remote sensing technology in measuring the earth system, there is a growing availability of high spatial resolution satellite observations (Stroud et al., 2010; Guinehut et al., 2012).
+One reason for the filter divergence is the variance of the ensemble forecast error becomes too small when there are unaccounted model errors (Lorenc, 2003); the other is due to the substantial mismatch between the estimated and true error covariance matrices (Houtekamer and Mitchell, 1998; Constantinescu et al., 2007).
+Along with advances in remote sensing technology in measuring the Earth system, there is a growing availability of high spatio-temporal resolution satellite observations (Stroud et al., 2010; Guinehut et al., 2012; Lellouche et al., 2013).
+Meanwhile, the high resolution dataset and forecasting are in increasing demands in modern meteorology and atmospheric systems like high resolution weather forecasting, global carbon calculation and ocean data assimilation, which leads to high dimensional geophysical models with high dimensional state vectors.
 The high dimensional state vectors lead to high dimensional  forecast error covariance and Kalman Gain matrices.
-The high dimensionality is known to cause the non-convergence of the sample forecast error covariance (Bai and Yin, 1993), which is felt in the form of filter divergence and unreliable data assimilation in practice.
-Geophysicists have proposed two methods to counter the problems.
-One is the localization method (Furrer et al., 2006; Furrer and Bengtsson, 2007) that sets covariances to zero if the geographical distances are bigger than a threshold; and the other is the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term (Wang and Bishop, 2003; Liang et al., 2012; Constantinescu et al., 2007).
-However, the statistical properties of these EnKF methods are generally lacking, despite their widespread applications.
+The high dimensionality is known to cause the non-convergence of the sample forecast error covariance (Bai and Yin, 1993), felt in the form of filter divergence and unreliable data assimilation in practice.
+Geophysicists have proposed two methods to counter the problems, the localization method (Furrer et al., 2006; Furrer and Bengtsson, 2007) that sets covariances to zero for the geographical distances bigger than a threshold and the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term (Wang and Bishop, 2003; Constantinescu et al., 2007; Liang et al., 2012).
+However, the statistical properties of these methods are yet to be established.
+Tong (2018) investigated the performance of the local EnKF with domain localization under linear systems.
 Katzfuss et al.
-(2020) proposed Gibbs Ensemble Kalman Filter and Smoother, and particle Ensemble Kalman Filter and Smoother algorithms for high dimensional hierarchical dynamic models from a Bayesian perspective.
-This paper proposes several high dimensional EnKF (HD-EnKF) methods based on consistent high dimensional covariance estimators and conducts a theoretical analysis of assimilated state vectors.
-It is shown that consistent estimation of the forecast error covariance leads to the consistency of the Kalman gain matrix by establishing an inequality that bounds the error of estimation of the Kalman Gain matrix by that of the forecast error covariance.
-This implies that, even at the high dimensional setting, the mean square errors (MSEs) of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators.
-Thus, consistent recovery of the underlying analysis states can be attained by the proposed HD-EnKF procedure.
-We establish the mean square error bounds and weak convergence of the analysis states from the HD-EnKF to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models.
-The MSE bounds for the multiple-step forecasting and data assimilation are also provided.
-Simulation studies suggest that the proposed HD-EnKF algorithm could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method.
-We apply the proposed HD-EnKF to assimilate sea temperature by combining a physical ocean model with satellite observations and use the ARGO data to validate the quality of the assimilated sea temperature.
-Both the simulation and the empirical study showed advantages of the HD-EnKF approach.
+(2020) proposed Gibbs EnKF and particle EnKF algorithms for high dimensional hierarchical dynamic models from a Bayesian perspective.
+Al-Ghattas and Sanz-Alonso (2024) developed a non-asymptotic analysis on the difference between the ensembles of single-step Ensemble Kalman Inversion using the true and the localized estimation of forecast error covariances under the framework of effective dimension of matrices and Al-Ghattas et al.
+(2024) explored the resampling EnKF under linear systems and effective dimension.
+While Al-Ghattas et al.
+(2025) studied the difference between the ensembles of single-step EnKF using the true and the thresholded estimation of forecast error covariance functions in a small correlation lengthscale regime.
+The EnKF methods and their theoretical analysis under high dimensional framework for general non-linear models, especially the theoretical performance of the multiple-step EnKF under both the correct and misspecified models, need further investigation.
+This paper proposes several high dimensional EnKF (HD-EnKF) methods and conducts a theoretical analysis of the one-step and multiple-step assimilated state vectors for general nonlinear models and imperfect models under both fixed and high dimensional cases.
+It is shown that even at the high dimensional setting, the mean square errors (MSEs) of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators.
+Thus, consistent recovery of the underlying analysis states can be attained by the proposed HD-EnKF.
+Specifically, we establish the MSE bounds and weak convergence of the analysis states from the HD-EnKF to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models.
+The MSE bounds for the multiple-step forecast and assimilation are also provided.
+The investigation of the HD-EnKF under the misspecified model gives a more realistic description of the performance of the EnKF.
+Under the misspecified model, the MSE not only depends on the estimation error of the forecast error covariance matrices but also the discrepancy between the misspecified and the true model.
+Simulation studies suggest that the proposed HD-EnKF could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method.
+This implies that the proposed HD-EnKF is applicable in a range of studies and prospects in earth science like global carbon calculation, weather forecasting and the production of high resolution data sets, considering both the theoretical guarantee and better performances in practice.
 The paper is structured as follows.
 Section 2 reviews the Kalman Filter and the EnKF.
-Section 3 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman gain matrix, and establishes the theoretical properties of the EnKF.
+Section 3 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman Gain, and establishes the theoretical properties of the EnKF.
 The effects of imperfect or misspecified models on the accuracy of the EnKF are provided in Section 4.
-Results of empirical data assimilation based on two well-known geophysical models are reported in Section 5 and the application to the sea temperature data assimilation is presented in Section 6, followed by a conclusion in Section 7.
-Details of the proposed HD-EnKF algorithms, additional theoretical and simulation results, and technical details are given in the Supplementary Material (SM).
+Results of empirical data assimilation based on two well-known geophysical models are reported in Section 5, followed by a conclusion in Section 6.
+The proposed HD-EnKF algorithms, additional theoretical and simulation results, and technical details are given in the Supplementary Material (SM).
 2 Review on Kalman and Ensemble Kalman Filters
 The Kalman Filter (Kalman, 1960) assumes the following linear Gaussian state-space model (which was generalized to nonlinear models in the EnKF)
-$\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathbf{M}_{t}\mathbf{x}_{t-1}+\mathbf{w}_{t},\quad\ \quad\mathbf{w}_{t}\sim\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\,\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}(\mathbf{0},\mathbf{R}_{t}),\end{array}\right.$ (2.1)
+$\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathbf{M}_{t}\mathbf{x}_{t-1}+\mathbf{w}_{t},\quad\ \quad\mathbf{w}_{t}\sim\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\,\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}(\mathbf{0},\mathbf{R}_{t}),\end{array}\right$.
+(2.1)
 where $\mathbf{x}_{t}$ is the $p$-dimensional not fully observed state vector of a dynamic system at time $t$, $\mathbf{M}_{t}$ is the linear model operator (also called evolution or forecast operator), $\mathbf{w}_{t}$ is the model error term with $\mathbf{Q}_{t}$ as its covariance, $\mathbf{y}_{t}$ is the $q$-dimensional observation vector at time $t$, $\mathbf{H}_{t}$ is a linear observation operator which relates the state $\mathbf{x}_{t}$ to the observation $\mathbf{y}_{t}$, and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix.
 Assuming $\mathbf{M}_{t}$, $\mathbf{Q}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are known, the Kalman Filter provides a two-step alternative updating framework based on the theory of multivariate normal distribution.
-In the forecast step, suppose that $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p}\left(\mathbf{x}_{t-1}^{a},\boldsymbol{\Sigma}_{t-1}^{a}\right)$, where $\mathbf{x}_{t-1}^{a}$ is the analysis state vector at the previous time $t-1$ and $\mathbf{y}_{1:t-1}$ denotes the observation $\{\mathbf{y}_{i}\}_{i=1}^{t-1}$ up to time $t-1$, the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{y}_{1:t-1}$ is
+In the forecast step, suppose that $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p}(\mathbf{x}_{t-1}^{a},\boldsymbol{\Sigma}_{t-1}^{a})$, where $\mathbf{x}_{t-1}^{a}$ is the analysis state vector at the previous time t1t-1 and $\mathbf{y}_{1:t-1}$ denotes the observation $\{\mathbf{y}_{i}\}_{i=1}^{t-1}$ up to time t1t-1, the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{y}_{1:t-1}$ is
 $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p}\big{(}\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f}\big{)}\equiv{\mathrm{N}}_{p}\big{(}{\mathbf{M}}_{t}\mathbf{x}_{t-1}^{a},{\mathbf{M}}_{t}\boldsymbol{\Sigma}_{t-1}^{a}{\mathbf{M}}_{t}^{\mathrm{T}}+\mathbf{Q}_{t}\big{)},$ (2.2)
@@
 $\boldsymbol{\Sigma}_{t}^{f}=\mathbf{M}_{t}\boldsymbol{\Sigma}_{t-1}^{a}\mathbf{M}_{t}^{\mathrm{T}}+\mathbf{Q}_{t},$ (2.3)
-is the variance of the forecast error $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$, also called the forecast error covariance matrix.
+is the forecast error covariance, the variance of the forecast error $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$.
 In the update (assimilation) step, the goal is to update $\mathbf{x}_{t}^{f}$ according to the new observation $\mathbf{y}_{t}$ and to reduce the difference between the true state $\mathbf{x}_{t}$ and the forecast state $\mathbf{x}_{t}^{f}$.
-The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is Gaussian, which implies that
-$\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right),$ (2.4)
-where, by the conditional Gaussian distribution formulae,
+The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is Gaussian, which implies that $\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right)$, where, by the conditional Gaussian distribution formulae,
 $\mathbf{x}_{t}^{a}$ $={\mathbf{x}}_{t}^{f}+\mathbf{K}_{t}\big{(}\mathbf{y}_{t}-\mathbf{H}_{t}{\mathbf{x}}_{t}^{f}\big{)},\,\boldsymbol{\Sigma}_{t}^{a}=\left(\mathbf{I}_{p}-\mathbf{K}_{t}\mathbf{H}_{t}\right)\boldsymbol{\Sigma}_{t}^{f}\text{ and }$ (2.4)
@@
 Note that the analysis state $\mathbf{x}_{t}^{a}$ is the minimum variance unbiased estimator for the mean of $\mathbf{x}_{t}$ under the linear state-space model (2.1).
-However, models in earth science or engineering are often nonlinear, that is the evolution model in (2.1) may not be linear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$, where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2).
-Besides, for large dynamic systems, the dimension $p$ of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store.
-To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error covariance is no longer updated as (2.3) but by the sample covariance based on $n$ forecast ensembles.
-It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing.
-The EnKF assumes possibly nonlinear model system with linear observation operator:
-$\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1}\right)+\mathbf{w}_{t},\quad\quad\ \mathbf{w}_{t}\sim\mathrm{N}_{p}(\mathbf{0},\mathbf{Q}_{t}),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\quad\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}\left(\mathbf{0},\mathbf{R}_{t}\right),\end{array}\right.$ (2.7)
-where $\mathcal{M}_{t}$ is the possibly nonlinear model operator, $\mathbf{H}_{t}$ is a linear observation operator.
-It further assumes that the forecast error $\boldsymbol{\eta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$ follows a Gaussian distribution given $\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$, where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)$.
-Here $\mathcal{M}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are assumed to be known for the time being as in the Kalman Filter.
+However, models in earth science or engineering are often nonlinear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$ in (2.1), where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2).
+Besides, for large dynamic systems, the dimension of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store.
+To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error covariance is no longer updated as (2.3) but by the sample covariance of $n$ forecast ensembles.
+It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing, and thus improves the computing efficiency and reduces the storage cost.
+The EnKF assumes possibly nonlinear model system with a linear observation operator, with the first equation in (2.1) replaced by
+$\mathbf{x}_{t}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1}\right)+\mathbf{w}_{t},$ (2.6)
+where $\mathcal{M}_{t}$ is the possibly nonlinear model operator.
 A key feature of the EnKF is to generate an ensemble of perturbed forecast states $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ of size $n$ from the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{y}_{1:t-1}$, which are used to estimate the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$.
@@
     Run the model forward in time to get the perturbed forecast state and their average:
-    $\mathbf{x}_{t,j}^{f}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1,j}^{a}\right)+\mathbf{w}_{t,j}\text{ and }\hat{\mathbf{x}}_{t}^{f}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f},\text{ where }\mathbf{w}_{t,j}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right).$ (2.8)
+    $\mathbf{x}_{t,j}^{f}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1,j}^{a}\right)+\mathbf{w}_{t,j},\,\hat{\mathbf{x}}_{t}^{f}=\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f},\,\mathbf{w}_{t,j}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right).$ (2.7)
   • Step (ii):
@@
     $\hat{\boldsymbol{\Sigma}}_{t}^{f}=(n-1)^{-1}\sum_{j=1}^{n}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}^{\mathrm{T}}\triangleq\mathbf{S}_{t}^{f},$ (2.8)
-    and obtain an estimate of the Kalman gain matrix $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$.
+    and obtain the Kalman Gain estimate $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$.
   • Step (iii):
     Calculate the perturbed residuals $\mathbf{d}_{t,j}=\mathbf{y}_{t}+\boldsymbol{\varepsilon}_{t,j}-{\mathbf{H}}_{t}\mathbf{x}_{t,j}^{f}$ where $\{\boldsymbol{\varepsilon}_{t,j}\}_{j=1}^{n}\stackrel{{\scriptstyle IID}}{{\sim}}{\mathrm{N}}\left(\mathbf{0},\mathbf{R}_{t}\right)$, and let $\mathbf{d}_{t}=n^{-1}\sum_{j=1}^{n}\mathbf{d}_{t,j}$.
-Compute the perturbed analysis states and their average, the assimilated (analysis) state,
+Compute the perturbed analysis states and their average, the analysis state,
     $\mathbf{x}_{t,j}^{a}$ $=\mathbf{x}_{t,j}^{f}+\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}\mathbf{d}_{t,j},$ (2.9)
@@
     which will be used for the data assimilation at the next time step.
+In practice, the algorithm is initialized at time $t=0$ by drawing $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ independently from ${\mathrm{N}}_{p}(\mathbf{x}_{0}^{a},\Sigma_{0}^{a})$ where $\mathbf{x}_{0}^{a}$ and $\Sigma_{0}^{a}$ are either chosen based on the physical model or by pre-running the physical model for a period of time.
+Then, the ensemble is propagated forward through time, alternating between the forecast and the analysis steps, according to the three-step procedure above.
+If $\mathbf{R}_{t}$ and $\mathbf{Q}_{t}$ are unknown, their estimates can be entertained to substitute them in (2.7) and (2.9).
 The EnKF does not require the expensive recursions (2.3) and (2.4)-(2.5) used in the Kalman Filter.
 Instead, it estimates the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance.
-For the case with non-linear model operator as in (2.7), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.5)-(2.6), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance defined by $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$.
-In fact, $\mathbf{x}_{t}^{a}$ defined by (2.5)-(2.6) is the solution to the minimizing problem (2.11)
-$J(\mathbf{x})=\frac{1}{2}\big{\{}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\big{\}}$ (2.11)
-with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$.
+For the case with non-linear model operator as in (2.6), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.4)-(2.5), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$.
+In fact, $\hat{\mathbf{x}}_{t}^{a}$ defined by (2.9) is the solution to the minimizing problem (2.10)
+$J(\mathbf{x})=\frac{1}{2}\big{\{}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\big{\}},$ (2.10)
+and $\mathbf{x}_{t}^{a}$ defined by (2.4)-(2.5) is the solution to the above minimizing problem with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$.
 For the ease of reference, and to avoid differentiating between the linear and the non-linear cases, we also call $\mathbf{x}_{t}^{a}$ as the analysis state of the Kalman Filter even in the non-linear model case.
@@
 3 Convergence of Ensemble Kalman Filter
-[…]
+We first provide the upper bound of the estimation error of the estimated Kalman Gain matrix.
+Throughout the paper, $\|\cdot\|$ denotes the $L_{2}$ (spectral) norm of a matrix or the $L_{2}$ norm of a vector, while other matrix norms will be specified explicitly.
+For two positive sequences $a_{n}$ and $b_{n}$, $a_{n}\asymp b_{n}$, $a_{n}\ll b_{n}$ and $a_{n}\gg b_{n}$ mean that $a_{n}/b_{n}$ is bounded away from zero and infinity, $a_{n}/b_{n}\rightarrow 0$ and $b_{n}/a_{n}\rightarrow 0$ as $n\rightarrow\infty$, respectively.
+The following assumptions are needed in the analysis.
 Assumption 1.
@@
 If model $\mathcal{M}_{t}$ is linear, it automatically satisfies both (i) and (ii) of Assumption 1.
-If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite interval.
+If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite intervals.
 However, strongly nonlinear systems may not satisfy the Lipschitz condition.
@@
 Assumption 3 puts a mild condition on the observation operator $\mathbf{H}_{t}$.
-The following proposition provides a basic result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$ is an estimator of the Kalman gain matrix assuming $\mathbf{R}_{t}$ is known.
-Proposition 3.1.
-Suppose Assumption 2 holds and $\lambda_{\min}(\mathbf{R}_{t})\geq\varepsilon_{0}$ for a positive $\varepsilon_{0}$,
+The following lemma provides a basic result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$ is an estimator of the Kalman Gain matrix assuming $\mathbf{R}_{t}$ is known.
+Closely related forms have appeared in the literature (Le Gland et al., 2009; Kwiatkowski and Mandel, 2015).
+Lemma 3.1.
+Suppose $\mathbf{R}_{t}$ is known and $\lambda_{\min}\left(\mathbf{R}_{t}\right)>0$, then
 $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|\leq\lambda_{\min}^{-1}\left(\mathbf{R}_{t}\right)\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|\leq{\lambda_{\min}^{-1}\left(\mathbf{R}_{t}\right)}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|.$
-The last inequality is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$.
-Proposition 3.1 suggests that the upper bound of the estimation error in $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$.
-Proposition 3.1 holds under weak conditions and actually is valid for both fixed and high dimensions.
+Lemma 3.1 suggests that the upper bound of the estimation error in $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$, which is valid for both fixed and high dimensions.
 Note that under fixed dimension, the sample covariance is consistent with the underlying population covariance at the $\sqrt{n}$ rate, that is $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$.
-The results for the high dimensions are shown later.
+The corresponding results for the high dimensions are shown later.
+Remark 1.
+The last inequality in Lemma 3.1 is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the conditional variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$.
+Let us take a close look at $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$.
+For the extreme case of $\mathbf{R}_{t}=\mathbf{0}$, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be 0 if $\mathbf{H}_{t}$ is invertible.
+Intuitively, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be close to 0 if $\mathbf{R}_{t}$ is close to $\mathbf{0}$.
+Suppose $p=q=1$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=(1+\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}/\sigma_{\mathbf{R}_{t}}^{2})^{-1}<1$.
+This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ decreases with the increase of $\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}$ or decrease of $\sigma_{\mathbf{R}_{t}}^{2}$.
+For $p,q>1$, if $\mathbf{H}_{t}=\mathbf{I}$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=\|\mathbf{I}-\boldsymbol{\Sigma}_{t}^{f}(\boldsymbol{\Sigma}_{t}^{f}+\mathbf{R}_{t})^{-1}\|=\varepsilon_{0}^{2}/(1+\varepsilon_{0}^{2})$ if we assume $\mathbf{R}_{t}=\varepsilon_{0}^{2}\boldsymbol{\Sigma}_{t}^{f}$.
+This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ will decrease with the decrease of $\varepsilon_{0}^{2}$, the ratio of $\mathbf{R}_{t}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
 In practice, the true observation error covariance matrix $\mathbf{R}_{t}$ may be unknown.
-Corollary S.3.1 of the SM provides an upper bound of the estimation error of the $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ when both $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
-Based on the result in Proposition 3.1, we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.5)-(2.6) by firstly establishing the convergence of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
-For the fixed dimension case, the sample covariance matrix is consistent such that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$.
-However, when the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.9) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ (Bai and Yin, 1993), which is a major threat to the validity of the EnKF method.
-Next, we present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and their covariance classes before we give the convergence of the EnKF.
-3.1 Consistent estimation of high dimensional $\boldsymbol{\Sigma}_{t}^{f}$
-The banding estimator (Bickel and Levina, 2008b) of the forecast error covariance matrix with bandwidth $k$ is $B_{k}(\mathbf{S}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|i-j|\leq k)\big{]}$, where $0\leq k<p$ and $\mathbf{S}_{t}^{f}=[s_{tij}^{f}]_{p\times p}$ is the sample covariance of the EnKF forecast ensembles.
+Let $\hat{\mathbf{R}}_{t}$ be an estimator of $\mathbf{R}_{t}$ based on either knowledge or historical data.
+The following Corollary 3.1 provides a result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{\hat{R}}_{t}\big{)}^{-1}$ is an estimator of the Kalman Gain matrix when both the forecast and observation error covariance matrices $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
+Corollary 3.1.
+If both the forecast and observation error covariance matrices $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively, then
+$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ $\leq\big{(}\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|\|\mathbf{H}_{t}\|+\|\mathbf{K}_{t}\|\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|\big{)}\|(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\hat{\mathbf{R}}_{t})^{-1}\|.$
+If $\mathbf{R}_{t}$ is invertible and $\hat{\mathbf{R}}_{t}$ is invertible with probability approaching $1$, then
+$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|\leq{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\lambda_{\min}^{-1}(\mathbf{R}_{t})\left\|\mathbf{H}_{t}\right\|\|\boldsymbol{\Sigma}_{t}^{f}\|\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|.$
+Compared with the upper bound in Lemma 3.1, there is an extra term related to the bound of the estimation error in the observation error covariance matrix $\mathbf{R}_{t}$.
+Most of the results established in the following of the paper are for the case of known $\mathbf{R}_{t}$ as is commonly assumed in the literature of the EnKF.
+Results for estimated $\mathbf{R}_{t}$ can be attained in the fashion shown in Corollary 3.1.
+Based on the result in Lemma 3.1 (or Corollary 3.1 for unknown $\mathbf{R}_{t}$), we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.4)-(2.5) by firstly establishing the convergence of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
+For the fixed or low dimensional state vectors $\mathbf{x}_{t}$ where $p$ is of a smaller order of the ensemble size $n$ such that $p/n\to 0$, the sample covariance matrix is consistent such that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$.
+However, remote sensing technology has been advancing the measurement of massive amounts of data sets for many geophysical processes, which produces high resolution observational data.
+Meanwhile, the increasing demands of high resolution forecast and dataset also make the spatiotemporal resolution of the state variables become much higher.
+Both make $\mathbf{x}_{t}$ high dimensional and bring high dimensional challenges on the validity of the EnKF.
+When the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.8) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ (Bai and Yin, 1993), which is a major threat to the validity of the EnKF.
+The last two decades have witnessed rapid development in high dimensional statistical inference.
+Consistent high dimensional  covariance estimators have been proposed, which include the banding, tapering and thresholding estimators (Bickel and Levina, 2008b, a; Cai et al., 2010).
+Next, we present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and their covariance classes before we evaluate the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ from the EnKF to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.
+3.1 Consistent estimation of high dimensional forecast error covariance matrix
+The banding estimator (Bickel and Levina, 2008b) of the forecast error covariance matrix with bandwidth $k$ is
+$B_{k}(\mathbf{S}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|i-j|\leq k)\big{]}_{p\times p},\,\,0\leq k<p,$
+where $\mathbf{S}_{t}^{f}=[s_{tij}^{f}]_{p\times p}$ is the sample covariance of the EnKF forecast ensembles.
 The covariance matrix class suitable for the banding estimator is the following bandable covariance class, $\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}$ for some positive $\alpha$, $C$ and $\varepsilon_{0}$.
@@
 Cai et al.
-(2010) updated the banding estimator with a tapering alteration which tapers down the level of the thresholding gradually
-$\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)=\mathbf{S}_{t}^{f}\circ\mathbf{W}(k)=\big{(}w_{ij}(k)s_{tij}^{f}\big{)}_{p\times p},$
-where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\left\{(k-|i-j|)_{+}-\left(k/2-|i-j|\right)_{+}\right\},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$.
+(2010) updated it with a tapering alteration which tapers down the level of the thresholding gradually
+$\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)=\mathbf{S}_{t}^{f}\circ\mathbf{W}(k)=\big{[}w_{ij}(k)s_{tij}^{f}\big{]}_{p\times p},$
+where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\big{\{}(k-|i-j|)_{+}-(k/2-|i-j|)_{+}\big{\}},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$.
 (a) $h(k)$ of $\boldsymbol{\Sigma}\in\mathcal{U}_{b}$
@@
 (d) Heatmap of $\boldsymbol{\Sigma}\in\mathcal{U}_{bc}$
-Figure 1: Bandable and circular bandable covariance matrices: the average energy function $h(k)$ in Panels (a) and (b) and their heatmaps in Panels (c) and (d), where $h(k)$ is the average of the squares of the $k^{th}$ sub-diagonal entries.
+Figure 1: Bandable $\mathcal{U}_{b}$ and circular bandable $\mathcal{U}_{bc}$ covariance matrices: the average energy function $h(k)$ in Panels (a) and (b) and their heatmaps in Panels (c) and (d), where $h(k)$ is the average of the squares of the $k^{th}$ sub-diagonal entries.
 In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or its poles, the state vectors are highly dependent for the variable components between the two ends of the state vector.
-Thus we propose a circular form of the bandable structure, $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$ Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariance matrices that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$, where $h(k)$ is the average of the squares of the $k^{th}$ subdiagonal entries of a covariance defined by $h(k):=\frac{1}{2(p-k)}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=\frac{1}{p-k}\sum_{l=1}^{p-k}\sigma_{ll+k}^{2}.$ Figures 1 (a)-(b) show that the $h(k)$ function of covariances in $\mathcal{U}_{b}$ is a $L$-type curve as the subdiagonal index $k$ increases, while it is an $U$-type curve for covariances in $\mathcal{U}_{bc}$.
-For covariances in the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, the proposed mid-banding forecast error covariance estimator with bandwidths $k_{1}$ and $k_{2}$ is $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})=[s_{tij}^{f}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))]$, where $0\leq k_{1}<(p-k_{2})\leq p$.
+Thus we propose a circular form of the bandable structure, $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$ Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariances that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$, where $h(k)$ is the average of the squares of the $k^{th}$ subdiagonal entries of a covariance defined by $h(k):=0.5{(p-k)^{-1}}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=(p-k)^{-1}\sum_{l=1}^{p-k}\sigma_{l,l+k}^{2}.$ Figures 1 (a)-(b) show that the $h(k)$ functions of covariances in $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$ are $L$-type and $U$-type curves as the subdiagonal index $k$ increases, respectively.
+For covariances in the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, the proposed mid-banding forecast error covariance estimator with bandwidths $k_{1}$ and $k_{2}$ is
+$B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})=[s_{tij}^{f}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))]_{p\times p},\,\,0\leq k_{1}<(p-k_{2})\leq p.$
 The bandable matrix class requires that the state vector $\mathbf{x}_{t}$ has an ordering (or a permutation) such that the correlation decays as two components of $\mathbf{x}_{t}$ are further apart.
 However, the presence of geo-spatial processes (weather front or hurricane) and geography (mountains or ocean) will interrupt the spatial dependence ordering prescribed by the banding covariance class.
-In such situations, the thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008a) may be employed, where $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i\text{ and }\lambda_{\min}(\boldsymbol{\Sigma})\geq\varepsilon_{0}>0\big{\}},\text{ for }0\leq\gamma<1.$ The thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|s_{tij}^{f}|\geq s)\big{]}$.
+In such situations, the thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008a) may be employed, where $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i\text{ and }\lambda_{\min}(\boldsymbol{\Sigma})\geq\varepsilon_{0}>0\big{\}},\text{ for }0\leq\gamma<1.$ The thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|s_{tij}^{f}|\geq s)\big{]}_{p\times p}$.
 Note that all of the banding, mid-banding, tapering and thresholding estimators depend on either the bandwidth $k$ or the threshold $s$ which can be selected by methods in Qiu and Chen (2015).
 For the bandwidths $k_{1},k_{2}$ of the mid-banding estimator, we propose a similar selection method and details can be found in Section S.2 of the SM.
+To apply the high dimensional estimators of the forecasting error covariance $\boldsymbol{\Sigma}_{t}^{f}$, we need the following distributional assumption.
+Assumption 4.
+The random vectors $\mathbf{x}_{t,i}^{f}=(x_{t,i1}^{f},\cdots,x_{t,ip}^{f})^{\mathrm{T}}$ are IID and each $x_{t,1j}^{f}\stackrel{{\scriptstyle d}}{{\sim}}F_{t,j}$ and $(x_{t,1j}^{f})^{2}\stackrel{{\scriptstyle d}}{{\sim}}G_{t,j}$, where $G_{t,j}(u)=F_{t,j}(\sqrt{u})-F_{t,j}(-\sqrt{u})$ for $u>0$.
+There exists a $\lambda_{0}>0$ such that $\max_{1\leq j\leq p}\mathbb{E}(e^{\lambda(x_{t,1j}^{f})^{2}})<\infty\textrm{ for }\lambda\in(-\lambda_{0},\lambda_{0}).$
+Assumption 4 effectively assume $(x_{t,1j}^{f})^{2}$ are sub-exponentially distributed, which include the Sub-Gaussian random variables as a special case.
 The following Proposition 3.1 summarizes the convergence rates of the forecast error covariance estimators by the banding, mid-banding, tapering and thresholding methods.
 Proposition 3.1.
-Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally IID Gaussian or sub-exponentially distributed as in Assumption S.3.1 of the SM.
+Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally sub-exponentially distributed as in Assumption 4.
   • (i)
-    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
+    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2\alpha+2)}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
   • (ii)
-    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
+    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2\alpha+2)}$, uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
   • (iii)
-    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$, suppose $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are sub-Gaussian, then for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$, $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
+    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$ and $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$, the optimal rate of convergence of the tapering estimator is $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
   • (iv)
-    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$, $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}.$
+    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$, $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)(n^{-1}{\log p})^{(1-\gamma)}.$
 The results in Proposition 3.1 (i), (iii) and (iv) are, respectively, from Bickel and Levina (2008b), Cai et al.
@@
 They imply that the high dimensional covariance estimators are consistent to $\boldsymbol{\Sigma}_{t}^{f}$ under the spectral norm if $\log(p)=o(n)$ or $p=o(e^{cn})$ for any positive constant $c$.
-According to Proposition 3.1, the convergence results in Proposition 3.2 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$.
+According to Lemma 3.1, the convergence results in Proposition 3.1 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$.
 Specifically,
-[…]
-[…]
-[…]
-As will be shown in the following, the estimation error of the Kalman gain matrix $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ plays a crucial role in quantifying the effects of the high dimensional  banding, the mid-banding, the tapering and the thresholding estimators for $\boldsymbol{\Sigma}_{t}^{f}$.
+$\text{ if }\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)\text{or }\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right),\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\asymp(\frac{\log p}{n})^{\frac{\alpha}{\alpha+1}}\,\text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k}(\mathbf{S}_{t}^{f})\text{ or }B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f});$
+$\text{ if }\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right),\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\asymp\min\big{\{}n^{\frac{-2\alpha}{2\alpha+1}}+\frac{\log p}{n},\;\frac{p}{n}\big{\}}\text{ for }\,\hat{\boldsymbol{\Sigma}}_{t}^{f}=\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k);$
+$\text{ if }\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right),\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\asymp c_{0}^{2}(p)(\frac{\log p}{n})^{(1-\gamma)}\ \text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=T_{s}(\mathbf{S}_{t}^{f}).$ (3.1)
+As will be shown later, the Kalman Gain estimation error $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ plays a crucial role in quantifying the effects of the high dimensional  banding, mid-banding, tapering and thresholding estimators for $\boldsymbol{\Sigma}_{t}^{f}$.
 3.2 Convergence of High Dimensional Ensemble Kalman Filter
 When the dimension $p$ of $\mathbf{x}_{t}$ is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 3.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$.
-Based on the consistent estimators of $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{K}_{t}$, we propose the HD-EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$.
+Based on the consistent estimators of $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{K}_{t}$, we propose the HD-EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ in the following.
 To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
 $\Omega_{t}=$ $\operatorname{tr}\big{(}{n}^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({{\mathbf{H}}_{t}\boldsymbol{\Sigma}_{t}^{f}{\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t}})\big{)}+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f})\|^{2}$
-$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}/n\big{)}+2\operatorname{tr}\big{(}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}\big{)}.$ (3.2)
-The following Theorem 3.1 presents the upper bounds of the discrepancy between the EnKF and the Kalman Filters for the one-step data assimilation.
+$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}/n\big{)}+2\operatorname{tr}\big{(}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}\big{)}.$
+The following Theorem 3.1 presents the upper bounds of the discrepancy between the EnKF analysis state and the underlying Kalman Filter analysis for the one-step data assimilation for a general case where the model operator $\mathcal{M}_{t}$ is possibly nonlinear.
 Theorem 3.1.
@@
 Then, the forecast and analysis states of the EnKF at time $t$ satisfy
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}$ $=$ ${n}^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}n^{-1}\big{)},$
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $=$ $\operatorname{tr}\big{(}n^{-1}p^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+p^{-1}\Omega_{t}$
-$=$ $O\big{(}n^{-1}+p^{-1}\Delta_{\mathbf{K},t}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}.$
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f}\|^{2}|\mathbf{y}_{1:t-1}\}={n}^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}n^{-1}\big{)},$
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=\operatorname{tr}\big{(}n^{-1}p^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+p^{-1}\Omega_{t}$ (3.2)
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\,=O\big{(}n^{-1}+p^{-1}\Delta_{\mathbf{K},t}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}.$
 Theorem 3.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (3.1).
+Note that the first term in (3.2), $n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=n^{-1}\boldsymbol{\Sigma}_{t}^{a}$, is the variance of the sample mean of the analysis ensemble using the true forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the second term $\Omega_{t}$ results from the estimation error of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$.
 Specifically, with the banding or mid-banding estimators, the tapering estimator and the thresholding estimator, the one-step EnKF data assimilation MSE are, respectively,
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\},$ (3.4)
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}q(n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p})\big{\}},$ (3.5)
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}qc_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}\big{\}}.$ (3.6)
-For fixed $p,q$, both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$ as $\Delta_{\mathbf{K},t}=O(n^{-1})$, as also shown in Theorem S.3.1 of the SM.
-Furthermore, Proposition S.3.1 implies that with linear model operator $\mathcal{M}_{t}$ and Gaussian initial state, the asymptotic distributions of $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}$ in the EnKF are the same as those of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t}|\mathbf{y}_{1:t-1}$ in the Kalman Filter, respectively.
-To show the impact of the dimensionality on the MSE, we take the banding and mid-banding estimators as examples.
-If $q\asymp p$, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O((n^{-1}{\log p})^{2\alpha/(2(\alpha+1))})$, which implies that the one-step assimilation MSE mainly depends on the estimation error of the Kalman gain matrix.
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\},$ (3.3)
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}q(n^{\frac{-2\alpha}{2\alpha+1}}+{n}^{-1}{\log p})\big{\}},$ (3.4)
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}qc_{0}^{2}(p)(n^{-1}{\log p})^{(1-\gamma)}\big{\}}.$ (3.5)
+It should also be noted that the above results are valid for general model operators $\mathcal{M}_{t}$ given in Assumption 1-(i) and (ii).
+Compared with the results in Al-Ghattas and Sanz-Alonso (2024), they investigated the difference between the ensembles of single-step Ensemble Kalman Inversion using the localized estimation and the true forecast error covariance matrices, under the framework of the effective dimension of a matrix $\mathbf{A}$ defined by $\operatorname{tr}(\mathbf{A})/\|\mathbf{A}\|$ (or the max-log effective dimension for a sparse covariance).
+For fixed $p,q$, both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$ because $\Delta_{\mathbf{K},t}=O(n^{-1})$, as also shown in Theorem S.3.1 of the SM.
+Furthermore, for linear model operator ${\mathbf{M}}_{t}$, Proposition S.3.1 of the SM implies that with Gaussian initial state, the asymptotic conditional distributions of the analysis ensemble $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}$ and the forecast ensemble $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}$ in the EnKF are the same as those of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t}|\mathbf{y}_{1:t-1}$ in the Kalman Filter, respectively.
+To show the impact of the dimensionality on the assimilation MSE, we take the banding and mid-banding estimators as examples.
+If $q\asymp p$, $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O((n^{-1}{\log p})^{2\alpha/(2\alpha+2)})$, which implies that the one-step assimilation MSE mainly depends on the estimation error of the Kalman Gain.
 Commonly in high resolution data assimilation, the dimension of the observation $\mathbf{y}_{t}$ is much smaller than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$.
-If $q$ is relatively large such that $p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\}\gg n^{-1}$, then the estimation error of the Kalman gain matrix will dominate, that is, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$.
+If $q$ is relatively large such that $p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\}\gg n^{-1}$, then the estimation error of the Kalman Gain matrix will dominate, that is, $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$.
 Otherwise, the first term $n^{-1}$ dominates.
-For the impact of the observation dimension $q$ on the MSE, a study reported in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease as the first term in (3.1) will decrease as $q$ increases.
-For the discrepancies between the EnKF analysis state and the true state, as $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ and $\mathbf{x}_{t}$ is a random variable with $\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ which is a constant not related to the ensemble size $n$ for a fixed $p$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ can not converge to zero.
-In practice, the data assimilation process can be repeated for multiple time steps ahead to make longer-time predictions.
+However, if sample covariance $\mathbf{S}_{t}^{f}$ is used in $\hat{\mathbf{K}}_{t}$ when $p\geq n$, then $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O(p/n^{2}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2})=O(pq/n^{2})$.
+For the impact of the observation dimension $q$ on the MSE, a study in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease since the first term in (3.2) decreases more.
+For the discrepancies between the EnKF analysis state and the true state, as $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}=\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}+\mathbb{E}\{\|\mathbf{x}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}$ and $\mathbf{x}_{t}$ is a random variable with $\mathbb{E}\{\|\mathbf{x}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ which is a constant not related to the ensemble size $n$ for a fixed $p$, then $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}$ can not converge to zero.
+This is due to the fact that the underlying state is a random variable.
+In contrast, $\mathbf{x}_{t}^{a}$ is a conditional mean and is possible to be estimated consistently.
+In practice, the ensemble forecast-analysis process in the EnKF repeats for multiple time steps ahead to make longer-time predictions.
+Thus, in addition to the one-step ahead data assimilation accuracy, one may also care about the quality of multiple-step forecasting and analysis.
 The following Theorem 3.2 gives the $h$-step MSE bounds between the forecast/analysis states of the EnKF in comparison to the underlying Kalman Filters.
@@
 $C_{t,h}^{A}=\frac{1-L_{t,h}^{h}A_{t,h}^{h}}{1-L_{t,h}A_{t,h}}\,\text{ and }C_{t,h}^{B}=\frac{1-L_{t,h}^{h}B_{t,h}^{h}}{1-L_{t,h}B_{t,h}},$ (3.6)
-where $A_{t,h}=\max_{t<k\leq t+h}\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|$, $B_{t,h}=\max_{t<k\leq t+h}\{\mathbb{E}\|\mathbf{I}-\hat{\mathbf{K}}_{k}\mathbf{H}_{k}\|\}$, and $L_{t,h}=\max_{0<k\leq h}\{\ell_{0},\ell_{t+k}\}$ for a fixed $t$, with $\ell_{0}$ and $\ell_{t}$ defined in Assumption 1.
+where $A_{t,h}=\max_{t<k\leq t+h}\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|$, $B_{t,h}=\max_{t<k\leq t+h}\{\mathbb{E}\|\mathbf{I}-\hat{\mathbf{K}}_{k}\mathbf{H}_{k}\|\}$, and $L_{t,h}=\max_{0<k\leq t+h}\{\ell_{0}$,
+$\ell_{k}\}$ with $\ell_{0}$ and $\ell_{k}$ being respectively the Lipschitz coefficient and the upper bound of the local linearization coefficient of the model operation defined in Assumption 1.
 We further denote
-$U_{k}=\big{\{}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\Delta_{\mathbf{K},k}+\operatorname{tr}\big{(}\mathbb{E}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})^{\mathrm{T}}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})({\mathbf{H}}_{k}\boldsymbol{\Sigma}_{k}^{f}{\mathbf{H}}_{k}^{\mathrm{T}}+\mathbf{R}_{k})\big{)}\big{\}}^{1/2}.$ (3.8)
+$U_{k}=\big{\{}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\Delta_{\mathbf{K},k}+\operatorname{tr}\big{(}\mathbb{E}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})^{\mathrm{T}}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})({\mathbf{H}}_{k}\boldsymbol{\Sigma}_{k}^{f}{\mathbf{H}}_{k}^{\mathrm{T}}+\mathbf{R}_{k})\big{)}\big{\}}^{{1}/{2}}.$
 Theorem 3.2.
@@
 By repeatedly performing the ensemble forecast-analysis process $h$ steps, the forecast and analysis states of the EnKF at the time $t+h$ satisfy
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f})|\mathbf{y}_{1:t+h-1}\|^{2}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{f})\}^{1/2}+L_{t,h}C_{t,h-1}^{B}\max_{t<k\leq t+h-1}U_{k}\Big{)}^{2}$
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f}\|^{2}|\mathbf{y}_{1:t+h-1}\}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{f})\}^{1/2}+L_{t,h}C_{t,h-1}^{B}\max_{t<k\leq t+h-1}U_{k}\Big{)}^{2}$
 $=O\Big{(}n^{-1}+p^{-1}L_{t,h}^{2}(C_{t,h-1}^{A})^{2}\max_{t<k\leq t+h-1}\big{\{}\Delta_{\mathbf{K},k}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\big{\}}\Big{)},$
@@
 $=O\Big{(}n^{-1}+p^{-1}(C_{t,h}^{A})^{2}\max_{t<k\leq t+h}\big{\{}\Delta_{\mathbf{K},k}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\big{\}}\Big{)}.$
-Compared with the one-step results given in Theorem 3.1, the term related to the estimation error of $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the $h$-step analysis MSE is inflated with constants $(C_{t,h}^{A})^{2}$ which are related to $L_{t,h}$ and $A_{t,h}$ defined before (3.7).
-However, for any fixed $h$, $\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}\asymp\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}.$
+Compared with the one-step results in Theorem 3.1, the term related to the estimation error in $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the $h$-step analysis MSE is inflated with constants $(C_{t,h}^{A})^{2}$ which are related to $L_{t,h}$ and $A_{t,h}$ defined after (3.6).
+This is because of the accumulation of the estimation error of $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the data assimilation, which can be seen from the proof of Theorem 3.2 in the SM.
+Using a similar analysis on the one-step results after Theorem 3.1 and according to (3.1), the $h$-step assimilation MSE between the analysis state of the EnKF $\hat{\mathbf{x}}_{t+h}^{a}$ and the $\mathbf{x}_{t+h}^{a}$ for $\boldsymbol{\Sigma}_{t}^{f}$ estimated by the four high dimensional estimators are, respectively,
+$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;(C_{t,h}^{A})^{2}{p}^{-1}q\big{(}{n}^{-1}{\log p}\big{)}^{\alpha/(\alpha+1)}\big{\}},$ (3.7)
+$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;(C_{t,h}^{A})^{2}{p}^{-1}q(n^{-\frac{2\alpha}{2\alpha+1}}+{n}^{-1}{\log p})\big{\}},$ (3.8)
+$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;(C_{t,h}^{A})^{2}{p}^{-1}qc_{0}^{2}(p)\big{(}n^{-1}{\log p}\big{)}^{(1-\gamma)}\big{\}},$ (3.9)
+in which the estimation error of $\mathbf{K}_{k}$ are inflated by $(C_{t,h}^{A})^{2}$ compared with that in the upper bound of the one-step assimilation MSE given in (3.3) - (3.5) in Theorem 3.1.
+According to Remark 1, $A_{t,h}=\max_{t<k\leq t+h}{\left\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\right\|}<1$.
+If $L_{t,h}A_{t,h}<1$, then $C_{t,h}^{A}$ converges to a small constant and $L_{t,h}^{h}A_{t,h}^{h}\rightarrow 0$ as $h\rightarrow\infty$, although $C_{t,h}^{A}$ becomes larger along with the increase of $h$.
+Thus, $\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$.
+If $L_{t,h}A_{t,h}>1$, then $C_{t,h}^{A}$ increases a lot with the increase of $h$, which suggests the quality of the data assimilation deteriorates as $h$ increases, although $\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ for a fixed $h$.
+However, it is possible to make $L_{t,h}$ be smaller by shortening the forecast time step or the integration time step.
+This underscores the necessity of the high-quality or high dimensional estimators of the forecast error covariance, considering the accumulation of estimation errors in multiple-step data assimilation.
+The MSE between the forecast state of the EnKF $\hat{\mathbf{x}}_{t+h}^{f}$ and the Kalman Filter $\mathbf{x}_{t+h}^{f}$ for the banding, mid-banding, tapering and thresholding estimators of $\boldsymbol{\Sigma}_{t}^{f}$ are similar to (3.7)-(3.9) but with slightly different inflation constants.
 Sometimes, the assimilation process proceeds every $k$ time steps not every step as one may not have observations at each step or the assimilation in every step is too computationally consuming.
@@
 […]
-Theorem 4.1 presents the one-step assimilation bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter when $\mathcal{M}_{t}$ is unknown. Denote
+Theorem 4.1 gives the one-step assimilation bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter when $\mathcal{M}_{t}$ is unknown. Denote
 $V_{t}$ $=\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\Xi_{\xi_{t}}\big{)}+\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2}$
-$+\operatorname{tr}\big{(}n^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({\mathbf{H}}_{t}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}}){\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t})\big{)}$
-$+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}-{\mathbf{H}}_{t}\mu_{\xi_{t}})\|^{2}+2\operatorname{tr}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}$ (4.1)
+$+\operatorname{tr}\big{(}n^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({\mathbf{H}}_{t}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}}){\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t})\big{)}$ (4.1)
+$+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}-{\mathbf{H}}_{t}\mu_{\xi_{t}})\|^{2}+2\operatorname{tr}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}$
 $-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\{\mathbf{H}_{t}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}})/n+(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}-{\mathbf{H}}_{t}\mu_{\xi_{t}})\mu_{\xi_{t}}^{\mathrm{T}}\}\big{)}.$
 Theorem 4.1.
-Suppose Assumptions 1(i) or 1(ii), 2, 3, 4 hold and $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t-1}^{a}$ and covariance $\boldsymbol{\Sigma}_{t-1}^{a}$ at time $t-1$.
-Then, the forecast and analysis states of the EnKF satisfy
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=n^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}})+p^{-1}\|\mu_{\xi_{t}}\|^{2},$ (4.2)
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=p^{-1}\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+V_{t}$ (4.3)
+[…]
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f}\|^{2}|\mathbf{y}_{1:t-1}\}=n^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}})+p^{-1}\|\mu_{\xi_{t}}\|^{2},$ (4.2)
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=p^{-1}\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+V_{t}$ (4.3)
 $=O\Big{(}n^{-1}+p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2}+p^{-1}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{(}\|\Xi_{\xi_{t}}\|^{2}+\tilde{\Delta}_{\mathbf{K},t}\big{)}\Big{)}$
-where $\tilde{\Delta}_{\mathbf{K},t}=\mathbb{E}\|\tilde{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ and $\tilde{\mathbf{K}}_{t}$ is the estimate of $\mathbf{K}_{t}$ with true model $\mathcal{M}_{t}$.
+where $\tilde{\Delta}_{\mathbf{K},t}=\mathbb{E}\|\tilde{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ and $\tilde{\mathbf{K}}_{t}$ is the estimated $\mathbf{K}_{t}$ with true model $\mathcal{M}_{t}$.
 The theorem suggests that the model approximation error creates extra terms related to $\mu_{\xi_{t}}$ and $\|\boldsymbol{\Xi}_{\xi_{t}}\|$, which are absent from the results in Theorem 3.1 where accurate model is assumed.
 Specifically, the assimilation MSEs for fixed $p,q$ and $p,q\rightarrow\infty$ are, respectively,
-$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}\text{ and }$ (4.4)
-$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
+$\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}$ (4.4)
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
 […]
+$\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O\big{(}n^{-1}+\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}(\|\Xi_{\xi_{t}}\|^{2}+n^{-1})\big{)}=O\big{(}n^{-1}+\|\Xi_{\xi_{t}}\|^{2})\big{)}.$ (4.5)
+Thus, if $\|\Xi_{\xi_{t}}\|^{2}=O(n^{-1})$, then $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O(n^{-1})$, that is, the model approximation error does not alter the order of magnitude of the assimilation MSEs of the EnKF.
+However, if $n\|\Xi_{\xi_{t}}\|^{2}\rightarrow\infty$, then $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O(\|\Xi_{\xi_{t}}\|^{2})$, which means the model approximation error dominates the assimilation MSE.
+When $p,q\rightarrow\infty$, $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}\asymp\max\{n^{-1},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
+[…]
+$\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}\text{ and }$ (4.6)
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
 The above analysis shows that for both the fixed and diverging dimensions and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the assimilation MSE, which suggests that a good approximation of the model operator is a significant aspect of data assimilation.
 While this may be obvious, Theorem 4.1 describes exactly how it happens.
-The accumulation of the model approximation error is more pronounced in the $h-$step assimilation MSEs as shown in Theorem S.3.3 of the SM.
-Thus, an iterative HD-EnKF algorithm is introduced in S.1.2 to account for the model approximation error to some extent since most geophysical models are imperfect in practice, of which the main idea is to substitute $\hat{\mathbf{x}}_{t}^{f}$ in the original forecast sample covariance matrix by the analysis state $\hat{\mathbf{x}}_{t}^{a}$ in the iteration.
-5 Simulation
+The accumulation of the model approximation error is more pronounced in the $h$-step assimilation MSEs as shown in Theorem 4.2 and more details of the $h$-step assimilation in the presence of the model error can be found in Theorem S.3.4 of the SM.
+Define $D_{t,h}=\max_{t<k\leq t+h}{\left\|\mathbf{I}-\check{\mathbf{K}}_{k}\mathbf{H}_{k}\right\|}$, where ${\check{\mathbf{K}}}_{k}=(\boldsymbol{\Sigma}_{k}^{f}+\Xi_{k})\mathbf{H}_{k}^{\mathrm{T}}\big{(}\mathbf{H}_{k}(\boldsymbol{\Sigma}_{k}^{f}+\Xi_{\xi_{k}})\mathbf{H}_{k}^{\mathrm{T}}+\mathbf{R}_{k}\big{)}^{-1}$, which is the underlying Kalman Gain matrix under the imperfect model.
+Denote a non-stochastic constant $C_{t,h}^{D}=(1-L_{t,h}^{h}D_{t,h}^{h})/(1-L_{t,h}D_{t,h})$.
+Theorem 4.2.
+[…]
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f}\|^{2}|\mathbf{y}_{1:t+h-1}\}$
+$=O\Big{(}n^{-1}+p^{-1}\|\mu_{\xi_{t+h}}\|^{2}+p^{-1}L_{t,h}^{2}(C_{t,h-1}^{D})^{2}\max_{t<k\leq t+h-1}\big{\{}\Delta_{\check{\mathbf{K}}_{k}}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}-\mathbf{H}_{k}\mu_{\xi_{k}}\|^{2}\big{\}}\Big{)},$
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}$
+$=O\Big{(}n^{-1}+p^{-1}\|(\mathbf{I}-\mathbf{K}_{t+h}\mathbf{H}_{t+h})\mu_{\xi_{t+h}}\|^{2}+p^{-1}\|\Xi_{\xi_{t+h}}\|^{2}\|\mathbf{y}_{t+h}-\mathbf{H}_{t+h}\mathbf{x}_{t+h}^{f}\|^{2}$
+$\quad+p^{-1}(C_{t,h}^{D})^{2}\max_{t<k\leq t+h}\big{\{}\Delta_{\check{\mathbf{K}}_{k}}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}-\mathbf{H}_{k}\mu_{\xi_{k}}\|^{2}\big{\}}\Big{)}.$
+Compared with the result in Theorem 3.2 under the correct model, there are extra terms related to $D_{t,h}=O(\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|+\|\Xi_{\xi_{k}}\|\|\mathbf{H}_{k}\|)$, $\|\mu_{\xi_{t+h}}\|$ and $\|\Xi_{\xi_{t+h}}\|$ resulting from the accumulation of the approximation error of the model operator $\mathcal{M}_{t}$ for the $h$-step assimilation MSE in Theorem 4.2.
+Compared with the one-step assimilation MSE in Theorem 4.1 under the imperfect models, the error term related to the approximation error of $\mathcal{M}_{t}$ is inflated by $(C_{t,h}^{D})^{2}$.
+The $h$-step assimilation MSE for the imperfect models under the fixed and high dimensions are, respectively,
+$\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\max\big{\{}\big{(}1+(C_{t,h}^{D})^{2}\big{)}n^{-1},\;\|\mu_{\xi_{t+h}}\|^{2},\;\|\Xi_{\xi_{t+h}}\|^{2}\big{\}},$
+$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\max\big{\{}n^{-1},\;p^{-1}\|\mu_{\xi_{t+h}}\|^{2},\;p^{-1}{q}\|\Xi_{\xi_{t+h}}\|^{2},\;(C_{t,h}^{D})^{2}p^{-1}{q}\Delta_{\check{\mathbf{K}}_{k}}\big{\}}.$ (4.7)
+Specifically, for the high dimensional case, the MSEs of the EnKF analysis state $\hat{\mathbf{x}}_{t+h}^{a}$ to the Kalman Filter $\mathbf{x}_{t+h}^{a}$ for the $h$-step assimilation using the banding or mid-banding, tapering and thresholding estimates of $\boldsymbol{\Sigma}_{t}^{f}$ can be obtained by replacing $\Delta_{\check{\mathbf{K}}_{k}}$ in equation (4.7) with the specific convergence rates $({n}^{-1}{\log p})^{\alpha/(\alpha+1)}$, $n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p}$ and $c_{0}^{2}(p)(n^{-1}{\log p})^{(1-\gamma)}$, respectively, according to (3.1).
+For the $h$-step assimilation, the approximation error of the model operator $\mathcal{M}_{t}$ has a stronger influence on the MSE of the data assimilation of the EnKF resulting from the accumulation of the approximation error.
+This highlights that a good approximation of the model operator is significant to data assimilation.
+In summary, the approximation error associated with the model operator $\mathcal{M}_{t}$ significantly affects the MSE in the assimilation process of the EnKF, particularly in scenarios involving multiple-step ahead data assimilation.
+Thus, an iterative HD-EnKF algorithm is introduced in S.1.2 of the SM to account for the model approximation error to some extent since most geophysical models are imperfect in practice, of which the main idea is to substitute $\hat{\mathbf{x}}_{t}^{f}$ in the original forecast sample covariance matrix by the analysis state $\hat{\mathbf{x}}_{t}^{a}$ in the iteration.
+5 Simulation Studies
 This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed HD-EnKF algorithms.
-These two models are often used for testing new approaches in atmospheric sciences for both data assimilation and the spatial-temporal forecasts.
-The proposed HD-EnKF methods are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models.
+These two models are often used for testing new approaches in atmospheric sciences for data assimilation and the spatial-temporal forecasts.
+The proposed HD-EnKFs are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models.
 Details of the proposed HD-EnKF algorithms are given in Section S.1 of the SM, including the consistent HD-EnKF update in S.1.1 and the iterative HD-EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators.
+The iterative HD-EnKF algorithm in S.1.2 of the SM is to account for the model approximation error in the imperfect models.
 5.1 Lorenz-96 model
 Lorenz-96 model (L96; Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere.
-It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation
-${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),\,\text{ for }j=1,\ldots,p,$ (5.1)
-with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denotes the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $-\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is a forcing term.
-L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016).
+It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation, for $j=1,\ldots,p$,
+${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),\,$ (5.1)
+with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denotes the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is a forcing term.
+L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016), which approximates the model as $\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)=\mathbf{x}_{t}+h(k_{1}+2k_{2}+2k_{3}+k_{4})/{6},$ where $k_{1}=\mathbf{f}\left(\mathbf{x}_{t},t\right)$, $k_{2}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{1}/2,t+h/2\right)$, $k_{3}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{2}/2,t+h/2\right)$, $k_{4}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{3},t+h\right)$ and $\mathbf{f}=\left(f_{1},f_{2},\dots,f_{p}\right)^{\mathrm{T}}$.
 An approximation to the true state was simulated by (i) implementing the RK4 method with a finer time step, say $h=0.05$; and (ii) adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$.
@@
 The $(i,j)$-element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,q-|i-j|\right\}}$, which prescribes spatially correlated observation errors.
-The observation operator $\mathbf{H}_{t}$ was a $q\times p$ matrix with the $(j,i)$ element being $1$ if and only if the $j-$th component of $\mathbf{y}_{t}$ was the observed $i-$th dimension of the state $\mathbf{x}_{t}$.
+The observation operator $\mathbf{H}_{t}$ was a $q\times p$ matrix with the $(j,i)$ element being $1$ if the $j$-th component of $\mathbf{y}_{t}$ was the observed $i$-th dimension of the state $\mathbf{x}_{t}$.
 We simulated observations every $4$ time steps for a total of $2000$ steps and reported the results for the last $1000$ steps to allow the system to settle down.
@@
 All the experiments were repeated 500 times.
-We considered the HD-EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance to estimate the forecast error covariance and the inflation methods (Liang et al., 2012).
+We considered the HD-EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance and the inflation methods (Liang et al., 2012).
 The performances of these methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$.
-To gain accurate estimations of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance matrices to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman filter update formula to obtain $\mathbf{x}_{t}^{a}$.
+To gain accurate estimations of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman Filter update formula to obtain $\mathbf{x}_{t}^{a}$.
 Figure 2: Average RMSEs to the oracle analysis state and their $25-75\%$ quantile bands (color bands) as a function of forcing $F$ used in ${\cal{M}}_{t}$ (true value is 8) by the standard EnKF (blue), the modified EnKF methods: the inflation (red), banding (green), tapering (cyan) and thresholding (purple) for $p=q=40$ and $100$, $n=20,30,60,90$, $\sigma_{0}=0.1,0.25$ and $0.5$, respectively.
@@
 It shows that the HD-EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) models.
-Even when $F$ deviated more from the true value $8$, the HD-EnKFs still achieved analysis states that were much closer to the oracle than the inflation method although both strains of methods showed increased RMSEs with respect to $|F-8|$.
-The standard EnKF with the sample covariance estimator to $\boldsymbol{\Sigma}_{t}^{f}$ had much larger assimilation errors in all cases.
+Even when $F$ deviated more from the true value $8$, the HD-EnKFs still achieved analysis states much closer to the oracle than the inflation method, although both strains of methods showed increased RMSEs with respect to |F8||F-8|.
+The standard EnKF with the sample covariance estimator had much larger assimilation errors in all cases.
 Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), the advantages of the proposed HD-EnKFs over the inflation method were more apparent for the higher dimension $p=100$.
@@
 The figure also displays the percentages of divergence of the five methods.
-It shows that the inflation method endured quite high divergences which increased substantially with the model error (e.g.
+The inflation method endured quite high divergences which increased substantially with the model error (e.g.
 reached 50% when $\sigma=0.5$), even higher than the standard EnKF.
@@
 Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ of dimension $p=3p^{\prime}=4650$.
-The initial height (depth) field was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~{}m},5.5\mathrm{~{}m}$ and $3.325\mathrm{~{}m}$, respectively.
+The initial height (depth) was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~m},5.5\mathrm{~m}$ and $3.325\mathrm{~m}$, respectively.
 The initial velocity field was derived from the initial height field with the geostrophic relation.
@@
 The ensemble size $n$ was 100, and the initial ensemble was generated by adding random noises from $\mathrm{N}_{p}(0,\mathbf{I}_{p})$ to the initial state $\mathbf{x}_{1}$.
-Figure 3: RMSEs to the oracle analysis state (the left panels) and the true state (the right panels) at each assimilation time step by the standard, inflation, banding, tapering and threshold under the correct model ($k=5\times 10^{4}$) and imperfect model ($k=10^{4}$) with $n=100$, $p=4650$ and $q=930$ on the SWE model.
-Table 2: Average RMSEs to the oracle analysis state and true state over the last 24 hours (2880 steps) of the simulation experiments and their $25\%$ and $75\%$ quantiles (in parentheses) by the five EnKF schemes for the SWE model under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models with $n=100$, and $p=4650$ and $q=930$.
+Figure 3: RMSEs to the oracle analysis state (left panels) and the true state (right panels) at each assimilation step by the standard, inflation, banding, tapering and threshold under the correct model ($k=5\times 10^{4}$) and imperfect model ($k=10^{4}$) with $n=100$, $p=4650$ and $q=930$ on the SWE model.
+Table 2: Average RMSEs to the oracle analysis state and true state over the last 24 hours (2880 steps) and their $25\%$ and $75\%$ quantiles (in parentheses) by the five EnKF schemes for the SWE model under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models with $n=100$, and $p=4650$ and $q=930$.
 method k = 50000 (correct model) k = 10000 (imperfect model)
@@
 Figure 3 shows the three proposed HD-EnKF methods outperformed the inflation and the standard EnKF methods for both the correct and imperfect models with much smaller RMSEs after the system settled down and that the superiority became more apparent over time, especially under the misspecified model.
-Specifically, the RMSEs at each assimilation time step of the three proposed HD-EnKF schemes became lower and lower over time compared to increased RMSEs for the standard EnKF for all cases and slight increases RMSEs for the inflation method under true models.
+Specifically, the RMSEs at each assimilation time step of the three proposed HD-EnKF schemes became lower and lower over time compared to increased RMSEs for the standard EnKF for all cases and for the inflation method under true models.
 It is obvious from Table 2 that the HD-EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands with more evident superiority under imperfect models.
-Specifically, the proposed HD-EnKFs were able to reduce the RMSE by 41% compared to the inflation method, implying that the proposed methods could attain more accurate and robust assimilated outcomes even under extremely sparse observations and ultra-high dimensional state variables.
-Thus, the proposed methods are highly appealing, especially considering that geophysical models in earth science often possess ultra-high dimensions.
-6 Sea Temperature Assimilation in Northwest Pacific
-This section demonstrates how the proposed method is applied to sea temperature assimilation based on the oceanography model named Regional Oceanic Modeling System (ROMS) with satellite data.
-ROMS is a three-dimensional, free-surface, terrain-following ocean general circulation model that solves the Reynolds-averaged Navier–Stokes equations using the hydrostatic and Boussinesq assumptions (Shchepetkin and McWilliams, 2005; Haidvogel et al., 2008).
-While the satellites produce high-resolution sea surface temperature data, they can not provide ocean measurement beneath the ocean surface where the physical model ROMS plays a key role.
-The data assimilation was conducted from January 1st, 2008 to March 31st, 2008 over a study area 25.5N-34.7N, 144.4E-154E.
-The spatial resolution of the ROMS was $0.56^{\circ}\times 0.49^{\circ}$ with $18\times 20$ horizontal grids and $18$ uneven vertical layers over the water column, making the dimension $p$ for the state variable of sea temperature ($\mathbf{x}_{t}$) to be $6480=18\times 20\times 18$.
-To scale down the computation, we did not include the other ocean variables namely salinity and current velocity as the assimilated variables, and let them be internally determined by ROMS at each step with the sea temperature as the assimilated variables.
-The assimilation time length was daily which was also the frequency of the observations.
-The satellite data were the daily Optimum Interpolation Sea Surface Temperature (OISST; https://www.ncei.noaa.gov/products).
-OISST is a long-term data records for Sea Surface Temperature (SST) that incorporate satellite observations into regular global grids using ships and buoys observations correct satellite SST biases (Banzon and Reynolds, 2013).
-As the OISST only provides sea surface temperature, the observation dimension was $q=360$.
-To run ROMS, we need to input the initial state of sea temperature $\mathbf{x}_{1}$, the boundary conditions and smoothed bathymetry data (sea bottom topography).
-The bathymetry data came from ETOPO2 (https://www.ncei.noaa.gov/products/etopo-global-relief-model), which is a global map of the earth’s seafloor and land elevations as shown in Figure S.1 of the SM.
-The initial state ($\mathbf{x}_{1}$) was from the Global Ocean Reanalysis and Simulations (GLORYS; https://data.marine.copernicus.eu/products) on January 1st.
-Thorough descriptions of boundary conditions of ROMS can be found in Marchesiello et al.
-(2001).
-As the ocean temperature, salinity and current fields have different spatial scales and resolution in the three dimension study region both in the ROMS and the physical ocean, we transform the study region to a scale down space to facilitate the bandwidth selection in the banding estimator.
-Specifically, the grid points in the study region are denoted by coordinates $\{(u,v,h),u=1,\cdots 18,v=1,\cdots 20,h=1,\cdots 18\}$ in the transformed space of unit rectangles, where the banding estimator is performed.
-The temperatures on the physical grids are labelled as $\{x_{u,v,h}\}$ in the standardized grids and the whole state vector can be labelled by $\mathbf{x}=(x_{1,1,1},\cdots,x_{18,20,18})^{\mathrm{T}}$ with the subscript $t$ hidden.
-The banding estimator with bandwidth $k$ means that we keep the covariances between $x_{u,v,h}$ and $x_{u^{\prime},{v^{\prime}},h^{\prime}}$ on the standardized grids if they satisfy $\max\{|u^{\prime}-u|,|v^{\prime}-v|,|h^{\prime}-h|\}\leq k$.
-We performed the HD-EnKF using the banding estimator with an ensemble size of $n=30$ to obtain the assimilated three-dimensional sea temperature for the study period and region.
-We also compared the proposed method with the standard EnKF and the inflation method.
-The out-of-sample performances of these methods were assessed on data collected under the ocean Argo project over the study period and region, accessed at https://argo.ucsd.edu/data/.
-The Argo project collects information from inside the ocean using floats that initially drift with the ocean currents and then surface up and down between the surface and mid-sea to collect measurements.
-As the observations provided by Argo are in-situ observations at various depths, they are more precise than the remote sensing data by the satellites, and hence, are used as substitutes for the ground truth for the out-of-sample assessment.
-The bandwidth $k$ used in the banding estimator was $k=1$, selected by minimizing the out-of-sample RMSE over the first month of the study period between assimilated temperatures and temperatures recorded by Argo at different depths.
-The inflation factor for the inflation method was similarly selected, which was $1.1$.
-Figure 4: Daily assimilation RMSEs and correlations to the OISST(left panels) and the Argo data (right panels) from Jan.
-1st 2008 to Mar.
-31st 2008 using the standard EnKF (blue), inflation (red) and HD-EnKF with banding (green), respectively.
-The dashed horizontal line represents the average assimilation RMSEs and correlations over the whole assimilation period.
-Table 3: The average assimilation RMSEs and correlations and their standard errors (in the parentheses) to the OISST data and the Argo data over the whole assimilation period.
-Method OISST Argo
-RMSE Corr RMSE Corr
-standard 0.677(0.0067) 0.943(0.0012) 1.242(0.0421) 0.984(0.0013)
-inflation 0.647(0.0071) 0.948(0.0011) 1.244(0.0421) 0.983(0.0012)
-banding 0.277(0.0043) 0.990(0.0003) 1.134(0.0425) 0.986(0.0014)
-Figure 5: The spatial distribution of the assimilated sea surface temperature on January 1st, February 15th and March 31st in 2008 over the study area, with the fill color from red to blue representing sea surface temperature from high to low.
-Figure 4 displays both the in-sample RMSEs and correlations between the assimilated sea surface temperature and the OISST observations, and the out-of-sample RMSEs and correlation between the assimilated sea water temperature and the temperatures recorded by Argos at various depths.
-From Figure 4, the HD-EnKF provided much lower and more stable assimilation RMSEs and much higher correlations in the entire assimilation period than those for the standard EnKF and the inflation methods.
-On the out-of-sample performance on the Argo data, the HD-EnKF produced assimilated sea temperatures with smaller RMSEs and higher correlations both on average and for most days during the assimilation period.
-These imply that the proposed HD-EnKF algorithm could offer more accurate and robust assimilation results in practice.
-Detailed numerical results of these methods are reported in Table 3, which summarizes the average assimilation RMSEs and correlations of the assimilated $\hat{\mathbf{x}}_{t}^{a}$ to the OISST and the Argo data over the whole assimilation period.
-The spatial distribution of the assimilated sea surface temperature over the study area on January 1st, February 15th and March 31st in 2008 is depicted in Figure 5, which varied over time and saw an overall downward trend during the study period.
+Specifically, the proposed HD-EnKFs could reduce the RMSE by 41% compared to the inflation method, implying that the proposed methods could attain more accurate and robust assimilated results even under extremely sparse observations and high dimensional states.
+Thus, the proposed methods are highly appealing, especially considering the high dimensional geophysical models in earth science.
 6 Discussion
@@
 The study shows that the proposed HD-EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations.
-One is that the established theoretical analysis on the consistency and the error bounds makes the proposed HD-EnKF methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies and the empirical study on sea temperature assimilation.
+One is that the established theoretical analysis on the consistency and the error bounds makes the proposed HD-EnKF methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies.
 The latter means that a smaller ensemble size for the EnKF can be used at a given level of accuracy which is particularly attractive as the running of the geophysical models is generally very expensive.
-The proposed HD-EnKF has great potential for applications in earth science where the geophysical models, the state variables or the observations are ultra-high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.
+The proposed HD-EnKF has great potential for applications in earth science where the geophysical models, the state variables or the observations are high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.
 SUPPLEMENTARY MATERIAL